### v1

# Bilateral facial reduction: qualification-free subdifferential calculus and exact duality in convex analysis

Abstract
This paper introduces a geometric framework to extend convex analysis results beyond standard qualification conditions such as intersecting relative interiors of domains. We define the joint facial subspace $T$ as the span of the face of CDC-D generated by 0, with its affine translation $T_{a}$ containing $C\cap D$. The intersections $C\cap T_{a}$ and $D\cap T_{a}$ are faces of the original sets, establishing $T_{a}$ as a bilateral facial reduction with parallels to Borwein-Wolkowicz facial reduction in conic programming. When reducing the domains of two convex functions to their joint facial subspace $T_{a}$, their relative interiors then always intersect, enabling unqualified application of classical theorems via localization. Key generalizations include subdifferential additivity, normal cones of intersections, a subdifferential chain rule, attained infimal convolution for $(f+g)^{*},$ and an exact Fenchel-Rockafellar dual. We characterize $T$ via generated faces and computationally through an iterative process converging in at most $n$ steps in $\mathbb{R}^{n}$. Proofs are self-contained and introduce novel concepts like facial subspaces and nested normals to describe convex boundaries and the lattice of faces.
1 Introduction
In convex analysis, the effective domain of the sum of two convex functions is the intersection of their effective domains. Many central results, including the summation rule of subdifferentials, the convex conjugate of sums of functions, or the strong duality of the Fenchel-Rockafellar dual, exclude the case where the domains of the convex functions intersect only on their relative boundaries. The conditions excluding such cases are called qualification conditions, or constraint qualifications depending on the context. One such condition, widespread in the seminal work [14] by Rockafellar, is that for two functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$,
$\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))\neq\emptyset.$
For instance, this condition appears in [14, Theorem 23.8], which establishes the summation rule of subdifferentials:
$\partial(f+g)(x)=\partial f(x)+\partial g(x).$
A second example is the strong duality of the Fenchel-Rockafellar dual: for $A\in\mathbb{R}^{m\times n}$, denoting by $f^{*}$ the convex conjugate of $f$,
$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)=-\inf_{y\in\mathbb{R}^{m}}[f^{*}(A^{T}y)+g^{*}(-y)].$
In [14, Corollary 31.2.1], this is guaranteed to hold if
$\operatorname{relint}(\operatorname{dom}(f))\cap A^{-1}(\operatorname{relint}\operatorname{dom}(g))\neq\emptyset,$
where $A^{-1}(\cdot)$ is the preimage for the linear map induced by the matrix $A$. This requirement is often called a constraint qualification, and is strikingly similar to the qualification condition for the subdifferential sum rule.
These and a number of other results share similar conditions. The object of this work is to generalize such results beyond the regime where these conditions hold. Specifically, in Section 5 we find generalizations for the two aforementioned results, as well as a few others: the attainment of infimal convolution as the convex conjugate of the sum of two convex functions, the normal cone of the intersection of two convex sets, and a chain rule for the subdifferential. There are many reasons such generalizations are worthwhile, and we list two here. From a theoretical perspective, extending such results makes the theory more elegant and intuitive. A second motivation is that such generalizations could lead to improved methods in optimization, as did facial reduction in conical programming (see [3]).
Qualification conditions are not merely theoretical artifacts; they are supported by counter examples, which indicates that these conditions reflect more or less accurately the true scope of the statements. We consider one such counter-example for the sum rule of subdifferentials. Take the function $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=\infty$ otherwise. This function has effective domain $\mathbb{R}_{+}$. Adding it to the indicator function $g(x)=\iota_{\mathbb{R}_{-}}(x)$, the sum rule states that $\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset$, when in fact, $\partial(f+g)(0)=\partial\iota_{0}(0)=\mathbb{R}$. The relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect. The qualification condition does not hold, so there is no contradiction with [14, Theorem 23.8], yet this still shows that the sum rule does not in fact apply to the sum of any two convex functions. Thus, an accurate generalization of the sum rule must offer a different formula for the subdifferential of the sum of two convex functions, and such a formula should ideally reduce to the classical sum rule when qualification conditions are satisfied. We provide such a formula in Theorem 5.1.
Our approach to relaxing qualification conditions is to focus on the geometry of the effective domain of the convex functions. Qualification conditions fail when relative interiors do not intersect, in which case only the boundaries of the domains meet. We show that the boundaries of two convex sets always meet in a structured way: each set presents a convex face to the other, and both of these faces are revealed by the intersection with a common affine subspace, which we name the joint facial subspace $T_{a}$. The main property of the joint facial subspace is that given convex sets $C,D\subseteq\mathbb{R}^{n}$, with $T_{a}$ defined as in Definition 4.1, whenever $C\cap D\neq\emptyset$,
$\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})\neq\emptyset.$
The results which shows this is Corollary 4.3. Therefore, qualification conditions always hold when localizing to the joint facial subspace. Roughly, we generalize results in convex analysis by replacing convex functions $f$ by their localized version $f+\iota_{T_{a}}$. The modified results then hold even without qualification conditions. And when qualification conditions do hold for the original results, then $T_{a}=\mathbb{R}^{n}$ (see Proposition 5.1), and the additional indicator function can be ignored.
For instance, in Theorem 5.1 we state that
$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x),$
without any condition on $f$ and $g$ other than they be convex functions. This modification works for the previously encountered counter example. In that problem, where $f(x)=-\sqrt{x}$ and $g(x)=\iota_{(-\infty,0]}$, the joint facial subspace from Definition 4.1 is $T_{a}=\{0\}$, and so
$\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x)=\partial\iota_{\{0\}}+\partial\iota_{\{0\}}=\mathbb{R},$
which is indeed $\partial(f+g)(x)$. One way to interpret our result is that there are subgradients that are “hiding” in the faces of the domains of the convex functions. Intersecting with the joint facial subspace reveals these hidden gradients.
The central theoretical concept introduced in this work is the joint facial subspace. We therefore focus much of our efforts in describing this object. This is the focus of Section 4, where we define the joint facial subspace in Definition 4.1 , provide two characterizations and state core property, that qualification conditions hold when localizing to it. Given two convex sets $C,D\subseteq\mathbb{R}^{n}$, the joint facial subspace is defined from the Minkowski difference $C-D=\{c-d\mid c\in C,d\in D\}$. In the first characterization, Theorem 4.1, it is made clear how the joint facial subspace relates to $C$ and $D$ individually. The third characterization is computational in nature, and describes how the joint facial subspace can be found from a number of steps that is bounded by the ambient dimension.
Another important contribution of the present work is in our proof techniques. In Section 6, we build a theoretical framework with nested normals (Definition 6.3) and facial subspaces (Definition 6.2), two objects which together describe the lattice of faces of a convex set. Their combination forms a dual object which is more descriptive than the standard normal cone (see Theorem 6.2). Our theory is therefore at least in part a theory about the boundary of convex sets. From this toolkit, we are able to describe the faces of differences of convex sets CDC-D in Section 6.2, from which we derive the various properties of the joint facial subspace in Section 4. These objects will be further explored and refined in future work.
The structure of the paper goes as follows. In Section 2 surveys prior works and Section 3 introduces some notation. Section 4 is as described above, with proofs deferred to Section 7. In Section 5, we provide generalizations of central results in convex analysis, together with their proofs, which use Corollary 4.3 from Section 4. In Section 6, we give a motivated introduction to our mathematical framework, introducing key concepts. In Section 7, we prove the results in Section 4 by using tools from Section 6.
2 Prior works
The existing research which is the most related to this work is Facial Reduction (FR) in conic programming  [2, 1, 3], which helped guide the development of the theory in this work. We briefly describe this method. Given a feasible cone and affine constraints, Slater’s condition (that the interior of the feasible cone has at least one point satisfying the constraints), FR narrows the ambient space of the problem to the span of a “minimal” face of the feasible cone. After narrowing the problem via this “preprocessing” step, the problem is guaranteed to satisfy Slater’s condition.
Our joint facial subspace does something similar, but for any two intersecting convex sets. We find a “minimal” face of the difference of the sets, CDC-D (see Definition 4.1), and this yields the joint facial subspace. The computation of the joint facial subspace can be similarly viewed as a processing step, after which the problem is reduced to the joint facial subspace. As specified in [3], FR can be applied with finite-step algorithms, with a number of steps bounded by the ambient dimension. This is very similar to our second characterization of the joint facial subspace, Theorem 7.1.
A result which is related (where the relation is examined in [11]), is the Extended Dual System of Ramana, a dual program for semidefinite programming for which strong duality always holds [13]. The corresponding result in our work is the exact Fenchel-Rockafellar dual in Theorem 5.5.
From these strong parallels, we conjecture that many aspects of Facial Reduction and the Extended Dual System can be recovered—and potentially generalized—from results in this work.
Prior efforts to obtain a qualification-free sum rule for subdifferentials have developed powerful limiting approaches that achieve exact characterizations without traditional qualifications. For instance, $\epsilon$-subdifferentials define approximate subgradients satisfying the subgradient inequality up to an $\varepsilon$-error, recovering the exact subdifferential as $\varepsilon$ approaches zero [4, 12, 10]. Similarly, fuzzy sum rules approximate the subdifferential by perturbing the point $x$ at which the subdifferential is taken, deriving exact additivity rules in the limit of vanishing perturbations [6, 5]. These methods, leveraging local neighborhood behavior, are related to ideas in variational analysis, and offer versatile tools for theoretical and computational analysis across general spaces. They can often be implemented with fixed tolerances for practical computation, though the exactness relies on infinite limiting processes to eliminate approximation errors. Though our approach is tailored to pairs of functions, our formula for the subdifferential of a sum of convex functions is exact, and can be computed in finite steps.
One theoretical insight that was critical to developing our theory was the lexicographical characterization of facial subspaces from [8]. This work shows that every face of a convex set can be obtained by iteratively exposing faces. It is a sequel of a preceding result [9], which shows a separation theorem for convex sets. Though the lexicographical framework was not flexible enough for our needs, prompting us to develop an independent theory, we did draw inspiration from the results and the proofs. In this paper, we recover this lexicographical representation, though in different terms, primarily in Lemma 6.5. Our proof is direct, and does not make use of a separation theorem.
3 Preliminaries
For two sets $A,B\subseteq\mathbb{R}^{n}$, $A+B:=\{a+b\mid a\in A,b\in B\}$ is the Minkowski sum, with $A-B:=\{a-b\mid a\in A,b\in B\}$, and for $z\in\mathbb{R}^{n}$, let $A+z:=A+\{z\}$. For a set $S\subseteq\mathbb{R}^{n}$, we denote by $\iota_{S}(x)$ the indicator function which is 0 if $x\in S$ and $\infty$ otherwise.
For a set $S\subseteq\mathbb{R}^{n}$, denote by $\operatorname{conv}(S)$ its convex hull. For $y_{1},y_{2}\in\mathbb{R}^{n}$, let $y_{1},y_{2}]:=\operatorname{conv}\{y_{1},y_{2}\}$, and $y_{1},y_{2}[:=\operatorname{conv}\{y_{1},y_{2}\}\setminus\{y_{1},y_{2}\}$.
Intuitively the notion of a face of a convex set $C$ is a subset of $C$ for which there is no line segment in $C$ that “goes through” $F$. The following definitions can be found inline in [16, p. 16] and [16, p.75] respectively.
Definition 3.1 (Face of a convex set).
Let $C$ be convex. A subset $F\subseteq C$ is defined to be a face of $C$ when
$\forall y,z\in C,(\exists x\in]y,z[\cap F\implies y,z\in F).$
Definition 3.2 (Exposed face of a convex set).
Let $C\subseteq\mathbb{R}^{n}$ be convex. A subset $F\subseteq C$ is defined to be an exposed face of $C$ when
$\exists u\in\mathbb{R}^{n}\setminus\{0\}:F=\operatorname*{\arg\max}_{c\in C}\langle c,u\rangle.$
For any set $S\subseteq C$, we denote by $F_{C}(S)$ the face of $C$ generated by $S$; it is the smallest face of $C$ that contains $S$ (it is also the intersection of all faces containing $S$, which, it can be shown, is itself a face).
We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cap\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
$\partial f(x):=\{v\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{n},\langle v,y-x\rangle\leq f(y)\},$
and the convex conjugate is $f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x)$.
4 The joint facial subspace
In this section we investigate the following object.
Definition 4.1 (The joint facial subspace).
The joint facial subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ is
$T(C,D):=\operatorname{\mathrm{span}}F_{C-D}(0).$
We also define $T_{a}(C,D):=T(C,D)+C\cap D$, the joint affine facial subspace.
The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact a simple translation. This follows from the next corollary.
Corollary 4.1 (Containment of the intersection).
Let $C,D\in\mathbb{R}^{n}$ be convex. Then
$C\cap D-C\cap D\subseteq T(C,D).$
This result follows from Lemma 6.10; we defer the proof of Corollary 4.1 to the end of Section 6.2.
From Corollary 4.1, we see that picking any $x\in C\cap D$, we have $T_{a}=T+x$, independently of the choice of $x$. Corollary 4.1 also implies that $C\cap D\subseteq T_{a}$. The affine subspace $T_{a}$ is therefore the translation of $T$ that contains $C\cap D$.
In this section we provide two characterizations of the joint facial subspace. The next result shows the relation of the joint facial subspace to the individual convex sets, and builds towards the first characterization.
Proposition 4.1 (The joint facial subspace reveals faces).
For convex sets $C,D\subseteq\mathbb{R}^{n}$, and $T_{a}=T_{a}(C,D)$,
$T_{a}\cap C=F_{C}(C\cap D).$
For the first characterization we require a subspace that is related to $F_{C}(C\cap D)$. This subspace, the generated facial subspace, is central to our theory, as is made clear in Section 6. There, we will re-introduce this object in a motivated way, in Definition 6.5.
Definition 4.2 (Preliminary definition of the generated facial subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, define the facial subspace of $C$ generated by $S\subseteq C$ to be
$H_{C}(S):=\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S)).$
The next theorem is our first characterization of the joint facial subspace.
Theorem 4.1 (Characterization through generated faces).
The joint facial subspace of $C,D\subseteq\mathbb{R}^{n}$ is
$T(C,D)=H_{C}(C\cap D)+H_{D}(C\cap D).$
Definition 4.1 and Theorem 4.1 both specify the joint facial subspace via analytical formulas. Our second characterization instead specifies the joint facial subspace as the result of a process with a number of steps bounded by the ambient dimension, which is more amenable to computation. This result, Theorem 7.1, is best stated in terms of nested normals a concept that we introduce in Definition 6.3, and motivate throughout Section 6. For this reason, we defer it to Section 7.3. For the moment, we state a direct corollary which does not require this object.
Corollary 4.2 (Iterative bilateral facial reduction).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex sets and fix any $x\in C\cap D$. Beginning with $T_{0}=\mathbb{R}^{n}$, let
$T_{i+1}=T_{i}\cap(N_{C}(x)\cap[-N_{D}(x)]\cap T_{i})^{\perp}.$
Let $\ell$ be the smallest number such that $T_{\ell+1}=T_{\ell}$. Then:
  1. 1.
    $T_{\ell}=T(C,D)$.
  2. 2.
    $\ell\leq n$.
Now that the joint facial subspace has been thoroughly identified, we move on to describing its core property. Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, a common qualification condition is that for convex sets $C:=\operatorname{dom}(f),D:=\operatorname{dom}(g)$,
$\operatorname{relint}(C)\cap\operatorname{relint}(D)\neq\emptyset.$ (1)
It was shown in [14, theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains. We define the notions of separation and proper separation below.
Definition 4.3 (Separation of sets).
Two convex sets $C,D\subseteq\mathbb{R}^{n}$ are said to be separated if there exists some $v\in\mathbb{R}^{n}\setminus\{0\}$ such that
$\forall c\in C,d\in D,v\cdot c\leq v\cdot d.$ (2)
The concept of proper separation is slightly stronger.
Definition 4.4 (Proper separation of sets).
Sets $C,D\in\mathbb{R}^{n}$ are said to be properly separated when there exists some $v\in\mathbb{R}^{n}$ such that Equation 2 holds, and such that there is some pair $(c,d)\in C\times D$ for which the inequality in Equation 2 is strict.
Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in some affine hyperplane.
We will show that inside the joint facial subspace, we always have a “lack of separation” of the convex sets. Since this regularity condition is stronger than a “lack of proper separation”, it is stronger than Equation 1.
Theorem 4.2 (Qualification conditions always holds in the joint facial subspace).
For two convex sets $C,D\subseteq\mathbb{R}^{n}$, $C\cap T_{a}(C,D)$ and $D\cap T_{a}(C,D)$ are not separated.
The proof of Theorem 4.2 is the subject of Section 7.1. The next corollary follows immediately from Theorem 4.2.
Corollary 4.3 (rint qualification holds in joint facial subspace).
For convex sets $C,D\subseteq\mathbb{R}^{n}$,
$\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})\neq\emptyset.$
With Corollary 4.3 we can use standard results of convex analysis from [14] by first taking intersections with $T_{a}$. We do this in the proof of each of the results in Section 5. Also note that if we only had access to weaker results than those available in [14], results that would have the stronger assumption of “no separation of the sets” instead of “intersection of relative interiors is non-empty”, the theory in this paper would still hold, and indeed we would recover the stronger results with the weaker assumption, that the intersection of relative interiors is non-empty.
5 Main results
In this section we show five generalizations of important results of convex analysis: additivity of subdifferentials, normal cone of the intersection of two convex sets, attainment of the infimal convolution in the conjugate of a sum of two functions, the chain rule of the subdifferential, and the Fenchel-Rockafellar dual. All proofs are deferred to Section 7, because they require the machinery introduced in Section 6.
Proposition 5.1 (Recovery of the classical results).
In all results that follow, the classical analogue is recovered when qualification conditions hold. Indeed, given two convex functions $f,g$ such that their relative interiors intersect, we have $\operatorname{dom}(f),\operatorname{dom}(g)\subseteq T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$, and therefore $f+\iota_{T_{a}}=f$ and $g+\iota_{T_{a}}=g$.
For a proper convex function $f$, the singular subdifferential for some $x\in\mathbb{R}^{n}$ is defined as $\partial^{\infty}f(x):=N_{\operatorname{dom}(f)}(x)$ when $x\in\operatorname{dom}(f)$, and is empty when $x\notin\operatorname{dom}(f)$ (see, e.g. [10, 15]). Singular subdifferentials capture the notion of an infinite slope in a particular direction.
Theorem 5.1 (Qualification-free subdifferential additivity).
For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper convex, $x\in\operatorname{dom}(f)\cap\operatorname{dom}(g)$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x),$
$\partial^{\infty}(f+g)(x)=\partial^{\infty}(f+\iota_{T_{a}})(x)+\partial^{\infty}(g+\iota_{T_{a}})(x).$
Proof of Theorem 5.1.
By Corollary 4.3 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$,
$\operatorname{relint}\operatorname{dom}(f+\iota_{T_{a}})\cap\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})$ (3)
$=$ $\operatorname{relint}(\operatorname{dom}(f)\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (4)
Therefore, a qualification condition is satisfied for the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$, and the first part of the result follows by application of [14, Theorem 23.8].
The second part of the result follows from applying [14, Corollary 23.8.1] to the convex sets $\operatorname{dom}(f)\cap T_{a}$ and $\operatorname{dom}(g)\cap T_{a}$, by which
$N_{(\operatorname{dom}(f)\cap T_{a})\cap(\operatorname{dom}(g)\cap T_{a})}(x)$ $=N_{\operatorname{dom}(f)\cap T_{a}}(x)+N_{\operatorname{dom}(g)\cap T_{a}}(x)$
$=N_{\operatorname{dom}(f+\iota_{T_{a}})}(x)+N_{\operatorname{dom}(g+\iota_{T_{a}})}(x).$
And since $N_{(\operatorname{dom}(f)\cap T_{a})\cap(\operatorname{dom}(g)\cap T_{a})}(x)=N_{\operatorname{dom}(f)\cap\operatorname{dom}(g)}(x)=N_{\operatorname{dom}(f+g)}(x)$, the second part of the result follows. ∎
Theorem 5.2 (Normal cone of the intersection of convex sets).
For $C,D\subseteq\mathbb{R}^{n}$ convex, and $x\subseteq C\cap D$, for $T_{a}=T_{a}(C,D)$,
$N_{C\cap D}(x)=N_{C\cap T_{a}}(x)+N_{D\cap T_{a}}(x).$
Proof of Theorem 5.2.
With Corollary 4.3 we can use [14, Corollary 23.8.1], from which we find that
$N_{C\cap D}(x)=N_{(C\cap T_{a})\cap(D\cap T_{a})}(x)=N_{C\cap T_{a}}(x)+N_{D\cap T_{a}}(x).$
Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the infimal convolution is
$(f\,\square g)(z):=\inf_{x,y\in\mathbb{R}^{n}:x+y=z}f(x)+g(y).$
In the following results, we denote by $f^{*}$ the convex conjugate of $f$.
Theorem 5.3 (Attained infimal convolution in the conjugate of a sum).
For two proper convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$(f+g)^{*}=(f+\iota_{T_{a}})^{*}\square(g+\iota_{T_{a}})^{*},$
where the infimum in the infimal convolution is attained.
Proof of Theorem 5.3.
When $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, $T_{a}=\emptyset$ and the equality holds trivially. Otherwise, by Corollary 4.3, $\operatorname{relint}(\operatorname{dom}(f+\iota_{T_{a}}))\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))\neq\emptyset$, and the result follows from [14, Theorem 20.1]. ∎
Theorem 5.4 (Generalized subdifferential chain rule).
Let $g\in\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ is a matrix. Then for $T=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
$\partial(g)(Ax)=A^{T}\partial(g+\iota_{T})(Ax).$
Proof of Theorem 5.4.
Denote $T_{a}=T_{a}(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$. Note that $\forall x\in\mathbb{R}^{n}$,
$(g\circ A)(x)=([g+\iota_{T_{a}}]\circ A)(x),$
because $T_{a}\supseteq\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)$. So it suffices to calculate the subdifferential of the r.h.s. of the above equation.
If $\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)=\emptyset,$ then the result holds trivially because $g(Ax)\equiv\infty$, and $(g+\iota_{T_{a}})(Ax)\equiv\infty$ as well. Otherwise, by Corollary 4.3 $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$. Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Proposition 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$. Since $0\in\operatorname{\mathrm{range}}(A)$, this in fact implies that $T_{a}=T$. The statement then follows by applying [14, Theorem 23.9] to the composition $g+\iota_{T}]\circ A$. ∎
In the following result, for a matrix $A\in\mathbb{R}^{m\times n}$ and a set $S\subseteq\mathbb{R}^{m}$, we denote by $A^{-1}(S)$ the preimage of $S$ under the linear map associated with the matrix $A$, making no assumption on the invertibility of the matrix $A$.
Theorem 5.5 (Exact Fenchel-Rockafellar dual).
For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper, lower semi-continuous convex functions, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. Then
$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)=-\inf_{y\in\mathbb{R}^{m}}[(f+\iota_{A^{-1}(T_{a})})^{*}(A^{T}y)+(g+\iota_{T_{a}})^{*}(-y)].$
Proof of Theorem 5.5.
In this proof we use the notation of left multiplication of functions by a matrix
$(Af)(u):=\inf\{f(x)\mid Ax=u\}.$
Note that $Af+\iota_{T_{a}}\equiv A(f+\iota_{A^{-1}(T_{a})})$. Indeed, both functions have the domain $(A\operatorname{dom}(f))\cap T_{a}$, where $A\operatorname{dom}(f):=\{Ax\mid x\in\operatorname{dom}(f)\}$, and $\forall y\in A\operatorname{dom}(f)\cap T_{a}$,
$(Af+\iota_{T_{a}})(y)$ $=\inf\{f(x)\mid Ax=y\}$
$=\inf\{f(x)+\iota_{Ax\in T_{a}}\mid Ax=y\}$
$=A(f+\iota_{A^{-1}(T_{a})})(y).$
Then, employing the change of variable $u=Ax$,
$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)$ $=\inf_{u\in\mathbb{R}^{m}}(Af)(u)+g(u)$
$=\inf_{u\in\mathbb{R}^{m}}(Af+\iota_{T_{a}})(u)+(g+\iota_{T_{a}})(u)$
$=\inf_{u\in\mathbb{R}^{m}}(A(f+\iota_{A^{-1}{T_{a}}}))(u)+(g+\iota_{T_{a}})(u)$
$=\inf_{x\in\mathbb{R}^{n}}(f+\iota_{A^{-1}{T_{a}}})(x)+(g+\iota_{T_{a}})(Ax).$
Note that by Corollary 4.3, $\exists u\in\operatorname{relint}\operatorname{dom}(Af)\cap\operatorname{relint}\operatorname{dom}(g)$. Note that $\operatorname{dom}(Af)=A\operatorname{dom}(f)$, and by [14, Theorem 6.6], $\operatorname{relint}A\operatorname{dom}(f)=A\operatorname{relint}\operatorname{dom}(f)$. Therefore, $u\in A\operatorname{relint}\operatorname{dom}(f)\cap\operatorname{relint}\operatorname{dom}(g)$, i.e. there exists an $x\in\operatorname{relint}\operatorname{dom}(f)$ such that $Ax\in\operatorname{relint}\operatorname{dom}(g)$. With this, we satisfy a sufficient requirement of [14, Corollary 31.2.1], which guarantees strong duality and completes the proof. ∎
6 Nested normals and facial subspaces
Inspired by the lexicographic characterization of convex faces of [8], we develop a framework to describe the lattice of convex faces at a point $x\in C$. By using this theory, we will then be able to show the results of Section 4.
We begin with a characterization of faces of convex sets. In the next few results, we denote $X$ for a real vector space.
Theorem 6.1 (Facial subspace characterisation of convex faces).
Let $C\subseteq X$. A subset $F\subseteq C$ is a face of $C$ if, and only if the two following conditions hold.
  1. 1.
    $\operatorname{aff}(F)\cap C=F$,
  2. 2.
    $C\setminus F$ is convex.
Proof of Theorem 6.1.
Let $F$ be a face. We argue by contradiction that $\operatorname{aff}(F)\cap[C\setminus F]=\emptyset$; let us assume that there is some $z\in\operatorname{aff}(F)\cap[C\setminus F]$.
If $F$ is a singleton, then $\operatorname{aff}(F)=F$, and $\operatorname{aff}(F)\cap[C\setminus F]=\emptyset$. Otherwise, we may assume that $0\in F\setminus\{z\}$ without loss of generality, because the faces of a convex set are invariant under translation. With this assumption, $\operatorname{aff}(F)=\operatorname{\mathrm{span}}(F)=\operatorname{cone}(F-F)$. Then $z\in\operatorname{cone}(F-F)$, so there are $y_{1},y_{2}\in F$ such that $z=\alpha(y_{1}-y_{2})$ for some $\alpha\geq 0$. We can rule out the possibility that $z\in[y_{1},y_{2}]$, because by convexity of $F$ it would follow that $z\in F$. The remaining possibility is that $y_{1}\in[y_{2},z[$. Further, note that $y_{1}\neq y_{2}$, because otherwise $z=0$, which disagrees with our assumption that $z\in F\setminus\{0\}$. Therefore, $y_{1}\in]y_{2},z[$. The line segment $y_{2},z]$ has a point $y_{1}\in]y_{2},z[\cap F$. Since $y_{2},z\in C$, the fact that $F$ is a face implies that $z$ must then be in $F$, in contradiction with our definition of $z$. We have therefore shown the claim.
We next argue that $C\setminus F$ is convex. $\forall y,z\in C\setminus F$, any $x\in]y,z[$ is in $C$ by convexity of $C$. Also, $x$ is not in $F$ because $F$ is a face. So any $x\in]y,z[$ is also in $C\setminus F$, which means that $C\setminus F$ is convex.
We now show the converse: let $F\subseteq C$ be such that $C\setminus F$ is convex and $\operatorname{aff}(F)\cap C=F$, we show that $F$ is a face of $C$.
Take any $y,z\in C$, $x\in F\cap]y,z[$. If $y\in\operatorname{aff}(F)$ then $z\in\operatorname{aff}(F)$ as well, because $z=\alpha(x-y)+y$ for some $\alpha>0$. So either $z,y$ are both in $\operatorname{aff}(F)$, or neither are. But if neither are, then $z,y\in C\setminus\operatorname{aff}(F)$, and by convexity of $C\setminus\operatorname{aff}(F)$, $x\in C\setminus\operatorname{aff}(F)$ also, which is a contradiction with $x\in F$. Therefore, $z,y$ are both in $F$, which shows that $F$ is a face. ∎
This characterization, surprisingly, appears to be novel to the knowledge of the author. It allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set. We call such face-selecting subspaces affine facial subspaces.
Definition 6.1 (Affine facial subspace).
Let $C\subseteq X$ be convex. An affine subspace $L\subseteq V$ is an affine facial subspace of $C$ if $C\setminus L$ is convex.
The next result describes the relationship between affine facial subspaces and faces.
Corollary 6.1 (Facial subspace relation to faces).
Let $C\subseteq X$ be convex and $S\subseteq C$ non-empty.
  1. 1.
    For every affine facial subspace $L$ of $C$, the set $F:=C\cap L$ is a face of $C$.
  2. 2.
    Conversely, for every face $F$ of $C$, there exists an affine facial subspace $L$ such that $F=L\cap C$ (specifically, $L=\text{aff}(F)$).
Proof of Corollary 6.1.
For the first part of the result, we show that if $L$ is an affine facial subspace, then $F:=C\cap L$ satisfies the conditions of Theorem 6.1.
  1. 1.
    That $\operatorname{aff}(F)\cap C=F$: note that $\operatorname{aff}(F)\subseteq L$ because $F\subseteq L$, and so $\operatorname{aff}(F)\cap C\subseteq L\cap C=:F$. On the other hand, $L\cap C=(L\cap C)\cap C\subseteq\operatorname{aff}(L\cap C)\cap C$, and so $F\subseteq\operatorname{aff}(F)\cap C$.
  2. 2.
    That $C\setminus F$ is convex: note that
    $C\setminus F=C\setminus[C\cap L]=C\setminus L,$
    which is convex by definition of $L$ as an affine facial subspace.
We show the second part of the result. For $F$ a face of $C$, by Theorem 6.1 $F=C\cap\operatorname{aff}(F)$ and $C\setminus F$ is convex. But then $C\setminus\operatorname{aff}(F)=C\setminus(\operatorname{aff}(F)\cap C)=C\setminus F$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, which makes $\operatorname{aff}(F)$ an affine facial subspace. ∎
In this work, we focus on “local” affine facial subspaces; the subset of affine facial subspaces that contain a point $x\in C$. These subsets have some additional nice algebraic properties. We find it useful to generalize this idea of “locality” to affine facial subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all affine subspaces that contain $S$. To emphasize this connection, we use the word “at” to describe the containment of $S$: ”a facial subspace is at SS”. The reader should keep in mind the important special case of $S$ being a singleton $S=\{x\}$, for which the results are localized at a point.
Any affine facial subspace $L$ at $S$ has its affine offset determined by $S$. Recall that LLL-L is simply the affine subspace $L$ translated in such a way that it becomes a (non-affine) subspace. With this, it becomes clear that L=(LL)+SL=(L-L)+S. We can always deduce the correct affine offset from in $S$. Therefore, we find it convenient to describe the set of affine facial subspaces that contain $S$ through their translated, non-affine counterparts.
Definition 6.2 (Facial subspace at S).
Let $C\subseteq X$ be a convex set and $S\subseteq X$ a non-empty subset. A subspace $U\subseteq X$ is defined to be a facial subspace of $C$ at $S$ if:
  1. 1.
    $S-S\subseteq U$
  2. 2.
    $U+S$ is an affine facial subspace of $C$.
The set of all facial subspaces of $C$ at $S$ is denoted $\mathcal{F}(C;S)$.
When $S$ is the singleton $\{x\}$, the requirement that $S-S\subseteq U$ is trivially satisfied. When $S$ is not a singleton, this assumption guarantees that $\forall s_{1},s_{2}\in S,U+s_{1}=U+s_{2}$. This implies that $\forall s\in S,U+s=U+S$, and so the Minkowski sum $U+S$ is a simple affine offset of $U$, for which $S\subseteq U+S$. In other words, the affine facial subspace $U+S$ contains $S$.
Next, we introduce a generalization of the notion of a normal cone. A normal cone of a convex set $C\subseteq\mathbb{R}^{n}$ is defined as
$N_{C}(x):=\{z\in\mathbb{R}^{n}\mid\forall y\in C\langle y-x,z\rangle\leq 0\}.$
When $C$ is a cone and $x=0$, the normal cone $N_{C}(0)=:C^{\circ}$ is called the polar of $C$. The polar is a dual operation, in the sense that
${C^{\circ}}^{\circ}=\operatorname{cl}(C).$
We can see that a normal cone is a dual object which “describes” the original set only up to its closure.
We require a dual object which has sufficient descriptive power to go beyond the closure, encoding information about the boundary. To this end, we introduce the notion of a nested normal cone. Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
$N_{C}^{U}(x):=N_{C\cap U}(x)\cap U.$
We would also like for our normal cones to located at a set $S\subseteq C$, similarly to facial subspaces.
$N_{C}(S):=\{v\in\mathbb{R}^{n}\mid\forall s\in S,c\in C,\langle c-s,v\rangle\leq 0\}.$
The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$. Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [7], where normal cones are at faces of polytopes, instead at points. The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at any point in the relative interior of a convex set $S\subseteq C$ will match the normal cone of $S$.
Combining the two generalizations, we get the following.
Definition 6.3 (nested normals).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set, and $S$ a non-empty subset of $C$. Then for a subspace $U\subseteq\mathbb{R}^{n}$ such that $S-S\subseteq U$, and with $C^{\prime}:=C\cap(U+S)$,
$N_{C}^{U}(S):=\{v\in U\mid\forall c\in C^{\prime},s\in S,\langle c-s,v\rangle\leq 0\}$
is a nested normal cone of $C$ at $S$, and we label any $v\in N_{C}^{U}(S)$ to be a nested normal of $C$ at $S$ in $U$.
The idea of normal cone defined at a set has an implicit relation with the notion of relative interior.
Lemma 6.1 (nested normals in the relative interior are representative).
Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ a non-empty convex set. Let $U\subseteq\mathbb{R}^{n}$ be a subspace, and $x\in\operatorname{relint}(S)$. Then
$N_{C}^{U}(S)=N_{C}^{U}(x).$
Proof of Lemma 6.1.
That $N_{C}^{U}(S)\subseteq N_{C}^{U}(x)$ follows from simple inclusion $x\subseteq S$.
For the containment ($\supseteq$), WLOG let $U=\mathbb{R}^{n}$. We can make this assumption because the normal cones are invariant under the translation $U+S\to U$, and so we have an equivalent normal cone inside the subspace $U$. Since $U$ is a finite dimensional vector space, it is isomorphic to $\mathbb{R}^{n}$ for some $n\in\mathbb{N}$.
Let $v\in N_{C}(x)$. Then $\forall s\in S,\langle v,s\rangle\leq\langle v,x\rangle$ because $S\subseteq C$. We show that this inequality is in fact always an equality. Arguing by contradiction, assume that there is an $s\in S$ such that $\langle v,s\rangle<\langle v,x\rangle$. Then since $x\in\operatorname{relint}(S)$, there is a $\varepsilon>0$ such that $(\varepsilon B+x)\cap\operatorname{aff}(S)\subseteq S$ for $B$ the unit ball in $\mathbb{R}^{n}$. There is a $\lambda>0$ such that $s^{\prime}:=x+\lambda(s-x)\in\varepsilon B+x$. Then note that $s^{\prime}\in(\varepsilon B+x)\cap\operatorname{aff}(S)\subseteq S$, and
$\langle v,s^{\prime}\rangle=\langle v,x\rangle-\lambda\langle v,s-x\rangle.$
The second term is strictly positive, therefore $\langle v,s^{\prime}\rangle>\langle v,x\rangle$. But this is a contradiction with $v\in N_{C}(x)$. Therefore, we find that $\forall s\in S,$ $\langle v,s\rangle=\langle v,x\rangle$. Then $\forall c\in C,s\in S$ $\langle c-s,v\rangle=\langle c-x,v\rangle\leq 0$, and therefore $v\in N_{C}(S)$. ∎
The nested normals at a set are orthogonal with that set.
Lemma 6.2 (Matching sets means paired normals).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set, and $U\subseteq\mathbb{R}^{n}$ a subspace. Then for any non-empty $S\subseteq C$ such that $S-S\subseteq U$,
$N_{C}^{U}(S)\subseteq(S-S)^{\perp}\cap U.$
Further, when $S=C\cap(U+S)$, the sets are equal.
Proof of Lemma 6.2.
Let $y\in N_{C}^{U}(S)$. By Definition 6.3, $\forall z,x\in S,$ since $x\in S$ and $z\in C$, $\langle x-z,y\rangle\leq 0,$ and since $x\in C$ and $z\in S$ $\langle x-z,y\rangle\geq 0$. Then $y\perp S-S$, which shows the inclusion.
Next we show the containment ($\supseteq$), in the case that $S=C\cap(U+S)$. Let $y\in(S-S)^{\perp}\cap U$. For any $x\in S,z\in C\cap(U+S)$ we need to show that $\langle z-x,y\rangle\leq 0$. Note that $z\in S$, so $x-z\in S-S$, and by definition of $y$, $\langle z-x,y\rangle=0$, which proves the statement. ∎
As previously discussed, one motivating property of the nested normals is that they can encode information about the boundary of convex sets. We make this idea concrete.
The tangent cone of a convex set $C$ at $x\in C$ is the smallest closed cone $T$ such that $C\subseteq x+T$. We can think of this cone as a local description of $C$ around $x$. The normal cone is the polar of hthe tangent cone, and therefore “describes” the convex set $C$ around $x$ to the same extent as the tangent cone. In contrast, nested normals in facial subspaces of $C$ at $x$ describe the cone of feasible directions.
Definition 6.4 (Cone of feasible directions).
For a convex set $C\in\mathbb{R}^{n}$ with $x\in C$, we define the cone of feasible directions of $C$ at $x$ to be $D_{C}(x):=\{v\in\mathbb{R}^{n}\mid\exists\lambda>0:\lambda v\in C\}$.
The closure of the cone of feasible directions is the tangent cone. The difference is that the cone of feasible directions captures additional information about the boundary of $C$ around $x$. The following result shows that any convex cone (including the cone of feasible directions at any point) can be recovered from the nested normals in facial subspaces.
Theorem 6.2 (Local polar without closure).
Given any convex cone $T$ with $0\in T$,
$T=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\subseteq N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}.$
Proof of Theorem 6.2.
Let $T^{\prime}:=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\subseteq N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}$. By the definition of conditional normals, $T\subseteq T^{\prime}$. For the containment ($\supseteq$), we argue by contradiction: we assume that there is some $x\in T^{\prime}\setminus T$. If $x\in\mathrm{int}(T^{c})$, then $x$ is strictly separated from $T$ by a supporting hyperplane of $T$ (e.g., see [14, Theorem 11.3]). Any such hyperplane induces a nested normal of the form $v\in N_{T}^{\mathbb{R}^{n}}(0)$, for which $v^{\perp}$ is the hyperplane, and $\langle v,x\rangle>0$. But then $x\notin T^{\prime}$, and so $x\notin\mathrm{int}(T^{c})$. Since we assumed that $x\notin T$, the only possibility is that $x\in\partial T$. Then $x$ is contained in a supporting hyperplane, so there is some nested normal $v\in N_{T}^{\mathbb{R}^{n}}(0)$ such that $\langle v,x\rangle=0$. Let $T_{1}=T\cap v^{\perp}$, and narrow the ambient space to the subspace $T_{1}$. Inside $T_{1}$, we repeat the same argument, and keep narrowing the ambient space in this manner. This process must eventually reduce the ambient space to $\operatorname{\mathrm{span}}(x)$, because $\mathbb{R}^{n}$ is finite-dimensional. In this one-dimensional space, with the narrowed cone $T_{\ell}=T\cap\operatorname{\mathrm{span}}(x)$, note that $T_{\ell}$ is a closed cone. Therefore, either $x\in T_{\ell}$ or $x\in\mathrm{int}(T_{\ell}^{c})$, both of which yields a contradiction, from which the result follows. ∎
Another important property of nested normals is that they describe the containment structure of facial subspaces, which is closely related to the lattice structure of the faces. We make this precise in the next two results.
Lemma 6.3 (Nested normals induce facial subspaces).
Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, $U\subseteq\mathbb{R}^{n}$ a subspace, and $v\in N_{C}^{U}(S)$. Let $V\in\mathcal{F}(C;S)$ be such that $V\subseteq U$. Then
$V\cap v^{\perp}\in\mathcal{F}(C;S).$
Lemma 6.4 (Nested normals induce facial subspaces).
First, note that if $v=0$, then $V\cap v^{\perp}=V$, and there result holds. We now consider the case where $v\neq 0$.
By Lemma 6.2, $v\perp S-S$, and therefore $S-S\subseteq V\cap v^{\perp}$.
We make two observations.
First, since $v\in N_{C}^{U}(S)$ and $V\subseteq U$, $\forall y\in C\cap(V+S)$, $\forall s\in S,$ we have that $\langle y,v\rangle\leq\langle s,v\rangle$. Second, this inequality is tight precisely for those values of $y$ that are also contained in $v^{\perp}+S$, i.e.,
$\{y\in C\cap(V+S)\mid\forall s\in S,\langle y,v\rangle=\langle s,v\rangle\}$ (5)
$=$ $((V+S)\cap C)\cap(v^{\perp}+S)=(V\cap v^{\perp}+S)\cap C.$ (6)
These two facts imply that
$(V\cap v^{\perp}+S)\cap C=\operatorname*{\arg\max}_{y\in C\cap(V+S)}\langle y,v\rangle.$
The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, i.e., $V\cap v^{\perp}\in\mathcal{F}(C;S)$.
We next show that all facial subspaces are “reached” by nested normals in facial subspaces.
Lemma 6.5 (Completeness of conditional normal proof composition).
Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, and $V,U\in\mathcal{F}(C;S)$ such that $U\subseteq V$. Then there is a sequence $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)$ with $k=\dim(V)-\dim(U)$ such that $V_{0}=V$, $V_{k}=U$, and for each $i\in\{1,\ldots,k\}$ there is a $v_{i}\in N_{C}^{V_{i-1}}(S)\setminus\{0\}$ such that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$.
This result is similar to the lexicographic characterization of the faces of convex sets in [8, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face. Let us state a technical lemma that we will need for the proof.
Lemma 6.6 (Affine slices preserve facial subspaces).
Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ non-empty. Let $U\in\mathcal{F}(C;S)$, and $V$ any subspace such that $S-S\subseteq V$. Then $U\cap V\in\mathcal{F}(C\cap(V+S);S)$.
Proof of Lemma 6.6.
Since $C\setminus(U+S)$ is convex, $(V+S)\cap(C\setminus(U+S))$ is also convex. We can write this set as $(V+S)\cap C]\setminus[(V+S)\cap(U+S)]$, which is the same as $((V+S)\cap C)\setminus((V\cap U)+S)$, because $S-S\in V\cap U$. Since this set is convex, and $S-S\subseteq V\cap U$, we find that $V\cap U\in\mathcal{F}((V+S)\cap C;S)$. ∎
Proof of Lemma 6.5.
It suffices to show the following statement: given any $U,V\in\mathcal{F}(C;S)$ such that $V\supsetneq U$, there exists a $v\in N_{C}^{V}(S)\setminus\{0\}$ such that $v\perp U$. If this statement is shown to hold, applying it recursively generates the sequence of facial subspaces $V_{1},\ldots,V_{k}$ in a finite number of steps. Note that the dimension of the facial subspaces decreases by one at each iteration while always containing $U$, which implies that $U\subseteq V_{k}$ with $\dim(U)=\dim(V_{k})$, and therefore $U=V_{k}.$
We now show the above statement. Let $C^{\prime}=C\cap(V+S)$, $\bar{C}=\operatorname{\Pi}_{U^{\perp}}C^{\prime}$, and $\bar{U}=\operatorname{\Pi}_{U^{\perp}}(U+S)$. Note that $\bar{C}$ is a convex set and $\bar{U}$ a singleton $\{x\}\subseteq\bar{C}$. We argue that $\bar{U}$ is a face of $\bar{C}$. Arguing by contradiction, we hypothesize the $\bar{U}$ is not a face of $\bar{C}$, in which case there are $\bar{y},\bar{z}\in\bar{C}\setminus\bar{U}$ such that $x=\lambda\bar{y}+(1-\lambda)\bar{z}$ for $\lambda\in(0,1)$. This implies that there are elements $y,z\in C^{\prime}$ such that $\bar{y}=\operatorname{\Pi}_{U^{\perp}}y$ and $\bar{z}=\operatorname{\Pi}_{U^{\perp}}z$, with $w=\lambda y+(1-\lambda)z$. Note that the line segment $[y,z]$ has endpoints $y,z\in C^{\prime}\setminus(U+S)$ with a point $w\in]y,z[$ that is also contained in $C^{\prime}\cap(U+S)$. This implies that $C^{\prime}\cap(U+S)$ is not a face of $C^{\prime}$, which, by Corollary 6.1, means that $U\notin\mathcal{F}(C^{\prime};S)$. This is a contradiction because in fact, $U\in\mathcal{F}(C^{\prime};S)$, as we now show. We know that $U\in\mathcal{F}(C;S)$, so by Lemma 6.6, $U=U\cap V\in\mathcal{F}(C\cap(V+S);S)$, i.e., $U\in\mathcal{F}(C^{\prime};S)$. Therefore, we indeed have a contradiction, from which we find that $\bar{U}$ is a face of $\bar{C}$. Note that $\operatorname{\Pi}_{U^{\perp}}(V+S)\subseteq V+S$. This becomes apparent from the following equality: $V+S=(V+S)+U=\operatorname{\Pi}_{U^{\perp}}(V+S)+U,$ where the first equality holds because $U\subseteq V$ implies that $V=V+U$.
Then since $U+S,C\cap(V+S)\subseteq V+S$, we find that $\bar{U},\bar{C}\subseteq(V+S)\cap U^{\perp}$. Since $\bar{U}$ is a face of $\bar{C}$, there exists a supporting hyperplane $H$ of $\bar{C}$, where the hyperplane is relative to the affine subspace $(V+S)\cap U^{\perp}$, such that $\bar{U}\subseteq H$. Indeed, if there is no such hyperplane $H$, $\bar{U}$ would then be a singleton in the interior of $\bar{C}$ (where the interior is relative to $(V+S)\cap U^{\perp}$, i.e., with the induced topology in $(V+S)\cap U^{\perp}$).
Take $\bar{v}\in V\cap U^{\perp}$ to be the unit normal to $H$ that is also in $N_{\bar{C}}^{V\cap U^{\perp}}(x)$. Then note that $\bar{v}\in N_{C}^{V}(S)\cap U^{\perp}$, which is the vector we had to find. ∎
Together, Lemma 6.3 and Lemma 6.5 enables a powerful tool that is central to our proofs: we can perform structural induction on the tree of facial subspaces, using only nested normals. Indeed, the set of facial subspaces forms a tree, where the root is $\mathbb{R}^{n}$ (note that it is always a facial subpace, and that it contains any other facial subspace), and every edge of the tree is associated with a nested normal. The induction statement is the following.
Proposition 6.1 (Induction by nested normal).
Let $C$ be a convex set, $S\subseteq C$ a non-empty subset, and a statement $A:\mathcal{F}(C;S)\to\{\text{true},\text{false}\}$. If we have show that:
  1. 1.
    $A(\mathbb{R}^{n}).$
  2. 2.
    $\forall U\in\mathcal{F}(C;S)$, $v\in N^{U}_{C}(S)$, $A(U)\implies A(U\cap v^{\perp})$. Then it follows that
    $\forall V\in\mathcal{F}(C;S),\,A(W).$
6.1 Minimal facial subspace
Recall that the smallest face of a convex set $C$ containing some $x\in C$ is $F_{C}(x)$, the face of $C$ generated by $x$. It is the only face with $x$ in its relative interior. In this section we develop an analogous notion for facial subspaces, and characterize this facial subspace by using nested normals.
Definition 6.5 (The generated facial subspace).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, we define the generated facial subspace $H_{C}(S)$ of $C$, generated by $S\subseteq\mathbb{R}^{n}$, to be the facial subspace that is contained in every other facial subspace in $\mathcal{F}(C;S)$.
We show existence and uniqueness for the above definition, starting with a lemma.
Lemma 6.7 (Facial subspaces are closed under intersections).
For $C\subseteq\mathbb{R}^{n}$ convex and $S\subseteq C$ non-empty, let $U,V\in\mathcal{F}(C;S)$. Then $U\cap V\in\mathcal{F}(C;S)$.
Proof of Lemma 6.7.
The subspace $U\cap V$ is such that
$(U\cap V+S)\cap C=((U+S)\cap C)\cap((V+S)\cap C).$
The r.h.s. is the intersection of two convex faces of $C$. Their intersection is then also a face (this is immediate from Definition 3.1). By Theorem 6.1, $C\setminus[(U\cap V+S)\cap C]$ is convex. The convexity of this set, which we can write as $C\setminus((U\cap V)+S)$, implies that $U\cap V\in\mathcal{F}(C;S)$. ∎
Theorem 6.3 (Existence and uniqueness of the generated facial subspace).
The generated facial subspace of Definition 6.5 exists and is unique.
Proof of Theorem 6.3.
We argue that there exists a facial subspace which contains no other. Arguing for the existence by contradiction, if every facial subspace of $C$ at $S$ properly contains some other facial subspace of $C$ at $S$, starting at any facial subspace, we can find an infinite chain of facial subspaces with dimension decreasing by one at each step. This is not possible because the dimension is finite. Therefore this chain of subspaces must terminate at some subspace which contains no other.
We now argue that there is only one subspace that contains no other. Let $V_{1},V_{2}$ be two distinct minimal facial subspaces of $C$ at $S$. Then by Lemma 6.7, $V_{1}\cap V_{2}$ is a facial subspace of $C$ at $S$ properly contained in both $V_{1}$ and $V_{2}$, which contradicts $V_{1}$ and $V_{2}$ being minimal.
We have shown the existence and uniqueness of a facial subspace that contains no other. From any facial subspace, we can start a proper containment chain as in our existence argument. The subspace at the end of this chain is the subspace which contains no other, which concludes the proof. ∎
As might be expected, the generated facial subspace is closely related to the generated face.
Lemma 6.8 (Relation of the generated face to the generated facial subspace).
With $C\subseteq\mathbb{R}^{n}$ convex and $S\subseteq C$ non-empty,
  1. 1.
    $F_{C}(S)=C\cap(H_{C}(S)+S).$
  2. 2.
    $H_{C}(S)=\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S)).$
Proof of Lemma 6.8.
By Theorem 6.1, $H_{C}(S)$ being the smallest facial subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own facial subspace (by Corollary 6.1), which would have to be properly contained in $H_{C}(S)$.
For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a facial subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is an affine facial subspace, as implied by Theorem 6.1. Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other facial subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$. ∎
We can now reformulate Definition 4.1 into an equivalent definition.
Definition 6.6 (Equivalent definition of the joint facial subspace).
The joint facial subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ from Definition 4.1 is
$T(C,D):=H_{C-D}(0).$
The next result characterizes generated facial subspaces in terms of nested normals.
Lemma 6.9 (The generated facial subspace is the only facial subspace with no nested normals).
Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated facial subspace $H_{C}(S)$ is the only facial subspace of $C$ at $S$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
Proof of Lemma 6.9.
The generated facial subspace does not contain any non-trivial nested normals, because any non-trivial nested normal $v\in N_{C}^{H(C)}(S)\setminus\{0\}$ induces the facial subspace $H_{C}(S)\cap v^{\perp}\subsetneq H_{C}(S)$ by Lemma 6.3.
For any facial subspace $V\in\mathcal{F}(C;S)$ such that $N_{C}^{V}(S)=\{0\}$, we argue that $V=H_{C}(S)$. Note that $H_{C}(S)\subseteq V$ by Definition 6.5. The inclusion cannot be proper, because otherwise, by Lemma 6.5, there would be a nontrivial nested normal $v\in N_{C}^{V}(S)\setminus\{0\}$. ∎
6.2 Nested normals of a difference of sets
In this section, we show the following important result.
Lemma 6.10 (Characterization of the conditional normals of a difference of sets).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex with $S\subseteq C\cap D$ non-empty, and $U\in\mathcal{F}(C-D;0)$. Then $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and
$N^{U}_{C-D}(0)=N^{U}_{C}(S)\cap(-N^{U}_{D}(S)).$
To show Lemma 6.10, we require the following lemma.
Lemma 6.11 (Normal cone of the difference of convex sets).
Given convex sets $C,D\subseteq\mathbb{R}^{n}$ with $S\subseteq C\cap D$ non-empty,
$N_{C-D}(0)=N_{C}(S)\cap(-N_{D}(S)).$
Proof of Lemma 6.11.
We start with the inclusion $(\subseteq)$. Any $v\in N_{C-D}(0)$ has the property that $\forall c\in C,d\in D$, $\langle v,(c-d)-0\rangle\leq 0$. Then $\forall c\in C,s\in S$, $\langle v,c-s\rangle\leq 0$ because $s\in D$, and therefore $v\in N_{C}(S)$. By a similar argument, $v\in N_{D}(S)$, which shows the inclusion.
For the containment ($\supseteq$), fix some $s\in S$. let $v\in N_{C}(S)\cap(-N_{D}(S))$. Then $\forall c\in C,d\in D$
$\langle v,c-d\rangle=\langle v,c-s\rangle+\langle v,s-d\rangle\leq 0,$
and so $v\in N_{C-D}(0)$. ∎
Proof of Lemma 6.10.
We argue by structured induction over the facial subspaces of CDC-D, as described in Proposition 6.1. Fixing some $S\subseteq C\cap D$, the statement $A:\mathcal{F}(C-D;0)\to\{\text{true},\text{false}\}$ that we show is the following: $A(U)$ is true if, and only if, all of the following hold.
  1. 1.
    $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S).$
  2. 2.
    $N^{U}_{C-D}(0)=N^{U}_{C}(S)\cap(-N^{U}_{D}(S)).$
  3. 3.
    $C\cap(U+S)-D\cap(U+S)=(C-D)\cap U.$ The additional point 3) has a supportive role in this proof.
The base case, where $U=\mathbb{R}^{n}$, holds because $\mathbb{R}^{n}$ is always a facial subspace. Point 2) is given by Lemma 6.11, and point 3) holds trivially.
For the inductive step, let $V\in\mathcal{F}(C-D;0)$ be such that $A(V)$ is true. Let $v\in N_{C-D}^{V}(0)\setminus\{0\}$, and $U=V\cap v^{\perp}$. In the rest of the proof, we show that $A(U)$ is true. Note that by point 2) of $A(V)$, $v\in N_{C}^{V}(S)\cap(-N_{D}^{V}(S))$, which by Lemma 6.3 implies that $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and this shows point 1) of $A(U)$.
We next show that for $C^{\prime}=C\cap(V+S)$, $D^{\prime}=D\cap(V+S)$,
$(C^{\prime}-D^{\prime})\cap U=(C^{\prime}\cap(U+S))-(D^{\prime}\cap(U+S)).$ (8)
We start with the inclusion ($\subseteq$): Take $z\in C^{\prime}$, $y\in D^{\prime}$ with $z-y\in U$. Then since $U\subseteq v^{\perp}$, $\langle z-y,v\rangle=0$, and therefore $\forall s\in S$, $\langle z-s,v\rangle-\langle y-s,v\rangle=0$. That $\langle z-s,v\rangle\leq 0$ follows from $v\in N_{C}^{V}(S)$ since $z\in C^{\prime}\subseteq C$ and $z-s\in V$, and similarly, $\langle y-s,v\rangle\geq 0$ because $v\in N_{D}^{V}(S)$. But then the only way that $\langle z-s,v\rangle-\langle y-s,v\rangle=0$ is if $\langle z-s,v\rangle=0$ and $\langle y-s,v\rangle=0$ individually. It follows that $z,y\in U+S$, and so the inclusion holds. The containment ($\supseteq$) is trivial, so we have shown Equation 8.
By point 3) of $A(V)$, $C^{\prime}-D^{\prime}=(C-D)\cap V$. Also, since $U\subseteq V$, $C^{\prime}\cap(U+S)$ = $C\cap(U+S)$, and similarly for $D$. Applying these identities on both sides of Equation 8, we find that
$(C-D)\cap U=C\cap(U+S)-D\cap(U+S).$
This shows that point 3) of $A(U)$ is true.
To show point 2) of $A(U)$, we compute
$N_{C}^{U}(S)\cap(-N_{D}^{U}(S))$ $=[N_{C\cap(U+S)}(S)\cap(-N_{D\cap(U+S)}(S))]\cap U$ (9)
$=N_{C\cap(U+S)-D\cap(U+S)}(0)\cap U$ (10)
$=N_{(C-D)\cap U}(0)\cap U$ (11)
$=N_{C-D}^{U}(0),$ (12)
where Equation 9 holds by definition of nested normals, Equation 10 follows from Lemma 6.11 applied to the sets $C\cap(U+S)$ and $D\cap(U+S)$, and Equation 11 follows from point 3) of $A(U)$. This shows that point 2) of $A(U)$ holds for $U$, and the proof is complete. ∎
We now have the tools to prove Corollary 4.1.
Proof of Corollary 4.1.
Considering Definition 4.1 in the light of Corollary 6.1, $T(C,D)=H_{C-D}(0)$. So $T\in\mathcal{F}_{C-D}(0)$, which by Lemma 6.10 implies that, taking $S=C\cap D$, $T\in\mathcal{F}_{C}(C\cap D)\cap\mathcal{F}_{D}(C\cap D)$. By the definition of facial subspaces (Definition 6.2), $C\cap D-C\cap D\subseteq T$. ∎
7 Proofs
We first show a smaller result, after which each subsection is dedicated to proving one result from Section 4.
Proof of Proposition 5.1.
First we show that for $C\subseteq\mathbb{R}^{n}$ convex, and $x\in\operatorname{relint}(C)$, $F_{C}(x)=C$. Note that
$N_{C}^{\operatorname{\mathrm{span}}(C-C)}(x)=N_{C}^{\operatorname{\mathrm{span}}(C-C)}(C)=(C-C)^{\perp}\cap\operatorname{\mathrm{span}}(C-C)=\{0\},$
where the first equality follows by Lemma 6.1 and the second by Lemma 6.2. Then by Lemma 6.9, $H_{C}(x)=\operatorname{\mathrm{span}}(C-C)$, and by Lemma 6.8, $F_{C}(x)=(\operatorname{\mathrm{span}}(C-C)+x)\cap C=C$.
From Proposition 4.1, for $x\in\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))$ and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
$T_{a}\cap\operatorname{dom}(f)=F_{\operatorname{dom}(f)}(\operatorname{dom}(f)\cap\operatorname{dom}(g))\supseteq F_{\operatorname{dom}(f)}(x)=\operatorname{dom}(f).$
From this the results follows. ∎
7.1 Regularity in the joint facial subspace
We now show the main property of the joint facial subspace, Theorem 4.2. We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [10, Theorem 2.19].
Lemma 7.1 (Equivalence of collisions and separation).
For convex sets $C,D\subseteq\mathbb{R}^{n}$, and $S\subseteq C\cap D$ non-empty, the following statement holds.
$N_{C}(S)\cap(-N_{D}(S))\neq\{0\}\iff C\text{ is separated from }D.$ (13)
Proof of Lemma 7.1.
We start with $(\implies)$. There exists a $v\in[N_{C}(S)\cap(-N_{D}(S))]\setminus\{0\}$, and from Lemma 6.2 $v\perp S-S$, whence the hyperplane $v^{\perp}+S$ separates $C$ from $D$.
[…]
  1. 1.
    $\forall s\in S,\langle v,s\rangle=\beta$.
  2. 2.
    $\forall c\in C$, $\langle y,v\rangle\geq\beta$.
  3. 3.
    $\forall d\in D$, $\langle d,v\rangle\leq\beta$. Then $\forall y\in C,s\in S$, $\langle v,y-s\rangle\geq 0,$ and so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, so $v\in N_{C}(S)\cap(-N_{D}(S))$, which shows that $N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$.
Proof of Theorem 4.2.
Recall that from Definition 6.6, $T=H_{C-D}(0)$. Then by Lemma 6.9, $N_{C-D}^{T}(C\cap D)=\{0\}$. Applying Lemma 6.10, we find that $N^{T}_{C}(C\cap D)\cap(-N^{T}_{D}(C\cap D))=\{0\}$.
By applying Lemma 7.1 in the space $T_{a}$, the result follows. Note that Lemma 7.1 is stated in $\mathbb{R}^{n}$, yet we can apply it in $T_{a}$. Indeed, $T$ is a finite-dimensional subspace, and thereby isomorphic to $\mathbb{R}^{n}$, and Lemma 7.1 is invariant when translating $T$ to $T_{a}$. ∎
7.2 First characterization
In this section we prove Proposition 4.1 and Theorem 4.1.
Lemma 7.2 (nested constraint subspaces for nested normals).
Let $C\subseteq\mathbb{R}^{n}$ be a convex set, $S\subseteq C$ be non-empty, and $U,V\subseteq\mathbb{R}^{n}$ be subspaces such that $S-S\subseteq U\subseteq V$ and $C\cap(V+S)\subseteq U+S$. Then
$N_{C}^{U}(S)+(U^{\perp}\cap V)=N_{C\cap(U+S)}^{V}(S).$
Proof of Lemma 7.2.
We start with the containment ($\supseteq$). Let $x\in N_{C}^{V}(S).$ Then $\forall s\in S,y\in C\cap(V+S)$, note that $y-s\in U$, and so
$\langle\operatorname{\Pi}_{U}x,y-s\rangle=\langle x,y-s\rangle\leq 0.$
This shows that $\operatorname{\Pi}_{U}x\in N_{C}^{U}(S)$. Now $x=\operatorname{\Pi}_{U}x+\operatorname{\Pi}_{U^{\perp}}x$, and $x,\operatorname{\Pi}_{U}x\in V\implies\operatorname{\Pi}_{U^{\perp}}x\in V$, therefore $x\in N_{C}^{U}(S)+(U^{\perp}\cap V)$.
For the inclusion ($\subseteq$), we consider any $q\in N^{U}_{C}(S)$ and $z\in U^{\perp}\cap V$. Let $y\in C\cap(U+S),s\in S$, and note that $\langle z,y-s\rangle=0$ because $y-s\in U$. Then
$\langle q+z,y-s\rangle=\langle q,y-s\rangle\leq 0.$
Therefore $q+z\in N_{C\cap(U+S)}^{V}(S)$, which shows the inclusion and completes the proof. ∎
Lemma 7.3 (Conical projections yield facial subspaces).
Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, and $U\in\mathcal{F}(C;S)$. Then for any $K\subseteq N^{U}_{C}(S)$,
$U\cap K^{\perp}\in\mathcal{F}(C;S).$
Proof of Lemma 7.3.
Since $K$ is finite dimensional, there is a finite number of vectors $v_{1},\ldots,v_{m}\in K$ such that $\operatorname{\mathrm{span}}(K)=\operatorname{\mathrm{span}}(v_{1},\ldots,v_{m})$. Then since $v_{1},\ldots,v_{m}\in N_{C}^{U}(S)$, by repeated application of Lemma 6.3, $U\cap v_{1}^{\perp}\ldots\cap v_{m}^{\perp}$ is a facial subspace, which is the same set as $U\cap K^{\perp}$. ∎
Lemma 7.4 (Normal to the generated facial subspace of one convex set).
For convex sets $C,D\subseteq\mathbb{R}^{n}$, $T=T(C,D)$, $T_{a}=T+C\cap D$, we have
$H_{C}(C\cap D)=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}.$
Proof of Lemma 7.4.
Let $U=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}$. we first show that $U\in\mathcal{F}(C\cap D;C)$. Note that $T\in\mathcal{F}(C\cap D;C)$ by Lemma 6.10. Then we can use Lemma 7.3 because $N_{C}^{T}(C\cap T_{a})\subseteq N_{C}^{T}(C\cap D)$, which shows that $U\in\mathcal{F}(C\cap D;C)$ as well.
In the remainder of the proof we show that the assumption of Lemma 6.9 holds for $U$, namely that $N_{C}^{U}(C\cap D)=\{0\}$.
We first show that $N_{C}^{U}(C\cap T_{a})=\{0\}$. By Lemma 6.2, $N_{C}^{T}(C\cap T_{a})=(C\cap T_{a}-C\cap T_{a})^{\perp}\cap T$. Also, note that by the identity in Lemma 7.2, we find that $N_{C}^{U}(C\cap T_{a})=N_{C}^{T}(C\cap T_{a})\cap U$. Combining these two facts,
$N_{C}^{U}(C\cap T_{a})=N_{C}^{T}(C\cap T_{a})\cap U=(C\cap T_{a}-C\cap T_{a})^{\perp}\cap U.$
But note that $C\cap T_{a}-C\cap T_{a}\subseteq U$, therefore it fo $N_{C}^{U}(C\cap T_{a})=\{0\}$. That $C\cap T_{a}-C\cap T_{a}\subseteq U$ holds because $C\cap T_{a}-C\cap T_{a}\subseteq T$ and $(C\cap T_{a}-C\cap T_{a})\perp N_{C}^{T}(C\cap T_{a})$ by Lemma 6.2.
It only remains to show that $N_{C}^{U}(C\cap D)=N_{C}^{U}(C\cap T_{a})$. By Corollary 4.3, we may pick some $x\in\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})$. Then by Lemma 6.1,
$N_{C}^{U}(C\cap T_{a})=N_{C}^{U}(x)\supseteq N_{C}^{U}(C\cap D).$
Then $N_{C}^{U}(C\cap D)=\{0\}$, and so the result follows from Lemma 6.9. ∎
Proof of Proposition 4.1.
From Lemma 7.4, $H_{C}(C\cap D)=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}$. Since $(C\cap T_{a}-C\cap T_{a})\perp N_{C}^{T}(C\cap T_{a})$ by Lemma 6.2, $(C\cap T_{a}-C\cap T_{a})\subseteq H_{C}(C\cap D)$, i.e., $H_{C}(C\cap D)+C\cap T_{a}\supseteq C\cap T_{a}$. But $H_{C}(C\cap D)+C\cap T_{a}=H_{C}(C\cap D)+C\cap D$, therefore
$H_{C}(C\cap D)+C\cap D\supseteq C\cap T_{a}.$ (14)
Additionally, since we know that $T\in\mathcal{F}(C\cap D;C)$ from Lemma 6.10, by definition of the generated facial subspace we know that $H_{C}(C\cap D)\subseteq T$. Therefore,
$H_{C}(C\cap D)+C\cap D\subseteq T_{a}.$ (15)
From Equation 14 and Equation 15 it follows that $(H_{C}(C\cap D)+C\cap D)\cap C=C\cap T_{a}$. Then the statement follows from Lemma 6.8. ∎
Proof of Theorem 4.1.
From Proposition 4.1, we have that $T\in\mathcal{F}(C;C\cap D)$ is a facial subspace of $C$ that contains $C\cap D$, and so by Definition 6.5, $H_{C}(C\cap D)\subseteq T$. Therefore, for $^{\prime}:=H_{C}(C\cap D)+H_{D}(C\cap D), TT$$^{\prime}\subseteq T$.
For the other containment, we argue by contradiction: let $T^{\prime}\subsetneq T$. Then there is a unit vector $v\in T$ with $v\perp T^{\prime}$. Then $v\perp H_{C}(C\cap D)$, which by Lemma 7.4, implies that $v\in N_{C}^{T}(C\cap T)$. By Lemma 6.2, $v\in N_{C}^{T}(C\cap T)$ as well. And since $N_{C}^{T}(C\cap T)\subseteq N_{C}^{T}(C\cap D)$, we find that $v,-v\in N_{C}^{T}(C\cap D)$. By a similar argument, $v,-v\in N_{D}^{T}(C\cap D)$. But then $v\in N_{C}^{T}(C\cap D)\cap(-N_{D}^{T}(C\cap D))$, which by Lemma 6.10 means that $v\in N_{C-D}^{T}(0)$. But by Lemma 6.9, $T$ cannot contain such a nested normal because by Definition 6.6, $T=H_{C-D}(0)$. Therefore, by contradiction $T=T^{\prime}$. ∎
7.3 Second characterization
We can now state the second characterization in full.
Theorem 7.1 (Iterative bilateral facial reduction).
Let $C,D\subseteq\mathbb{R}^{n}$ be convex sets and $S\subseteq C\cap D$ non-empty. Beginning with $T_{0}=\mathbb{R}^{n}$, let
$T_{i+1}=T_{i}\cap(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])^{\perp}.$
Let $\ell$ be the first index for which $T_{\ell+1}=T_{\ell}$. Then:
  1. 1.
    $T_{\ell}=T(C,D)$.
  2. 2.
    $\ell\leq n$.
Proof of Theorem 7.1.
First, $(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])=N^{T_{i}}_{C-D}(S)$ by Lemma 6.10. Applying Lemma 7.3 at each step, we find that $\forall i\in\{0,\ldots,\ell\},\,T_{i}\in\mathcal{F}(C-D;S)$. Finally, note that whenever $N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)]\neq\{0\}$, $\dim(T_{i+1})\leq\dim(T_{i})-1$. Therefore, the dimension may be reduced at most $n$ times, which implies that $\ell\leq n$. On the other hand, the fact that $T_{\ell}=T_{\ell+1}$ implies that $N^{T_{\ell}}_{C-D}(S)=\{0\}$. Then from Lemma 6.9 it follows that $T_{\ell}=H_{C-D}(S)$, which is $T(C,D)$ from Definition 6.6. ∎
7.4 Aknowledgements
The author would like to thank Sharvaj Kubal, prof. Yaniv Plan, and prof. Ozgur Yilmaz for many fruitful discussions. A special thanks to prof. Michael Friedlander for pointing out that a generalization of the Fenchel-Rockafellar dual would be a promising avenue of research. This paper was produced with the Latex Exporter plugin in Obsidian.
range
pages36 rangepages12 rangepages13 rangepages12 rangepages25 rangepages17 rangepages34

### v3

@@
-Bilateral facial reduction: qualification-free subdifferential calculus and exact duality in convex analysis
+Qualification-free convex analysis via the joint supporting subspace
 Abstract
-This paper introduces a geometric framework to extend convex analysis results beyond standard qualification conditions such as intersecting relative interiors of domains.
-We define the joint facial subspace $T$ as the span of the face of CDC-D generated by 0, with its affine translation $T_{a}$ containing $C\cap D$.
-The intersections $C\cap T_{a}$ and $D\cap T_{a}$ are faces of the original sets, establishing $T_{a}$ as a bilateral facial reduction with parallels to Borwein-Wolkowicz facial reduction in conic programming.
-When reducing the domains of two convex functions to their joint facial subspace $T_{a}$, their relative interiors then always intersect, enabling unqualified application of classical theorems via localization.
-Key generalizations include subdifferential additivity, normal cones of intersections, a subdifferential chain rule, attained infimal convolution for $(f+g)^{*},$ and an exact Fenchel-Rockafellar dual.
-We characterize $T$ via generated faces and computationally through an iterative process converging in at most $n$ steps in $\mathbb{R}^{n}$.
-Proofs are self-contained and introduce novel concepts like facial subspaces and nested normals to describe convex boundaries and the lattice of faces.
+In convex analysis, qualification conditions (also termed constraint qualifications) help avoid pathological behavior at domain boundaries.
+In this work we remove the need for such conditions by localizing to an affine subspace—the joint supporting subspace—that contains the feasible region and ensures qualification conditions hold after localization.
+Our theory generalizes Borwein and Wolkowicz’s facial reduction beyond conic programs to convex programs of the form $f(x)+g(Ax)$.
+Intuitively, the joint supporting subspace corresponds to a bilateral facial reduction between any two convex sets.
+It enables simple qualification-free generalizations for a host of central results of convex analysis.
+These include: an exact Fenchel-Rockafellar dual; Karush-Kuhn-Tucker (KKT) optimality conditions; attained infimal convolution for convex conjugates; subdifferential sum and chain rules; and a characterization of the normal cones of the intersection of two convex sets.
+All generalizations reduce seamlessly to their original formulations when qualification conditions hold.
+We offer a number of characterizations for the joint supporting subspace, one of which is constructive.
+Our proofs are self-contained, and introduce a novel theoretical framework consisting of nested normals and supporting subspaces, which simultaneously describe both the boundary of convex sets and the lattice of faces.
 1 Introduction
-In convex analysis, the effective domain of the sum of two convex functions is the intersection of their effective domains.
-Many central results, including the summation rule of subdifferentials, the convex conjugate of sums of functions, or the strong duality of the Fenchel-Rockafellar dual, exclude the case where the domains of the convex functions intersect only on their relative boundaries.
-The conditions excluding such cases are called qualification conditions, or constraint qualifications depending on the context.
-One such condition, widespread in the seminal work [14] by Rockafellar, is that for two functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$,
-$\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))\neq\emptyset.$
-For instance, this condition appears in [14, Theorem 23.8], which establishes the summation rule of subdifferentials:
-$\partial(f+g)(x)=\partial f(x)+\partial g(x).$
-A second example is the strong duality of the Fenchel-Rockafellar dual: for $A\in\mathbb{R}^{m\times n}$, denoting by $f^{*}$ the convex conjugate of $f$,
-$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)=-\inf_{y\in\mathbb{R}^{m}}[f^{*}(A^{T}y)+g^{*}(-y)].$
-In [14, Corollary 31.2.1], this is guaranteed to hold if
-$\operatorname{relint}(\operatorname{dom}(f))\cap A^{-1}(\operatorname{relint}\operatorname{dom}(g))\neq\emptyset,$
-where $A^{-1}(\cdot)$ is the preimage for the linear map induced by the matrix $A$.
-This requirement is often called a constraint qualification, and is strikingly similar to the qualification condition for the subdifferential sum rule.
-These and a number of other results share similar conditions.
-The object of this work is to generalize such results beyond the regime where these conditions hold.
-Specifically, in Section 5 we find generalizations for the two aforementioned results, as well as a few others: the attainment of infimal convolution as the convex conjugate of the sum of two convex functions, the normal cone of the intersection of two convex sets, and a chain rule for the subdifferential.
-There are many reasons such generalizations are worthwhile, and we list two here.
-From a theoretical perspective, extending such results makes the theory more elegant and intuitive.
-A second motivation is that such generalizations could lead to improved methods in optimization, as did facial reduction in conical programming (see [3]).
-Qualification conditions are not merely theoretical artifacts; they are supported by counter examples, which indicates that these conditions reflect more or less accurately the true scope of the statements.
-We consider one such counter-example for the sum rule of subdifferentials.
+In convex analysis, many results depend on qualification conditions, of which one prominent example found in the seminal work [rockafellarConvexAnalysis1970] by Rockafellar, is that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$,
+$\operatorname{relint}\operatorname{dom}(f)\cap\operatorname{relint}\operatorname{dom}(g)\neq\emptyset,$ (1)
+where $\operatorname{relint}$ denotes the relative interior.
+This condition ensures that the feasible region $\operatorname{dom}(f+g)$ is not contained in the relative boundary of either $\operatorname{dom}(f)$ or $\operatorname{dom}(g)$.
+When Equation 1 does not hold, the geometry of the boundary of the convex sets $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$ becomes important, and requires special treatment.
+Such cases are typically excluded from consideration with qualification conditions because of the added complexity, and also because such cases are often avoidable in practice.
+The sum rule of subdifferentials,
+$\partial(f+g)(x)=\partial f(x)+\partial g(x),$
+is guaranteed by [rockafellarConvexAnalysis1970, Theorem 23.8], subject to Equation 1 (amongst other possible sufficient conditions).
+Other results of convex analysis with similar conditions include, e.g., the subdifferential chain rule, attainment of the infimal convolution in convex conjugates, strong duality of the Fenchel-Rockafellar dual, and Karush-Kuhn-Tucker (KKT) optimality conditions.
+For results concerned with optimization problems, this type of condition is usually called constraint qualifications.
+Qualification conditions are not merely theoretical artifacts; they are justified by counterexamples, and therefore reflect more or less accurately the true scope of applicability of the associated rules.
+We introduce one such counterexample for the sum rule of subdifferentials, to later show how our theory resolves this case.
 Take the function $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=\infty$ otherwise.
-This function has effective domain $\mathbb{R}_{+}$.
-Adding it to the indicator function $g(x)=\iota_{\mathbb{R}_{-}}(x)$, the sum rule states that $\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset$, when in fact, $\partial(f+g)(0)=\partial\iota_{0}(0)=\mathbb{R}$.
-The relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect.
-The qualification condition does not hold, so there is no contradiction with [14, Theorem 23.8], yet this still shows that the sum rule does not in fact apply to the sum of any two convex functions.
-Thus, an accurate generalization of the sum rule must offer a different formula for the subdifferential of the sum of two convex functions, and such a formula should ideally reduce to the classical sum rule when qualification conditions are satisfied.
-We provide such a formula in Theorem 5.1.
-Our approach to relaxing qualification conditions is to focus on the geometry of the effective domain of the convex functions.
-Qualification conditions fail when relative interiors do not intersect, in which case only the boundaries of the domains meet.
-We show that the boundaries of two convex sets always meet in a structured way: each set presents a convex face to the other, and both of these faces are revealed by the intersection with a common affine subspace, which we name the joint facial subspace $T_{a}$.
-The main property of the joint facial subspace is that given convex sets $C,D\subseteq\mathbb{R}^{n}$, with $T_{a}$ defined as in Definition 4.1, whenever $C\cap D\neq\emptyset$,
+Consider the indicator function $g(x)=\iota_{\mathbb{R}_{-}}(x)$, which is 0 on $\mathbb{R}_{-}$ and $\infty$ elsewhere.
+Naive application of the sum rule yields
+$\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset.$
+In fact, $\partial(f+g)(0)=\partial\iota_{0}(0)=\mathbb{R}$ so the sum rule does not hold.
+There is no conflict with [rockafellarConvexAnalysis1970, Theorem 23.8] because qualification conditions are violated; the relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect.
+The standard sum rule fails in this case, though we will show that it is recovered after localization to the joint supporting subspace.
+We tackle the problem of qualification conditions directly by developing geometric results describing the boundary of convex sets, allowing us to generalize a collection of results from convex analysis in one fell swoop in Section 5, with the classical formulations trivially recovered when qualifications hold (see Proposition 5.1).
+The key geometric insight is as follows.
+For any two intersecting convex sets $C,D\subseteq\mathbb{R}^{n}$ in $\mathbb{R}^{n}$, there exists a special affine subspace $T_{a}$—the joint supporting subspace—that contains $C\cap D$ and that is uniquely determined by $C$ and $D$, i.e., $T_{a}:=T_{a}(C,D)$.
+We provide multiple characterizations of the joint supporting subspace in Section 4.
+Supporting subspaces (see Definition 6.1) are a generalization of supporting hyperplanes, first defined, to our knowledge, in [brondstedIntroductionConvexPolytopes1983, Exercise 5.4].
+As intersections with supporting hyperplanes reveal exposed faces of convex sets, so do intersections with supporting subspaces reveal faces—which need not be exposed.
+As the name suggests, $T_{a}$ is a supporting subspace of both $C$ and $D$, which means that two faces of $C$ and $D$ are simultaneously revealed by intersection with $T_{a}$.
+Each of the convex sets presents a face to the other, and they meet within $T_{a}$.
+The utility of $T_{a}$ in convex analysis stems from two simple facts:
+  1. 1.
+    $C\cap D\subseteq T_{a}.$
+  2. 2.
     $\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})\neq\emptyset.$
-The results which shows this is Corollary 4.3.
-Therefore, qualification conditions always hold when localizing to the joint facial subspace.
-Roughly, we generalize results in convex analysis by replacing convex functions $f$ by their localized version $f+\iota_{T_{a}}$.
-The modified results then hold even without qualification conditions.
-And when qualification conditions do hold for the original results, then $T_{a}=\mathbb{R}^{n}$ (see Proposition 5.1), and the additional indicator function can be ignored.
-For instance, in Theorem 5.1 we state that
-$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x),$
-without any condition on $f$ and $g$ other than they be convex functions.
-This modification works for the previously encountered counter example.
-In that problem, where $f(x)=-\sqrt{x}$ and $g(x)=\iota_{(-\infty,0]}$, the joint facial subspace from Definition 4.1 is $T_{a}=\{0\}$, and so
+We prove these facts in Corollary 4.1 and Corollary 4.5 respectively.
+These two properties can be understood as follows: when reducing the ambient space to $T_{a}$, the the reduced ( localized, intersected) sets $C^{\prime}:=C\cap T_{a}$ and $D^{\prime}:=D\cap T_{a}$ preserve the original intersection ($C^{\prime}\cap D^{\prime}=C\cap D$) while satisfying the qualification condition $\operatorname{relint}(C^{\prime})\cap\operatorname{relint}(D^{\prime})\neq\emptyset$.
+Similarly, for convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, with $T_{a}:=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$, the localized functions $f^{\prime}:=f+\iota_{T_{a}},g^{\prime}:=g+\iota_{T_{a}}$ preserve the original sum ($f^{\prime}+g^{\prime}=f+g$) while satisfying Equation 1.
+Standard convex analysis results can then be applied to the localized functions $f^{\prime}$ and $g^{\prime}$, while still accurately describing the original sum $f+g$.
+For instance, in Corollary 5.1 we state that
+$\partial(f+g)(x)=\partial f^{\prime}(x)+\partial g^{\prime}(x),$ (2)
+without any condition on $f$ and $g$ other than they be proper convex functions. A host of other generalizations are obtained via minor variations on this procedure in Section 5.
+This modification to the sum rule resolves the previously encountered counterexample.
+In that problem, where $f(x)=-\sqrt{x}$ and $g(x)=\iota_{(-\infty,0]}$, the joint supporting subspace from Definition 4.1 is $T_{a}=\{0\}$, and so
 $\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x)=\partial\iota_{\{0\}}+\partial\iota_{\{0\}}=\mathbb{R},$
-which is indeed $\partial(f+g)(x)$.
-One way to interpret our result is that there are subgradients that are “hiding” in the faces of the domains of the convex functions.
-Intersecting with the joint facial subspace reveals these hidden gradients.
-The central theoretical concept introduced in this work is the joint facial subspace.
-We therefore focus much of our efforts in describing this object.
-This is the focus of Section 4, where we define the joint facial subspace in Definition 4.1 , provide two characterizations and state core property, that qualification conditions hold when localizing to it.
-Given two convex sets $C,D\subseteq\mathbb{R}^{n}$, the joint facial subspace is defined from the Minkowski difference $C-D=\{c-d\mid c\in C,d\in D\}$.
-In the first characterization, Theorem 4.1, it is made clear how the joint facial subspace relates to $C$ and $D$ individually.
-The third characterization is computational in nature, and describes how the joint facial subspace can be found from a number of steps that is bounded by the ambient dimension.
-Another important contribution of the present work is in our proof techniques.
-In Section 6, we build a theoretical framework with nested normals (Definition 6.3) and facial subspaces (Definition 6.2), two objects which together describe the lattice of faces of a convex set.
-Their combination forms a dual object which is more descriptive than the standard normal cone (see Theorem 6.2).
-Our theory is therefore at least in part a theory about the boundary of convex sets.
-From this toolkit, we are able to describe the faces of differences of convex sets CDC-D in Section 6.2, from which we derive the various properties of the joint facial subspace in Section 4.
-These objects will be further explored and refined in future work.
-The structure of the paper goes as follows.
-In Section 2 surveys prior works and Section 3 introduces some notation.
-Section 4 is as described above, with proofs deferred to Section 7.
-In Section 5, we provide generalizations of central results in convex analysis, together with their proofs, which use Corollary 4.3 from Section 4.
-In Section 6, we give a motivated introduction to our mathematical framework, introducing key concepts.
-In Section 7, we prove the results in Section 4 by using tools from Section 6.
+which indeed matches $\partial(f+g)(x)$.
+1.1 Facial reduction and computational outlook
+As we show in Section 5.3.1, our theory is a generalization of facial reduction, a preprocessing technique with strong theoretical guarantees in conic programming.
+Facial reduction reduces the problem to the smallest face of the conic constraint that contains the feasible set [massamOptimalityConditionsConeConvex1979, borweinFacialReductionConeconvex1981, borweinRegularizingAbstractConvex1981].
+After reducing to this face, Slater’s condition is provably satisfied.
+As mentioned in [drusvyatskiyManyFacesDegeneracy2017], facial reduction leads to both a reduction in the size of the problem and improved numerical stability, and has consequently been implemented in solvers for semi-definite programs (e.g., in the MOSEK solver [permenterSolvingConicOptimization2017, fribergFacialReductionMOSEK]).
+The impact of facial reduction in computation inspires hope that our theory could lead to improvements in optimization algorithms in non-conic settings where constraint qualifications may fail.
+Where facial reduction reduces to the minimal face of a cone, we reduce jointly to the smallest faces of two convex sets, both selected by intersection with the single joint supporting subspace.
+For simplicity, we frame our results as a reduction of the relevant objects (convex sets or convex functions), while keeping the ambient space $\mathbb{R}^{n}$ fixed.
+Our results are however naturally suited to a reduction of the ambient space to the joint supporting subspace, i.e., the joint supporting subspace would then become the new ambient space (see, e.g., Theorem 5.1).
+This alternative mathematical formulation would better model a reduction of the problem in a preprocessing step.
+Reduction of the ambient space to the joint supporting subspace in a preprocessing step may seem counterintuitive, as results in convex analysis are commonly specified only for real vector spaces, not affine spaces.
+This discrepancy can be mitigated by translating the affine subspace jointly with all the objects it contains until it reaches origin.
+But even this minor adjustment seems redundant, because, we believe, affine spaces seem to be the natural home for many results in convex analysis.
+In full abstraction, an affine space is a pair $(L,U)$ consisting of a set $L$ and a real vector space $U$ (the tangent space), equipped with an addition operation $L\times U\to L$ with the property that $\forall l_{1},l_{2}\in L$ there exists a unique $u\in U$ such that $l_{1}+u=l_{2}$.
+For $L$ an affine subspace of $\mathbb{R}^{n}$, the tangent space is $U=L-L=\{l_{1}-l_{2}\mid l_{1},l_{2}\in L\}$.
+Results in convex analysis can then be generalized by distinguishing tangent vectors from points in the affine space.
+For example, that affine subspaces are closed under convex combinations follows from reformulating $\lambda x+(1-\lambda)y$ for $x,y\in L$ as $y+\lambda(x-y)$, by noticing that $y\in L$ and that $\lambda(x-y)$ is in the tangent space LLL-L.
+The tangent space appears in our definition for the separation of convex sets (Definition 4.3), and justifies our repeated use of self-differences of sets SSS-S (see, e.g., Definition 6.2, Corollary 4.1, or Definition 4.2).
+1.2 Theoretical contribution
+An important contribution of the present work is the theoretical framework that we build in Section 6.
+The central objects are nested normals (Definition 6.3) and supporting subspaces (Definition 6.2), which describe local subsets of the lattice of faces of a convex set.
+Their combination forms a mathematical object that is an exact dual for convex cones (see Theorem 6.2).
+As such, the nested normals can also be understood as a local description of the boundary of convex sets, an idea that we discuss in Section 6.
+With these objects, we derive results describing the intersection of convex sets, as well as the facial structure of differences of convex sets.
+Notable technical results include Theorem 6.3, Theorem 6.4, and Theorem 6.6.
+We therefore contribute a geometric framework describing the boundaries, intersections, and differences of convex sets.
+1.3 Paper outline
+Section 2 surveys prior works, notably facial reduction in conic programming, limiting approaches to exact subdifferential calculus, and the lexicographic characterization of the faces of convex sets.
+In Section 3 we introduce some notation.
+In Section 4 we state our results describing the joint supporting subspace, with proofs deferred to Section 7.
+There, we define the joint supporting subspace in Definition 4.1, characterize it with a distinct analytical formula in Theorem 4.1, and with a constructive iterative procedure in Corollary 5.7 which terminates in a number of steps no more than the ambient dimension.
+In Theorem 4.2 we state the main property of the joint supporting subspace: that qualification conditions hold when localizing to it.
+In Section 5, we state and prove generalizations of central results in convex analysis.
+In Section 5.3.1, we show that results of facial reduction in conic programming can be recovered from our results.
+In Section 6, we provide a narrative, motivated introduction to our mathematical framework, introducing key concepts and proving associated results along the way.
+In Section 7, we prove the results of Section 4 with the tools from Section 6.
 2 Prior works
-The existing research which is the most related to this work is Facial Reduction (FR) in conic programming  [2, 1, 3], which helped guide the development of the theory in this work.
-We briefly describe this method.
-Given a feasible cone and affine constraints, Slater’s condition (that the interior of the feasible cone has at least one point satisfying the constraints), FR narrows the ambient space of the problem to the span of a “minimal” face of the feasible cone.
-After narrowing the problem via this “preprocessing” step, the problem is guaranteed to satisfy Slater’s condition.
-Our joint facial subspace does something similar, but for any two intersecting convex sets.
-We find a “minimal” face of the difference of the sets, CDC-D (see Definition 4.1), and this yields the joint facial subspace.
-The computation of the joint facial subspace can be similarly viewed as a processing step, after which the problem is reduced to the joint facial subspace.
-As specified in [3], FR can be applied with finite-step algorithms, with a number of steps bounded by the ambient dimension.
-This is very similar to our second characterization of the joint facial subspace, Theorem 7.1.
-A result which is related (where the relation is examined in [11]), is the Extended Dual System of Ramana, a dual program for semidefinite programming for which strong duality always holds [13].
-The corresponding result in our work is the exact Fenchel-Rockafellar dual in Theorem 5.5.
-From these strong parallels, we conjecture that many aspects of Facial Reduction and the Extended Dual System can be recovered—and potentially generalized—from results in this work.
-Prior efforts to obtain a qualification-free sum rule for subdifferentials have developed powerful limiting approaches that achieve exact characterizations without traditional qualifications.
-For instance, $\epsilon$-subdifferentials define approximate subgradients satisfying the subgradient inequality up to an $\varepsilon$-error, recovering the exact subdifferential as $\varepsilon$ approaches zero [4, 12, 10].
-Similarly, fuzzy sum rules approximate the subdifferential by perturbing the point $x$ at which the subdifferential is taken, deriving exact additivity rules in the limit of vanishing perturbations [6, 5].
-These methods, leveraging local neighborhood behavior, are related to ideas in variational analysis, and offer versatile tools for theoretical and computational analysis across general spaces.
-They can often be implemented with fixed tolerances for practical computation, though the exactness relies on infinite limiting processes to eliminate approximation errors.
-Though our approach is tailored to pairs of functions, our formula for the subdifferential of a sum of convex functions is exact, and can be computed in finite steps.
-One theoretical insight that was critical to developing our theory was the lexicographical characterization of facial subspaces from [8].
-This work shows that every face of a convex set can be obtained by iteratively exposing faces.
-It is a sequel of a preceding result [9], which shows a separation theorem for convex sets.
-Though the lexicographical framework was not flexible enough for our needs, prompting us to develop an independent theory, we did draw inspiration from the results and the proofs.
-In this paper, we recover this lexicographical representation, though in different terms, primarily in Lemma 6.5.
-Our proof is direct, and does not make use of a separation theorem.
+Constraint qualifications (CQs) in finite dimensions evolved from Slater’s condition [slaterLagrangeMultipliersRevisited1959], later weakened to the algebraic Mangasarian-Fromovitz CQ [mangasarianFritzJohnNecessary1967] and to the geometric Abadie CQ [abadieNonlinearProgramming1967].
+The weak Slater condition, introduced by Rockafellar [rockafellarConvexAnalysis1970], uses the notion of relative interiors.
+Subsequent relaxed CQs include the constant rank constraint qualification condition [janinDirectionalDerivativeMarginal1984], further relaxed in [qiConstantPositiveLinear2000, andreaniRelaxedConstantPositive2012], and a CQ by [robinsonFirstOrderConditions1976], which takes the form of a stability condition now known as linear metric subregularity.
+We note that in his seminal work [rockafellarConvexAnalysis1970], Rockafellar used the same qualification conditions for optimization problems and subdifferential calculus, which enabled our unified treatment of optimization and subdifferential calculus.
+There has been two major research directions generalizing results beyond qualification conditions.
+The geometric stream, to which our work belongs, focuses on the geometry of constraints and feasible sets.
+The sequential stream considers limits of increasingly accurate approximations, with exact results in the limit, and is related to variational analysis [mordukhovich_VariationalAnalysisApplications_2018, rockafellarVariationalAnalysis1997].
+The geometric stream focuses on constraint qualifications for optimization problems, while the sequential stream emphasizes subdifferential calculus, with extensions to sequential optimality conditions.
+Our work extends the geometric stream to subdifferential calculus.
+We survey qualification-free results in finite dimensions that are comparable to our results in Section 5.
+2.1 Geometric stream
+In the setting of inequality-constrained convex programs, the Fritz-John condition [johnExtremumProblemsInequalities2014] is an optimality condition that holds without qualification conditions.
+Unlike the KKT condition, it is not a sufficient condition for optimality, and is satisfied either when KKT conditions hold, or when some degeneracy in the constraints is detected, akin to applying the KKT conditions with a suitable constraint qualification.
+A qualification-free necessary and sufficient condition was formulated by [ben-talCharacterizationOptimalityConvex1976], and extended to the nonsmooth case by [wolkowiczGeometryOptimalityConditions1980].
+These works add a term to the KKT condition to deal with pathological cases; see [dharaOptimalityConditionsConvex2011, Chapter 6] for a treatment.
+Our qualification-free KKT condition in Corollary 5.6 is formulated for programs of the form $f(x)+g(Ax)$, instead of inequality-constrained convex programs, and uses faces of convex sets instead of an additional term.
+Facial Reduction (FR) generalized preprocessing ideas from linear programming to the setting of conic programming [massamOptimalityConditionsConeConvex1979, borweinFacialReductionConeconvex1981, borweinRegularizingAbstractConvex1981], with subsequent developments providing efficient computational methods for the reduction.
+An early method iteratively reduces the dimension of the problem [borweinRegularizingAbstractConvex1981, luoDualityResultsConic1997, wakiFacialReductionAlgorithms2013] until the minimal face is reached.
+Ideas of facial reduction can alternatively be framed with exact dual problems [borweinCharacterizationOptimalityAbstract1981], with important applications to semidefinite programming [ramanaExactDualityTheory1997, patakiStrongDualityConic2013].
+For in-depth literature reviews of facial reduction and its numerous applications, we refer the reader to [tuncelStrongDualityMinimal2012, drusvyatskiyManyFacesDegeneracy2017, permenterReductionMethodsSemidefinite2017].
+Many results in our work have analogues in facial reduction.
+For instance, the notion of a minimal face appears in different forms in Definition 4.1 and Theorem 4.1.
+The iterative reduction procedure of [borweinRegularizingAbstractConvex1981] finds its analogue in Theorem 7.1.
+We derive an exact Fenchel-Rockafellar dual in Corollary 5.5, echoing exact duals in conic programming.
+Finally, our qualification-free optimality condition in Corollary 5.6 generalizes similar results in the conic setting.
+2.2 Sequential stream
+The sequential approach to relaxing qualification conditions began with subdifferential calculus.
+One approach is that of $\varepsilon$-subdifferentials, which defines subgradients satisfying the subgradient inequality up to an $\varepsilon$-error, recovering the exact subdifferential in the limit of vanishing $\varepsilon$ [hiriarturruty_SubdifferentialCalculusUsing_1993, penotSubdifferentialCalculusQualification1996, mordukhovich_VariationalAnalysisApplications_2018].
+A second, similar approach is that of fuzzy sum rules, which approximate the subdifferential by perturbing the point $x$ at which the subdifferential is taken, deriving exact additivity rules in the limit of vanishing perturbations [jules_FormulasSubdifferentialsSums_2002, ioffeFuzzyPrinciplesCharacterization1998].
+These ideas were then applied to optimality conditions, where the condition of optimality takes the form of the existence of some sequence converging to the optimal point.
+One approach is to consider sequences of points converging to the point of interest [thibaultSequentialConvexSubdifferential1997], similarly to fuzzy sum rules, while another is analogous to $\varepsilon$-subdifferentials [jeyakumarInequalitySystemsGlobal1996, jeyakumarNewSequentialLagrange2003] relating them to the epigraphs of conjugate functions.
+2.3 Theory of convex faces
+A geometric insight that was critical to the development of our theory was the lexicographical characterization of faces [martinez-legazLexicographicalSeparationRn1987, martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_].
+[martinez-legazLexicographicalSeparationRn1987] derive a separation theorem from which the characterization follows.
+It turned out that the lexicographical framework was not flexible enough for our needs, prompting us to develop an analogous yet independent and self-contained theory.
+In this paper, we recover the lexicographical characterization of convex faces in Theorem 6.4, though in different terms.
+Our proof is direct, and does not make use of a separation theorem, though some separation ideas are implicit in the definition of supporting subspaces (Definition 6.1).
 3 Preliminaries
@@
 For a set $S\subseteq\mathbb{R}^{n}$, we denote by $\iota_{S}(x)$ the indicator function which is 0 if $x\in S$ and $\infty$ otherwise.
-For a set $S\subseteq\mathbb{R}^{n}$, denote by $\operatorname{conv}(S)$ its convex hull.
+For a set $S\subseteq\mathbb{R}^{n}$, denote by $\operatorname{conv}(S)$ its convex hull (smallest containing convex set).
 For $y_{1},y_{2}\in\mathbb{R}^{n}$, let $y_{1},y_{2}]:=\operatorname{conv}\{y_{1},y_{2}\}$, and $y_{1},y_{2}[:=\operatorname{conv}\{y_{1},y_{2}\}\setminus\{y_{1},y_{2}\}$.
+We also denote $\operatorname{cone}(S)$ for the conical hull of a set; the smallest cone containing $S$.
 Intuitively the notion of a face of a convex set $C$ is a subset of $C$ for which there is no line segment in $C$ that “goes through” $F$.
-The following definitions can be found inline in [16, p.
-16] and [16, p.75] respectively.
+The following definitions can be found inline in [schneider_ConvexBodiesBrunn_2013, p.
+16] and [schneider_ConvexBodiesBrunn_2013, p.75] respectively.
 Definition 3.1 (Face of a convex set).
@@
 For any set $S\subseteq C$, we denote by $F_{C}(S)$ the face of $C$ generated by $S$; it is the smallest face of $C$ that contains $S$ (it is also the intersection of all faces containing $S$, which, it can be shown, is itself a face).
-We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cap\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
-$\partial f(x):=\{v\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{n},\langle v,y-x\rangle\leq f(y)\},$
-and the convex conjugate is $f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x)$.
-4 The joint facial subspace
+For $x\in\mathbb{R}^{n}$ and a convex set $C\subseteq\mathbb{R}^{n}$, the normal cone of $C$ at $x$ is
+$N_{C}(x)=\{y\mid\forall c\in C,\langle y,c-x\rangle\leq 0\}$
+if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise.
+We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
+$\partial f(x):=\{v\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{n},f(x)+\langle v,y-x\rangle\leq f(y)\},$
+with $\partial f(x)=\emptyset$ when $x\notin\operatorname{dom}(f)$. The convex conjugate is
+$f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x).$
+4 The joint supporting subspace
+Figure 1: In each figure, the green subspace is the joint supporting subspace for the red and blue convex sets.
 In this section we investigate the following object.
-Definition 4.1 (The joint facial subspace).
-The joint facial subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ is
+Definition 4.1 (The joint supporting subspace).
+The joint supporting subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ is
 $T(C,D):=\operatorname{\mathrm{span}}F_{C-D}(0).$
-We also define $T_{a}(C,D):=T(C,D)+C\cap D$, the joint affine facial subspace.
-The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact a simple translation. This follows from the next corollary.
+We also define its affine counterpart: $T_{a}(C,D):=T(C,D)+C\cap D$.
+The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact an affine translation. This follows from the next corollary.
 Corollary 4.1 (Containment of the intersection).
-Let $C,D\in\mathbb{R}^{n}$ be convex. Then
+Let $C,D\subseteq\mathbb{R}^{n}$ be convex. Then
 $C\cap D-C\cap D\subseteq T(C,D).$
-This result follows from Lemma 6.10; we defer the proof of Corollary 4.1 to the end of Section 6.2.
+All proofs for results presented in this section are deferred to Section 7, because they require the machinery introduced in Section 6.
+Corollary 4.1 follows from Theorem 6.6; for details, the proof can be found at the end of Section 6.2.
 From Corollary 4.1, we see that picking any $x\in C\cap D$, we have $T_{a}=T+x$, independently of the choice of $x$.
-Corollary 4.1 also implies that $C\cap D\subseteq T_{a}$.
-The affine subspace $T_{a}$ is therefore the translation of $T$ that contains $C\cap D$.
-In this section we provide two characterizations of the joint facial subspace.
-The next result shows the relation of the joint facial subspace to the individual convex sets, and builds towards the first characterization.
-Proposition 4.1 (The joint facial subspace reveals faces).
+Corollary 4.1 also implies that $C\cap D\subseteq T_{a}$, showing that $T_{a}$ is the unique translation of $T$ that contains $C\cap D$.
+In this section we provide two main characterizations of the joint supporting subspace.
+The next result builds towards the first characterization, describing the relation of the joint supporting subspace with the individual convex sets.
+Lemma 4.1 (The joint supporting subspace reveals faces).
 For convex sets $C,D\subseteq\mathbb{R}^{n}$, and $T_{a}=T_{a}(C,D)$,
 $T_{a}\cap C=F_{C}(C\cap D).$
-For the first characterization we require a subspace that is related to $F_{C}(C\cap D)$.
-This subspace, the generated facial subspace, is central to our theory, as is made clear in Section 6.
-There, we will re-introduce this object in a motivated way, in Definition 6.5.
-Definition 4.2 (Preliminary definition of the generated facial subspace).
-Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, define the facial subspace of $C$ generated by $S\subseteq C$ to be
+Next we introduce the generated supporting subspace, which is central to our theory. This is made clear in Section 6, where we re-introduce this object in motivated fashion in Definition 6.5.
+Definition 4.2 (Preliminary definition of the generated supporting subspace).
+Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, define the supporting subspace of $C$ generated by $S\subseteq C$ to be
 $H_{C}(S):=\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S)).$
-The next theorem is our first characterization of the joint facial subspace.
+The next theorem is our first characterization of the joint supporting subspace.
 Theorem 4.1 (Characterization through generated faces).
-The joint facial subspace of $C,D\subseteq\mathbb{R}^{n}$ is
+The joint supporting subspace of $C,D\subseteq\mathbb{R}^{n}$ is
 $T(C,D)=H_{C}(C\cap D)+H_{D}(C\cap D).$
-Definition 4.1 and Theorem 4.1 both specify the joint facial subspace via analytical formulas.
-Our second characterization instead specifies the joint facial subspace as the result of a process with a number of steps bounded by the ambient dimension, which is more amenable to computation.
-This result, Theorem 7.1, is best stated in terms of nested normals a concept that we introduce in Definition 6.3, and motivate throughout Section 6.
+We provide two other characterizations that follow from Theorem 4.1.
+Corollary 4.2 (Joint supporting subspace as hull of faces).
+For convex sets $C,D\subseteq\mathbb{R}^{n}$,
+$T_{a}(C,D)=\operatorname{aff}(F_{C}(C\cap D)\cup F_{D}(C\cap D)).$
+Corollary 4.3 (Joint supporting subspace as difference of faces).
+For convex sets $C,D\subseteq\mathbb{R}^{n}$,
+$T(C,D)=\operatorname{\mathrm{span}}(F_{C}(C\cap D)-F_{D}(C\cap D)).$
+So far, we have characterized the joint supporting subspace with various analytical formulas.
+Our next characterization instead specifies how to obtain the joint supporting subspace constructively, from an iterative process with a number of steps bounded by the ambient dimension.
+This result, Theorem 7.1, is best stated in terms of nested normals, a concept that we introduce in Definition 6.3, and motivate throughout Section 6.
 For this reason, we defer it to Section 7.3.
@@
     $\ell\leq n$.
-Now that the joint facial subspace has been thoroughly identified, we move on to describing its core property.
+Now that the joint supporting subspace has been thoroughly identified, we move on to describing its core property.
 Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, a common qualification condition is that for convex sets $C:=\operatorname{dom}(f),D:=\operatorname{dom}(g)$,
 $\operatorname{relint}(C)\cap\operatorname{relint}(D)\neq\emptyset.$ (3)
-It was shown in [14, theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains. We define the notions of separation and proper separation below.
+It was shown in [rockafellarConvexAnalysis1970, Theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains.
+We define the notions of separation, and proper separation below.
+We define these notions in affine subspaces, an intuitive generalization of separation in Euclidean spaces.
 Definition 4.3 (Separation of sets).
-Two convex sets $C,D\subseteq\mathbb{R}^{n}$ are said to be separated if there exists some $v\in\mathbb{R}^{n}\setminus\{0\}$ such that
-$\forall c\in C,d\in D,v\cdot c\leq v\cdot d.$ (2)
+Given an affine subspace $V\subseteq\mathbb{R}^{n}$, two convex sets $C,D\subseteq V$ are said to be separated in $V$ when there exists some $u\in(V-V)\setminus\{0\}$ such that
+$\forall c\in C,d\in D,\langle u,c\rangle\leq\langle u,d\rangle.$ (4)
 The concept of proper separation is slightly stronger.
 Definition 4.4 (Proper separation of sets).
-Sets $C,D\in\mathbb{R}^{n}$ are said to be properly separated when there exists some $v\in\mathbb{R}^{n}$ such that Equation 2 holds, and such that there is some pair $(c,d)\in C\times D$ for which the inequality in Equation 2 is strict.
-Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in some affine hyperplane.
-We will show that inside the joint facial subspace, we always have a “lack of separation” of the convex sets.
-Since this regularity condition is stronger than a “lack of proper separation”, it is stronger than Equation 1.
-Theorem 4.2 (Qualification conditions always holds in the joint facial subspace).
-For two convex sets $C,D\subseteq\mathbb{R}^{n}$, $C\cap T_{a}(C,D)$ and $D\cap T_{a}(C,D)$ are not separated.
-The proof of Theorem 4.2 is the subject of Section 7.1. The next corollary follows immediately from Theorem 4.2.
-Corollary 4.3 (rint qualification holds in joint facial subspace).
-For convex sets $C,D\subseteq\mathbb{R}^{n}$,
+For an affine subspace $V\subseteq\mathbb{R}^{n}$, convex sets $C,D\subseteq\mathbb{R}^{n}$ are properly separated in $V$ when there exists some $u\in(V-V)\setminus\{0\}$ such that Equation 4 holds, and additionally such that there exists some pair $(c,d)\in C\times D$ for which the inequality in Equation 4 is strict.
+Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in a single hyperplane.
+Since separation follows from proper separation, and lack of proper separation is equivalent to intersection of relative interiors (by [rockafellarConvexAnalysis1970, Theorem 11.3]), the contrapositive statement is that
+$\text{lack of separation}\implies\text{intersection of relative interiors}.$ (5)
+In the joint supporting subspace there is always a “lack of separation” of the convex sets.
+Theorem 4.2 (Qualification conditions always hold in the joint supporting subspace).
+For two convex sets $C,D\subseteq\mathbb{R}^{n}$ with $T_{a}:=T_{a}(C,D)$, $C\cap T_{a}$ and $D\cap T_{a}$ are not separated in $T_{a}$.
+The proof of Theorem 4.2 is the subject of Section 7.1.
+Corollary 4.5 (rint qualification holds in joint supporting subspace).
+For any convex sets $C,D\subseteq\mathbb{R}^{n}$ such that $C\cap D\neq\emptyset$,
 $\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})\neq\emptyset.$
-With Corollary 4.3 we can use standard results of convex analysis from [14] by first taking intersections with $T_{a}$.
-We do this in the proof of each of the results in Section 5.
-Also note that if we only had access to weaker results than those available in [14], results that would have the stronger assumption of “no separation of the sets” instead of “intersection of relative interiors is non-empty”, the theory in this paper would still hold, and indeed we would recover the stronger results with the weaker assumption, that the intersection of relative interiors is non-empty.
+Proof of Corollary 4.5.
+The result follows from Theorem 4.2 by Equation 5, with the caveat that we defined separation in affine subspaces, which are not necessarily vector spaces, while in [rockafellarConvexAnalysis1970, theorem 11.3] this notion is defined in $\mathbb{R}^{n}$.
+Yet the argument still holds because the notions of separation, proper separation, and relative interiors are invariant under translation of $T_{a}$ to $T$.
+For any point $x_{0}\in T_{a}$, $C\cap T_{a}$ is (properly) separated from $C\cap T_{a}$ iff $C\cap T_{a}-x_{0}$ is (properly) separated from $D\cap T_{a}-x_{0}$ in $T$.
+A similar equivalence holds for intersection of relative interiors.
+∎
 5 Main results
-In this section we show five generalizations of important results of convex analysis: additivity of subdifferentials, normal cone of the intersection of two convex sets, attainment of the infimal convolution in the conjugate of a sum of two functions, the chain rule of the subdifferential, and the Fenchel-Rockafellar dual.
-All proofs are deferred to Section 7, because they require the machinery introduced in Section 6.
+With Corollary 4.5 we extend standard results of convex analysis by first localizing to $T_{a}$.
+We thus generalize the following beyond qualification conditions: subdifferential additivity; normal cones of the intersection of convex sets; attained infimal convolution for the conjugate of a sum $(f+g)^{*}$; the subdifferential chain rule; Fenchel-Rockafellar duality; and KKT conditions.
+The next result implies that when standard qualification conditions hold, the classical formulation of our generalizations is trivially recovered.
 Proposition 5.1 (Recovery of the classical results).
-In all results that follow, the classical analogue is recovered when qualification conditions hold.
-Indeed, given two convex functions $f,g$ such that their relative interiors intersect, we have $\operatorname{dom}(f),\operatorname{dom}(g)\subseteq T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$, and therefore $f+\iota_{T_{a}}=f$ and $g+\iota_{T_{a}}=g$.
-For a proper convex function $f$, the singular subdifferential for some $x\in\mathbb{R}^{n}$ is defined as $\partial^{\infty}f(x):=N_{\operatorname{dom}(f)}(x)$ when $x\in\operatorname{dom}(f)$, and is empty when $x\notin\operatorname{dom}(f)$ (see, e.g.
-[10, 15]).
-Singular subdifferentials capture the notion of an infinite slope in a particular direction.
-Theorem 5.1 (Qualification-free subdifferential additivity).
-For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper convex, $x\in\operatorname{dom}(f)\cap\operatorname{dom}(g)$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
-$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x),$
-$\partial^{\infty}(f+g)(x)=\partial^{\infty}(f+\iota_{T_{a}})(x)+\partial^{\infty}(g+\iota_{T_{a}})(x).$
-Proof of Theorem 5.1.
-By Corollary 4.3 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$,
-$\operatorname{relint}\operatorname{dom}(f+\iota_{T_{a}})\cap\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})$ (3)
-$=$ $\operatorname{relint}(\operatorname{dom}(f)\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (4)
-Therefore, a qualification condition is satisfied for the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$, and the first part of the result follows by application of [14, Theorem 23.8].
-The second part of the result follows from applying [14, Corollary 23.8.1] to the convex sets $\operatorname{dom}(f)\cap T_{a}$ and $\operatorname{dom}(g)\cap T_{a}$, by which
-$N_{(\operatorname{dom}(f)\cap T_{a})\cap(\operatorname{dom}(g)\cap T_{a})}(x)$ $=N_{\operatorname{dom}(f)\cap T_{a}}(x)+N_{\operatorname{dom}(g)\cap T_{a}}(x)$
-$=N_{\operatorname{dom}(f+\iota_{T_{a}})}(x)+N_{\operatorname{dom}(g+\iota_{T_{a}})}(x).$
-And since $N_{(\operatorname{dom}(f)\cap T_{a})\cap(\operatorname{dom}(g)\cap T_{a})}(x)=N_{\operatorname{dom}(f)\cap\operatorname{dom}(g)}(x)=N_{\operatorname{dom}(f+g)}(x)$, the second part of the result follows.
+If the relative interiors of two convex sets $C,D\subseteq\mathbb{R}^{n}$ intersect, then
+$C,D\subseteq T_{a}(C,D).$
+5.1 Subdifferential calculus
+Corollary 5.1 (Subdifferential sum rule).
+For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper convex, $x\in\mathbb{R}^{n}$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
+$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x).$
+Proof of Corollary 5.1.
+By Corollary 4.5 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$, the result follows by application of [rockafellarConvexAnalysis1970, Theorem 23.8] to the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$.
 ∎
-Theorem 5.2 (Normal cone of the intersection of convex sets).
-For $C,D\subseteq\mathbb{R}^{n}$ convex, and $x\subseteq C\cap D$, for $T_{a}=T_{a}(C,D)$,
+Corollary 5.2 (Normal cone of the intersection of convex sets).
+For $C,D\subseteq\mathbb{R}^{n}$ convex, and $x\in C\cap D$, for $T_{a}=T_{a}(C,D)$,
 $N_{C\cap D}(x)=N_{C\cap T_{a}}(x)+N_{D\cap T_{a}}(x).$
-Proof of Theorem 5.2.
-With Corollary 4.3 we can use [14, Corollary 23.8.1], from which we find that
+Proof of Corollary 5.2.
+With Corollary 4.5 we can use [rockafellarConvexAnalysis1970, Corollary 23.8.1], from which we find that
 $N_{C\cap D}(x)=N_{(C\cap T_{a})\cap(D\cap T_{a})}(x)=N_{C\cap T_{a}}(x)+N_{D\cap T_{a}}(x).$
-Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the infimal convolution is
-$(f\,\square g)(z):=\inf_{x,y\in\mathbb{R}^{n}:x+y=z}f(x)+g(y).$
-In the following results, we denote by $f^{*}$ the convex conjugate of $f$.
-Theorem 5.3 (Attained infimal convolution in the conjugate of a sum).
-For two proper convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
-$(f+g)^{*}=(f+\iota_{T_{a}})^{*}\square(g+\iota_{T_{a}})^{*},$
-where the infimum in the infimal convolution is attained.
-Proof of Theorem 5.3.
-When $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, $T_{a}=\emptyset$ and the equality holds trivially.
-Otherwise, by Corollary 4.3, $\operatorname{relint}(\operatorname{dom}(f+\iota_{T_{a}}))\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))\neq\emptyset$, and the result follows from [14, Theorem 20.1].
-∎
-Theorem 5.4 (Generalized subdifferential chain rule).
-Let $g\in\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ is a matrix. Then for $T=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
-$\partial(g)(Ax)=A^{T}\partial(g+\iota_{T})(Ax).$
-Proof of Theorem 5.4.
+In results that follow, we abuse notation by associating a matrix $A\in\mathbb{R}^{m\times n}$ with the corresponding linear map $A:\mathbb{R}^{n}\to\mathbb{R}^{m}$.
+For a set $S\subseteq\mathbb{R}^{n}$, we denote by $AS$ the set image $A(S)$, and by $A^{-1}(\cdot)$ the preimage of the linear map.
+We make no assumption on the invertibility of the matrix $A$.
+Corollary 5.3 (Subdifferential chain rule).
+Let $g\in\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ a matrix. Then for $T:=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
+$\partial(g\circ A)(x)=A^{T}\partial(g+\iota_{T})(Ax).$
+Proof of Corollary 5.3.
 Denote $T_{a}=T_{a}(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$. Note that $\forall x\in\mathbb{R}^{n}$,
@@
 If $\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)=\emptyset,$ then the result holds trivially because $g(Ax)\equiv\infty$, and $(g+\iota_{T_{a}})(Ax)\equiv\infty$ as well.
-Otherwise, by Corollary 4.3 $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$.
-Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Proposition 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$.
-Since $0\in\operatorname{\mathrm{range}}(A)$, this in fact implies that $T_{a}=T$.
-The statement then follows by applying [14, Theorem 23.9] to the composition $g+\iota_{T}]\circ A$.
+Otherwise, by Corollary 4.5, $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$, which lets us apply [rockafellarConvexAnalysis1970, Theorem 23.9] to the composition $g+\iota_{T_{a}}]\circ A$.
+Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Lemma 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$.
+Since $0\in\operatorname{\mathrm{range}}(A)$, this in fact implies that $T_{a}=T$, and the result follows.
 ∎
-In the following result, for a matrix $A\in\mathbb{R}^{m\times n}$ and a set $S\subseteq\mathbb{R}^{m}$, we denote by $A^{-1}(S)$ the preimage of $S$ under the linear map associated with the matrix $A$, making no assumption on the invertibility of the matrix $A$.
-Theorem 5.5 (Exact Fenchel-Rockafellar dual).
-For $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ proper, lower semi-continuous convex functions, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. Then
-$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)=-\inf_{y\in\mathbb{R}^{m}}[(f+\iota_{A^{-1}(T_{a})})^{*}(A^{T}y)+(g+\iota_{T_{a}})^{*}(-y)].$
+5.2 Infimal convolution
+Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the infimal convolution is
+$(f\,\square g)(z):=\inf_{x+y=z}f(x)+g(y).$
+In the following results, we denote by $f^{*}$ the convex conjugate of $f$.
+Corollary 5.4 (Attained infimal convolution).
+For two proper convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, and $T_{a}:=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
+$(f+g)^{*}=(f+\iota_{T_{a}})^{*}\square(g+\iota_{T_{a}})^{*},$
+where the infimum in the infimal convolution is attained.
+Proof of Corollary 5.4.
+Whenever $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, we have that $T_{a}=\emptyset$, in which case the equality holds trivially, with the infimum attained at any $y\in\mathbb{R}^{n}$.
+Otherwise, by Corollary 4.5 the result follows from [rockafellarConvexAnalysis1970, Theorem 20.1].
+∎
+5.3 Convex optimization
+We next examine the following general convex program.
+Problem 5.1 (Convex program).
+Given lower semicontinuous proper convex functions $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$, and a matrix $A\in\mathbb{R}^{m\times n}$,
+$\mathop{\mathrm{minimize}}_{x\in\mathbb{R}^{n}}f(x)+g(Ax).$
+A standard regularity condition for this problem is that
+$A(\operatorname{relint}\operatorname{dom}(f))\cap\operatorname{relint}\operatorname{dom}(g)\neq\emptyset,$
+which is a weakening of Slater’s condition. A program with this property is called strongly consistent [rockafellarConvexAnalysis1970, p. 300].
+The next result shows the sense in which 5.1 can be localized to the joint facial subspace $T_{a}$, guaranteeing strong consistency whenever the program is feasible, while retaining all the crucial information of the original program.
+Theorem 5.1 (Strong consistency of the localized program).
+For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ lower semicontinuous proper convex functions, a matrix $A\in\mathbb{R}^{n\times m}$, $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$, and any $x\in\mathbb{R}^{n}$,
+$f(x)+g(Ax)=(f(x)+\iota_{A^{-1}(T_{a})}(x))+(g(x)+\iota_{T_{a}}(x)).$ (6)
+And whenever $\operatorname{dom}(f+g\circ A)\neq\emptyset$,
+$A\operatorname{relint}\operatorname{dom}(f+\iota_{A^{-1}(T_{a})})\cap\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\neq\emptyset.$
 Proof of Theorem 5.1.
-In this proof we use the notation of left multiplication of functions by a matrix
-$(Af)(u):=\inf\{f(x)\mid Ax=u\}.$
-Note that $Af+\iota_{T_{a}}\equiv A(f+\iota_{A^{-1}(T_{a})})$.
-Indeed, both functions have the domain $(A\operatorname{dom}(f))\cap T_{a}$, where $A\operatorname{dom}(f):=\{Ax\mid x\in\operatorname{dom}(f)\}$, and $\forall y\in A\operatorname{dom}(f)\cap T_{a}$,
-$(Af+\iota_{T_{a}})(y)$ $=\inf\{f(x)\mid Ax=y\}$
-$=\inf\{f(x)+\iota_{Ax\in T_{a}}\mid Ax=y\}$
-$=A(f+\iota_{A^{-1}(T_{a})})(y).$
-Then, employing the change of variable $u=Ax$,
-$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)$ $=\inf_{u\in\mathbb{R}^{m}}(Af)(u)+g(u)$
-$=\inf_{u\in\mathbb{R}^{m}}(Af+\iota_{T_{a}})(u)+(g+\iota_{T_{a}})(u)$
-$=\inf_{u\in\mathbb{R}^{m}}(A(f+\iota_{A^{-1}{T_{a}}}))(u)+(g+\iota_{T_{a}})(u)$
-$=\inf_{x\in\mathbb{R}^{n}}(f+\iota_{A^{-1}{T_{a}}})(x)+(g+\iota_{T_{a}})(Ax).$
-Note that by Corollary 4.3, $\exists u\in\operatorname{relint}\operatorname{dom}(Af)\cap\operatorname{relint}\operatorname{dom}(g)$.
-Note that $\operatorname{dom}(Af)=A\operatorname{dom}(f)$, and by [14, Theorem 6.6], $\operatorname{relint}A\operatorname{dom}(f)=A\operatorname{relint}\operatorname{dom}(f)$.
-Therefore, $u\in A\operatorname{relint}\operatorname{dom}(f)\cap\operatorname{relint}\operatorname{dom}(g)$, i.e.
-there exists an $x\in\operatorname{relint}\operatorname{dom}(f)$ such that $Ax\in\operatorname{relint}\operatorname{dom}(g)$.
-With this, we satisfy a sufficient requirement of [14, Corollary 31.2.1], which guarantees strong duality and completes the proof.
+By Corollary 4.1, $(A\operatorname{dom}(f))\cap\operatorname{dom}(g)\subseteq T_{a}$, and therefore $\iota_{T_{a}}(Ax)=0$ for all $x\in\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)$.
+Additionally,
+$\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)\subseteq A^{-1}(A\operatorname{dom}(f)\cap\operatorname{dom}(g))\subseteq A^{-1}T_{a},$
+which means that $\iota_{A^{-1}T_{a}}(x)$ is 0 on $\operatorname{dom}(f+g\circ A)$.
+And for any $x\notin\operatorname{dom}(f+g\circ A)$, it is easily checked that both sides of Equation 6 evaluate to $\infty$, so Equation 6 holds for all $x\in\mathbb{R}^{n}$.
+For the second part of the result,
+$A\operatorname{relint}\operatorname{dom}(f+\iota_{A^{-1}(T_{a})})\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))$ (7)
+$=$ $\operatorname{relint}(A\operatorname{dom}(f+\iota_{A^{-1}(T_{a})}))\cap\operatorname{relint}(\operatorname{dom}(g+\iota_{T_{a}}))$ (8)
+$=$ $\operatorname{relint}(A(\operatorname{dom}(f)\cap A^{-1}(T_{a})))\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})$ (9)
+$=$ $\operatorname{relint}((A\operatorname{dom}(f))\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (10)
+where Equation 8 holds by [rockafellarConvexAnalysis1970, Theorem 6.6]; Equation 10 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and the last inequality is guaranteed by Corollary 4.5.
 ∎
-6 Nested normals and facial subspaces
-Inspired by the lexicographic characterization of convex faces of [8], we develop a framework to describe the lattice of convex faces at a point $x\in C$.
-By using this theory, we will then be able to show the results of Section 4.
-We begin with a characterization of faces of convex sets. In the next few results, we denote $X$ for a real vector space.
-Theorem 6.1 (Facial subspace characterisation of convex faces).
-Let $C\subseteq X$. A subset $F\subseteq C$ is a face of $C$ if, and only if the two following conditions hold.
+We present an exact Fenchel–Rockafellar dual for 5.1: a dual problem for which strong duality always holds and the dual attained.
+Corollary 5.5 (Exact Fenchel-Rockafellar dual).
+For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ l.s.c.
+proper convex functions, and a matrix $A\in\mathbb{R}^{n\times m}$, let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$.
+Then
+$\inf_{x\in\mathbb{R}^{n}}[f(x)+g(Ax)]=-\min_{y\in\mathbb{R}^{m}}[(f+\iota_{A^{-1}(T_{a})})^{*}(A^{T}y)+(g+\iota_{T_{a}})^{*}(-y)].$
+Proof of Corollary 5.5.
+The result follows from Theorem 5.1 with the use of [rockafellarConvexAnalysis1970, Corollary 31.2.1]. ∎
+Definition 5.1 (Minimizer).
+A minimizer of a given convex program is any point $x\in\mathbb{R}^{n}$ that attains the infimum of the program, provided that the infimum is finite.
+Next, we provide a characterization of the minimizers of 5.1, i.e., KKT optimality conditions.
+Our KKT condition holds without any further assumptions on the program; we do not require strong consistency.
+Corollary 5.6 (KKT conditions).
+Consider 5.1, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.1 if, and only if, the following conditions hold:
+  1. 1.
+    $Ax\in\partial(g+\iota_{T_{a}})^{*}(y).$
+  2. 2.
+    $A^{T}y\in\partial(f+\iota_{A^{-1}(T_{a})})(x)$.
+Proof of Corollary 5.6.
+From Theorem 5.1, the result follows from application of [rockafellarConvexAnalysis1970, Corollary 31.3.1] to the functions $f+\iota_{A^{-1}(T_{a})}$ and $g+\iota_{T_{a}}$. ∎
+Infeasible case as convention
+In the above results, the domains of the function under consideration may not intersect, in which case $f+g\circ A\equiv\infty$.
+Intersection is required for the existence of a meaningful joint supporting subspace, so it might be reasonable to make this non-trivial intersection a requirement.
+This kind of requirement was dubbed a “universal qualification condition” in the context of conic programming [tuncelStrongDualityMinimal2012].
+We observe that such a condition is not a true limitation of the theory, because the behaviour outside this condition is known and simple.
+Therefore, we instead use a convention: whenever $C\cap D=\emptyset$, we let $T_{a}(C,D):=\emptyset$ as well.
+As a result, $f+\iota_{A^{-1}(T_{a})}\equiv\infty$, and the equality in Corollary 5.5 trivially holds because both sides evaluate to $\infty$, with the value of $\infty$ attained in the dual.
+5.3.1 Facial reduction
+We derive two known facial reduction results, thereby showing that the joint supporting subspace is indeed a generalization of facial reduction.
+Problem 5.2 (Conic program).
+For $K\subseteq\mathbb{R}^{n}$ a closed convex cone, $A\in\mathbb{R}^{m\times n}$ a matrix, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$,
+$\mathop{\mathrm{minimize}}_{x\in\mathbb{R}^{n}}\langle x,c\rangle\quad\mathrm{s.t.}\quad Ax=b,\quad x\in K.$
+We apply Corollary 5.5 and Corollary 5.6 to this problem, yielding analogues of [tuncelStrongDualityMinimal2012, Theorem 4.3] and [tuncelStrongDualityMinimal2012, Corollary 4.4].
+The derived results differ from those in [tuncelStrongDualityMinimal2012] only in that our primal problem is their dual.
+Corollary 5.7 (Exact conic dual).
+For $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, $c\in\mathbb{R}^{n}$, and a closed convex cone $K\subseteq\mathbb{R}^{n}$,
+$\inf\{\langle x,c\rangle\mid x\in\mathbb{R}^{n},Ax=b,\text{ and }x\in K\}$
+$\max\{\langle b,y\rangle\mid y\in\mathbb{R}^{m}\text{ and }A^{T}y-c\in F^{\circ}\},$
+where $F=K\cap A^{-1}(H_{AK}(\{b\}))$.
+Proof of Corollary 5.7.
+We use Corollary 5.5 with $f(x):=\langle c,x\rangle+\iota_{K}(x)$, and $g:=\iota_{\{b\}}$. We find
+$T:=T(A\operatorname{dom}(f),\operatorname{dom}(g))=T(AK,\{b\})=H_{AK}(\{b\}),$
+where the last equality follows from Theorem 4.1.
+Since $AK$ is a cone, any face includes the origin, and therefore $0\in H_{AK}(\{b\})=T_{a}$, which means that $T_{a}=T$.
+The functions in the dual program are
+$f(x)+\iota_{A^{-1}(T)}(x)$ $=\langle c,x\rangle+\iota_{K}(x)+\iota_{A^{-1}(H_{AK}(\{b\}))}(x)$
+$=\langle c,x\rangle+\iota_{K\cap A^{-1}(H_{AK}(\{b\}))}(x)$
+and
+$g(Ax)+\iota_{T}(Ax)=\iota_{\{b\}}(Ax)+\iota_{H_{AK}(\{b\})}(Ax)=\iota_{\{b\}}(Ax).$
+Let $F=K\cap A^{-1}(H_{AK}(\{b\}))$, the subset of $K$ selected by the joint supporting subspace.
+The set $F$ is a cone, because $H_{AK}(\{b\})\cap AK$ is a face of the cone $AK$ by Definition 6.5 and Corollary 6.1, and $F$ is therefore itself a cone.
+Then with [rockafellarConvexAnalysis1970, Theorem 12.3, Theorem 14.1] we compute that
+$(f+\iota_{A^{-1}(T)})^{*}(A^{T}y)+(g+\iota_{T})^{*}(-y)$ $=(\langle c,\cdot\rangle+\iota_{F})^{*}(A^{T}y)+(\iota_{\{b\}})^{*}(-y)$
+$=\iota_{F^{\circ}}(A^{T}y-c)-\langle b,y\rangle.$
+We thereby find the dual problem, for which, by Corollary 5.5, strong duality holds, and the dual is attained. ∎
+Corollary 5.8 (Conic KKT conditions).
+A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.2 if, and only if, there exists some $y\in\mathbb{R}^{m}$ such that all of the following hold.
+  1. 1.
+    $Ax=b$.
+  2. 2.
+    $A^{T}y-c\in F^{\circ}$.
+  3. 3.
+    $\langle x,A^{T}y-c\rangle=0$.
+Proof of Corollary 5.8.
+We apply Corollary 5.6 to the case where $f(x):=\langle c,x\rangle+\iota_{K}(x)$ and $g:=\iota_{\{b\}}$.
+As in the proof of Corollary 5.7, we find that $f(x)+\iota_{A^{-1}(T_{a})}=\langle c,x\rangle+\iota_{F}(x)$ and $g(\bar{y})+\iota_{T_{a}}(\bar{y})=\iota_{\{b\}}(\bar{y})$.
+Then $(g+\iota_{T_{a}})^{*}(y)=\langle b,y\rangle$, and the first KKT condition, that $\partial(g+\iota_{T_{a}})^{*}(y)=\{b\}$ reduces to $Ax=b$. The second KKT condition is
+$A^{T}y$ $\in\partial(f+\iota_{A^{-1}(T_{a})})(x)$
+$=\partial(\langle c,\cdot\rangle+\iota_{F})(x)=c+N_{F}(x).$
+But since $N_{F}(x)=F^{\circ}\cap x^{\perp}$, this second condition holds if, and only if, both $x\perp A^{T}y-c$ and $A^{T}y-c\in F^{\circ}$. ∎
+6 A framework for local face lattices
+Inspired by the lexicographic characterization of convex faces of [martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_], we develop a framework describing the lattice of convex faces at a point $x\in C$, which we will then use to show the results of Section 4.
+We begin with a characterization of faces of convex sets.
+In the next few results, we denote $X$ for a real vector space.
+Theorem 6.1 (Characterization of convex faces by supporting subspaces).
+Let $C\subseteq X$ be a convex set. A subset $F\subseteq C$ is a face of $C$ if, and only if the two following conditions hold.
   1. 1.
@@
 ∎
-This characterization, surprisingly, appears to be novel to the knowledge of the author.
-It allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set.
-We call such face-selecting subspaces affine facial subspaces.
-Definition 6.1 (Affine facial subspace).
-Let $C\subseteq X$ be convex. An affine subspace $L\subseteq V$ is an affine facial subspace of $C$ if $C\setminus L$ is convex.
-The next result describes the relationship between affine facial subspaces and faces.
-Corollary 6.1 (Facial subspace relation to faces).
+Theorem 6.1 motivates the next mathematical object.
+Definition 6.1 (Supporting subspace).
+Let $C\subseteq X$ be convex. A supporting subspace $L\subseteq X$ of $C$ is an affine subspace such that $C\setminus L$ is convex.
+Though supporting subspaces are affine subspaces, we omit the adjective “affine” for brevity, mirroring the term “supporting hyperplane”.
+The notion of supporting subspaces seems to be surprisingly sparse in the literature; the only previous appearance of this notion the author could find is in [brondstedIntroductionConvexPolytopes1983, Exercise 5.4].
+This idea is of central importance to this paper; it allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set.
+The next result describes the relationship between supporting subspaces and faces.
+Corollary 6.1 (Relation of supporting subspaces to faces).
 Let $C\subseteq X$ be convex and $S\subseteq C$ non-empty.
   1. 1.
-    For every affine facial subspace $L$ of $C$, the set $F:=C\cap L$ is a face of $C$.
+    For any affine subspace $L\subseteq X$,
+[…]
   2. 2.
-    Conversely, for every face $F$ of $C$, there exists an affine facial subspace $L$ such that $F=L\cap C$ (specifically, $L=\text{aff}(F)$).
+    For any subset $F\subseteq C$,
+[…]
 Proof of Corollary 6.1.
-For the first part of the result, we show that if $L$ is an affine facial subspace, then $F:=C\cap L$ satisfies the conditions of Theorem 6.1.
+1), $(\implies)$ we show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Theorem 6.1:
   1. 1.
@@
     $C\setminus F=C\setminus[C\cap L]=C\setminus L,$
-    which is convex by definition of $L$ as an affine facial subspace.
-We show the second part of the result.
-For $F$ a face of $C$, by Theorem 6.1 $F=C\cap\operatorname{aff}(F)$ and $C\setminus F$ is convex.
-But then $C\setminus\operatorname{aff}(F)=C\setminus(\operatorname{aff}(F)\cap C)=C\setminus F$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, which makes $\operatorname{aff}(F)$ an affine facial subspace.
-∎
-In this work, we focus on “local” affine facial subspaces; the subset of affine facial subspaces that contain a point $x\in C$.
+    which is convex by definition of $L$ as a supporting subspace.
+[…]
+2), ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace.
+Indeed, by Theorem 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex.
+Note that $C\setminus F=C\setminus(C\cap\operatorname{aff}(F))=C\setminus\operatorname{aff}(F)$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, while $F=C\setminus\operatorname{aff}(F)$, which shows that $\operatorname{aff}(F)$ is a supporting subspace that selects $F$ by intersection.
+[…]
+In this work, we focus on “local” supporting subspaces; the subset of supporting subspaces that contain a point $x\in C$.
 These subsets have some additional nice algebraic properties.
-We find it useful to generalize this idea of “locality” to affine facial subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all affine subspaces that contain $S$.
-To emphasize this connection, we use the word “at” to describe the containment of $S$: ”a facial subspace is at SS”.
+We find it useful to generalize this idea of “locality” to supporting subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all supporting subspaces of $C$ that also contain $S$.
+To emphasize this connection, we say that a supporting subspace is ”at SS” to describe its containment of $S$.
 The reader should keep in mind the important special case of $S$ being a singleton $S=\{x\}$, for which the results are localized at a point.
-Any affine facial subspace $L$ at $S$ has its affine offset determined by $S$.
-Recall that LLL-L is simply the affine subspace $L$ translated in such a way that it becomes a (non-affine) subspace.
+Any supporting subspace $L$ at $S$ has its affine offset determined by $S$.
+Recall that LLL-L is the tangent space of $L$, i.e., the translation of $L$ that becomes a (non-affine) subspace.
 With this, it becomes clear that L=(LL)+SL=(L-L)+S.
-We can always deduce the correct affine offset from in $S$.
-Therefore, we find it convenient to describe the set of affine facial subspaces that contain $S$ through their translated, non-affine counterparts.
-Definition 6.2 (Facial subspace at S).
-Let $C\subseteq X$ be a convex set and $S\subseteq X$ a non-empty subset. A subspace $U\subseteq X$ is defined to be a facial subspace of $C$ at $S$ if:
+We can always deduce the correct affine offset from $S$ alone.
+Therefore, we find it convenient to describe the set of supporting subspaces that contain $S$ through their tangent spaces.
+We henceforth call the tangent spaces of supporting subspaces “supporting subspaces” as well, and whether we mean the affine subspace or its tangent space will be clear from context.
+Definition 6.2 (Supporting subspace at S).
+Let $C\subseteq X$ be a convex set and $S\subseteq X$ a non-empty subset. We call a subspace $U\subseteq X$ a supporting subspace of $C$ at $S$ if:
   1. 1.
@@
   2. 2.
-    $U+S$ is an affine facial subspace of $C$.
-The set of all facial subspaces of $C$ at $S$ is denoted $\mathcal{F}(C;S)$.
+    $U+S$ is an (affine) supporting subspace of $C$ as defined in Definition 6.1.
+We define $\mathcal{F}(C;S)$ to be the set of all subspaces $U$ satisfying both conditions above for fixed $C$ and $S$.
 When $S$ is the singleton $\{x\}$, the requirement that $S-S\subseteq U$ is trivially satisfied.
@@
 This implies that $\forall s\in S,U+s=U+S$, and so the Minkowski sum $U+S$ is a simple affine offset of $U$, for which $S\subseteq U+S$.
-In other words, the affine facial subspace $U+S$ contains $S$.
+In other words, the supporting subspace $U+S$ contains $S$.
 Next, we introduce a generalization of the notion of a normal cone. A normal cone of a convex set $C\subseteq\mathbb{R}^{n}$ is defined as
@@
 ${C^{\circ}}^{\circ}=\operatorname{cl}(C).$
-We can see that a normal cone is a dual object which “describes” the original set only up to its closure.
-We require a dual object which has sufficient descriptive power to go beyond the closure, encoding information about the boundary.
+We can see that a normal cone is a dual object which “describes” the original cone only up to its closure.
+We require a dual object which has sufficient descriptive power to go beyond the closure, by encoding information about the boundary.
 To this end, we introduce the notion of a nested normal cone.
@@
 $N_{C}^{U}(x):=N_{C\cap U}(x)\cap U.$
-We would also like for our normal cones to located at a set $S\subseteq C$, similarly to facial subspaces.
+We would also like for our normal cones to located at a set $S\subseteq C$, similarly to supporting subspaces.
 $N_{C}(S):=\{v\in\mathbb{R}^{n}\mid\forall s\in S,c\in C,\langle c-s,v\rangle\leq 0\}.$
 The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$.
-Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [7], where normal cones are at faces of polytopes, instead at points.
+Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [lu_NormalFansPolyhedral_2008], where normal cones are at faces of polytopes, instead at points.
 The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at any point in the relative interior of a convex set $S\subseteq C$ will match the normal cone of $S$.
@@
 The idea of normal cone defined at a set has an implicit relation with the notion of relative interior.
-Lemma 6.1 (nested normals in the relative interior are representative).
+Lemma 6.1 (Normal cones in the relative interior are representative).
 Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ a non-empty convex set. Let $U\subseteq\mathbb{R}^{n}$ be a subspace, and $x\in\operatorname{relint}(S)$. Then
@@
 Proof of Lemma 6.1.
-That $N_{C}^{U}(S)\subseteq N_{C}^{U}(x)$ follows from simple inclusion $x\subseteq S$.
+That $N_{C}^{U}(S)\subseteq N_{C}^{U}(x)$ follows from simple inclusion $x\in S$.
 For the containment ($\supseteq$), WLOG let $U=\mathbb{R}^{n}$.
@@
 The nested normals at a set are orthogonal with that set.
-Lemma 6.2 (Matching sets means paired normals).
+Lemma 6.2 (Normal cone everywhere is the orthogonal subspace).
 Let $C\subseteq\mathbb{R}^{n}$ be a convex set, and $U\subseteq\mathbb{R}^{n}$ a subspace. Then for any non-empty $S\subseteq C$ such that $S-S\subseteq U$,
@@
 We can think of this cone as a local description of $C$ around $x$.
-The normal cone is the polar of hthe tangent cone, and therefore “describes” the convex set $C$ around $x$ to the same extent as the tangent cone.
-In contrast, nested normals in facial subspaces of $C$ at $x$ describe the cone of feasible directions.
+The normal cone is the polar of the tangent cone, and therefore “describes” the convex set $C$ around $x$ to the same extent as the tangent cone.
+In contrast, nested normals in supporting subspaces of $C$ at $x$ describe the cone of feasible directions.
 Definition 6.4 (Cone of feasible directions).
-For a convex set $C\in\mathbb{R}^{n}$ with $x\in C$, we define the cone of feasible directions of $C$ at $x$ to be $D_{C}(x):=\{v\in\mathbb{R}^{n}\mid\exists\lambda>0:\lambda v\in C\}$.
+For a convex set $C\in\mathbb{R}^{n}$ with $x\in C$, we define the cone of feasible directions of $C$ at $x$ to be
+$D_{C}(x):=\operatorname{cone}(C-x).$
 The closure of the cone of feasible directions is the tangent cone.
 The difference is that the cone of feasible directions captures additional information about the boundary of $C$ around $x$.
-The following result shows that any convex cone (including the cone of feasible directions at any point) can be recovered from the nested normals in facial subspaces.
+The following result shows that any convex cone (including the cone of feasible directions at any point) can be recovered from the nested normals in supporting subspaces.
 Theorem 6.2 (Local polar without closure).
 Given any convex cone $T$ with $0\in T$,
-$T=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\subseteq N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}.$
+$T=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\in N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}.$
 Proof of Theorem 6.2.
-Let $T^{\prime}:=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\subseteq N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}$.
-By the definition of conditional normals, $T\subseteq T^{\prime}$.
+Let $T^{\prime}:=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\in N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}$.
+By the definition of nested normals, $T\subseteq T^{\prime}$.
 For the containment ($\supseteq$), we argue by contradiction: we assume that there is some $x\in T^{\prime}\setminus T$.
-If $x\in\mathrm{int}(T^{c})$, then $x$ is strictly separated from $T$ by a supporting hyperplane of $T$ (e.g., see [14, Theorem 11.3]).
+If $x\in\mathrm{int}(T^{c})$, then $x$ is strictly separated from $T$ by a supporting hyperplane of $T$ (e.g., see [rockafellarConvexAnalysis1970, Theorem 11.3]).
 Any such hyperplane induces a nested normal of the form $v\in N_{T}^{\mathbb{R}^{n}}(0)$, for which $v^{\perp}$ is the hyperplane, and $\langle v,x\rangle>0$.
@@
 ∎
-Another important property of nested normals is that they describe the containment structure of facial subspaces, which is closely related to the lattice structure of the faces.
+Another important property of nested normals is that they describe the containment structure of supporting subspaces, which is closely related to the lattice structure of the faces.
 We make this precise in the next two results.
-Lemma 6.3 (Nested normals induce facial subspaces).
+Theorem 6.3 (Nested normals induce supporting subspaces).
 Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, $U\subseteq\mathbb{R}^{n}$ a subspace, and $v\in N_{C}^{U}(S)$. Let $V\in\mathcal{F}(C;S)$ be such that $V\subseteq U$. Then
 $V\cap v^{\perp}\in\mathcal{F}(C;S).$
-Lemma 6.4 (Nested normals induce facial subspaces).
+Proof of Theorem 6.3.
 First, note that if $v=0$, then $V\cap v^{\perp}=V$, and there result holds. We now consider the case where $v\neq 0$.
@@
 Second, this inequality is tight precisely for those values of $y$ that are also contained in $v^{\perp}+S$, i.e.,
-$\{y\in C\cap(V+S)\mid\forall s\in S,\langle y,v\rangle=\langle s,v\rangle\}$ (5)
-$=$ $((V+S)\cap C)\cap(v^{\perp}+S)=(V\cap v^{\perp}+S)\cap C.$ (6)
+$\{y\in C\cap(V+S)\mid\forall s\in S,\langle y,v\rangle=\langle s,v\rangle\}$
+$=$ $((V+S)\cap C)\cap(v^{\perp}+S)=(V\cap v^{\perp}+S)\cap C.$
 These two facts imply that
 $(V\cap v^{\perp}+S)\cap C=\operatorname*{\arg\max}_{y\in C\cap(V+S)}\langle y,v\rangle.$
-The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, i.e., $V\cap v^{\perp}\in\mathcal{F}(C;S)$.
-We next show that all facial subspaces are “reached” by nested normals in facial subspaces.
-Lemma 6.5 (Completeness of conditional normal proof composition).
+The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, which by Corollary 6.1 implies that $V\cap v^{\perp}\in\mathcal{F}(C;S)$. ∎
+We next show that all supporting subspaces are “reached” by nested normals in supporting subspaces.
+Theorem 6.4 (Completeness of the composition of nested normals).
 Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, and $V,U\in\mathcal{F}(C;S)$ such that $U\subseteq V$.
 Then there is a sequence $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)$ with $k=\dim(V)-\dim(U)$ such that $V_{0}=V$, $V_{k}=U$, and for each $i\in\{1,\ldots,k\}$ there is a $v_{i}\in N_{C}^{V_{i-1}}(S)\setminus\{0\}$ such that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$.
-This result is similar to the lexicographic characterization of the faces of convex sets in [8, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face.
+This result is similar to the lexicographic characterization of the faces of convex sets in [martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face.
 Let us state a technical lemma that we will need for the proof.
-Lemma 6.6 (Affine slices preserve facial subspaces).
+Lemma 6.3 (Affine slices preserve supporting subspaces).
 Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ non-empty. Let $U\in\mathcal{F}(C;S)$, and $V$ any subspace such that $S-S\subseteq V$. Then $U\cap V\in\mathcal{F}(C\cap(V+S);S)$.
@@
 ∎
-Proof of Lemma 6.5.
+Proof of Theorem 6.4.
 It suffices to show the following statement: given any $U,V\in\mathcal{F}(C;S)$ such that $V\supsetneq U$, there exists a $v\in N_{C}^{V}(S)\setminus\{0\}$ such that $v\perp U$.
-If this statement is shown to hold, applying it recursively generates the sequence of facial subspaces $V_{1},\ldots,V_{k}$ in a finite number of steps.
-Note that the dimension of the facial subspaces decreases by one at each iteration while always containing $U$, which implies that $U\subseteq V_{k}$ with $\dim(U)=\dim(V_{k})$, and therefore $U=V_{k}.$
+If this statement is shown to hold, applying it recursively generates the sequence of supporting subspaces $V_{1},\ldots,V_{k}$ in a finite number of steps.
+Note that the dimension of the supporting subspaces decreases by one at each iteration while always containing $U$, which implies that $U\subseteq V_{k}$ with $\dim(U)=\dim(V_{k})$, and therefore $U=V_{k}.$
 We now show the above statement.
@@
 ∎
-Together, Lemma 6.3 and Lemma 6.5 enables a powerful tool that is central to our proofs: we can perform structural induction on the tree of facial subspaces, using only nested normals.
-Indeed, the set of facial subspaces forms a tree, where the root is $\mathbb{R}^{n}$ (note that it is always a facial subpace, and that it contains any other facial subspace), and every edge of the tree is associated with a nested normal.
-The induction statement is the following.
-Proposition 6.1 (Induction by nested normal).
-Let $C$ be a convex set, $S\subseteq C$ a non-empty subset, and a statement $A:\mathcal{F}(C;S)\to\{\text{true},\text{false}\}$. If we have show that:
+Together, Theorem 6.3 and Theorem 6.4 enables a powerful tool that is central to our proofs: we can perform structural induction on the tree of supporting subspaces, using only nested normals.
+Indeed, the set of supporting subspaces forms a tree, where the nodes are the supporting subspaces, the root is $\mathbb{R}^{n}$ (note that it is always a supporting subspace, and that it contains any other supporting subspace), and every edge of the tree is associated with a nested normal.
+Crucially, Theorem 6.4 shows that starting from the root, and generating the tree by using the nested normals, we attain all supporting subspaces.
+Induction on this tree then takes the following form.
+Corollary 6.2 (Induction by nested normal).
+Let $C$ be a convex set, $S\subseteq C$ a non-empty subset, and a statement $A:\mathcal{F}(C;S)\to\{\text{true},\text{false}\}$. If the following statements hold:
   1. 1.
-    $A(\mathbb{R}^{n}).$
+    $A(\mathbb{R}^{n}),$
   2. 2.
-    $\forall U\in\mathcal{F}(C;S)$, $v\in N^{U}_{C}(S)$, $A(U)\implies A(U\cap v^{\perp})$. Then it follows that
-    $\forall V\in\mathcal{F}(C;S),\,A(W).$
-6.1 Minimal facial subspace
+    $\forall U\in\mathcal{F}(C;S)$, $v\in N^{U}_{C}(S)$, $A(U)\implies A(U\cap v^{\perp})$, then it follows that
+    $\forall V\in\mathcal{F}(C;S),\,A(V).$
+Proof of Corollary 6.2.
+Consider any $U\in\mathcal{F}(C;S)$, and let the statement $A$ satisfy both stated conditions.
+Then by Theorem 6.4, we have supporting subspaces $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)$ and nested normals $v_{1},\ldots,v_{k}$ as in Theorem 6.4.
+$A(V_{0})=\mathrm{true}$ because $V_{0}=\mathbb{R}^{n}$.
+Also, $\forall i\in[k]$, $A(V_{i-1})\implies A(V_{i})$ because of the existence of the nested normal $v_{i}\in N_{C}^{V_{i-1}}(S)\setminus\{0\}$ with the property that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$.
+Then $A(U)$ holds by classical induction, because $U=V_{k}$.
+∎
+6.1 Minimal supporting subspace
 Recall that the smallest face of a convex set $C$ containing some $x\in C$ is $F_{C}(x)$, the face of $C$ generated by $x$.
 It is the only face with $x$ in its relative interior.
-In this section we develop an analogous notion for facial subspaces, and characterize this facial subspace by using nested normals.
-Definition 6.5 (The generated facial subspace).
-Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, we define the generated facial subspace $H_{C}(S)$ of $C$, generated by $S\subseteq\mathbb{R}^{n}$, to be the facial subspace that is contained in every other facial subspace in $\mathcal{F}(C;S)$.
+In this section, we develop an analogous notion for supporting subspaces, and characterize this supporting subspace by using nested normals.
+Definition 6.5 (The generated supporting subspace).
+Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, we define the generated supporting subspace $H_{C}(S)$ of $C$, generated by $S$, to be the supporting subspace that is contained in every other supporting subspace in $\mathcal{F}(C;S)$.
 We show existence and uniqueness for the above definition, starting with a lemma.
-Lemma 6.7 (Facial subspaces are closed under intersections).
+Lemma 6.4 (Supporting subspaces are closed under intersections).
 For $C\subseteq\mathbb{R}^{n}$ convex and $S\subseteq C$ non-empty, let $U,V\in\mathcal{F}(C;S)$. Then $U\cap V\in\mathcal{F}(C;S)$.
@@
 ∎
-Theorem 6.3 (Existence and uniqueness of the generated facial subspace).
-The generated facial subspace of Definition 6.5 exists and is unique.
-Proof of Theorem 6.3.
-We argue that there exists a facial subspace which contains no other.
-Arguing for the existence by contradiction, if every facial subspace of $C$ at $S$ properly contains some other facial subspace of $C$ at $S$, starting at any facial subspace, we can find an infinite chain of facial subspaces with dimension decreasing by one at each step.
-This is not possible because the dimension is finite.
-Therefore this chain of subspaces must terminate at some subspace which contains no other.
-We now argue that there is only one subspace that contains no other.
-Let $V_{1},V_{2}$ be two distinct minimal facial subspaces of $C$ at $S$.
-Then by Lemma 6.7, $V_{1}\cap V_{2}$ is a facial subspace of $C$ at $S$ properly contained in both $V_{1}$ and $V_{2}$, which contradicts $V_{1}$ and $V_{2}$ being minimal.
-We have shown the existence and uniqueness of a facial subspace that contains no other.
-From any facial subspace, we can start a proper containment chain as in our existence argument.
-The subspace at the end of this chain is the subspace which contains no other, which concludes the proof.
+Proposition 6.1.
+The generated supporting subspace of Definition 6.5 exists and is unique.
+Proof of Proposition 6.1.
+For existence, starting at any supporting subspace $U\in\mathcal{F}(C;S)$, either $U$ contains no other supporting subspace in $\mathcal{F}(C;S)$, or we have a $U^{\prime}\in\mathcal{F}(C;S)$ such that $U^{\prime}\subsetneq U$.
+Repeating this argument, we eventually find a supporting subspace that contains no other, because the dimension is reduced at every step, and the ambient dimension is finite.
+For uniqueness, let $V_{1},V_{2}\in\mathcal{F}(C;S)$ be such that they properly contain no other supporting subspace in $\mathcal{F}(C;S)$.
+Then by Lemma 6.4, $V_{1}\cap V_{2}\in\mathcal{F}(C;S)$ is properly contained in both $V_{1}$ and $V_{2}$, which is a contradiction.
 ∎
-As might be expected, the generated facial subspace is closely related to the generated face.
-Lemma 6.8 (Relation of the generated face to the generated facial subspace).
+As might be expected, the generated supporting subspace is closely related to the generated face.
+Lemma 6.5 (Relation between the generated face and the generated supporting subspace).
 With $C\subseteq\mathbb{R}^{n}$ convex and $S\subseteq C$ non-empty,
@@
 Proof of Lemma 6.5.
-By Theorem 6.1, $H_{C}(S)$ being the smallest facial subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own facial subspace (by Corollary 6.1), which would have to be properly contained in $H_{C}(S)$.
-For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a facial subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is an affine facial subspace, as implied by Theorem 6.1.
-Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other facial subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$.
+By Theorem 6.1, $H_{C}(S)$ being the smallest supporting subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own supporting subspace (by Corollary 6.1), which would have to be properly contained in $H_{C}(S)$.
+For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a supporting subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is a supporting subspace, as implied by Theorem 6.1.
+Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other supporting subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$.
 ∎
-We can now reformulate Definition 4.1 into an equivalent definition.
-Definition 6.6 (Equivalent definition of the joint facial subspace).
-The joint facial subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$ from Definition 4.1 is
+We can now restate Definition 4.1 with specialized terms.
+Definition 6.6 (Reformulated definition of the joint supporting subspace).
+The joint supporting subspace of two convex sets $C,D\subseteq\mathbb{R}^{n}$, as defined in Definition 4.1, is
 $T(C,D):=H_{C-D}(0).$
-The next result characterizes generated facial subspaces in terms of nested normals.
-Lemma 6.9 (The generated facial subspace is the only facial subspace with no nested normals).
-Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated facial subspace $H_{C}(S)$ is the only facial subspace of $C$ at $S$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
-Proof of Lemma 6.9.
-The generated facial subspace does not contain any non-trivial nested normals, because any non-trivial nested normal $v\in N_{C}^{H(C)}(S)\setminus\{0\}$ induces the facial subspace $H_{C}(S)\cap v^{\perp}\subsetneq H_{C}(S)$ by Lemma 6.3.
-For any facial subspace $V\in\mathcal{F}(C;S)$ such that $N_{C}^{V}(S)=\{0\}$, we argue that $V=H_{C}(S)$.
+The next result characterizes generated supporting subspaces in terms of nested normals.
+Theorem 6.5 (The generated supporting subspace is the only supporting subspace with no nested normals).
+Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated supporting subspace $H_{C}(S)$ is the only supporting subspace of $C$ at $S$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
+Proof of Theorem 6.5.
+The generated supporting subspace does not contain any non-trivial nested normals, because any non-trivial nested normal $v\in N_{C}^{H(C)}(S)\setminus\{0\}$ induces the supporting subspace $H_{C}(S)\cap v^{\perp}\subsetneq H_{C}(S)$ by Theorem 6.3.
+For any supporting subspace $V\in\mathcal{F}(C;S)$ such that $N_{C}^{V}(S)=\{0\}$, we argue that $V=H_{C}(S)$.
 Note that $H_{C}(S)\subseteq V$ by Definition 6.5.
-The inclusion cannot be proper, because otherwise, by Lemma 6.5, there would be a nontrivial nested normal $v\in N_{C}^{V}(S)\setminus\{0\}$.
+The inclusion cannot be proper, because otherwise, by Theorem 6.4, there would be a nontrivial nested normal $v\in N_{C}^{V}(S)\setminus\{0\}$.
 ∎
@@
 In this section, we show the following important result.
-Lemma 6.10 (Characterization of the conditional normals of a difference of sets).
+Theorem 6.6 (Characterization of the nested normals of a difference of sets).
 Let $C,D\subseteq\mathbb{R}^{n}$ be convex with $S\subseteq C\cap D$ non-empty, and $U\in\mathcal{F}(C-D;0)$. Then $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and
 $N^{U}_{C-D}(0)=N^{U}_{C}(S)\cap(-N^{U}_{D}(S)).$
-To show Lemma 6.10, we require the following lemma.
+To show Theorem 6.6, we require the following lemma.
 Lemma 6.6 (Normal cone of the difference of convex sets).
@@
 Then $\forall c\in C,s\in S$, $\langle v,c-s\rangle\leq 0$ because $s\in D$, and therefore $v\in N_{C}(S)$.
-By a similar argument, $v\in N_{D}(S)$, which shows the inclusion.
+By a similar argument, $v\in-N_{D}(S)$, which shows the inclusion.
 For the containment ($\supseteq$), fix some $s\in S$. let $v\in N_{C}(S)\cap(-N_{D}(S))$. Then $\forall c\in C,d\in D$
@@
 and so $v\in N_{C-D}(0)$. ∎
-Proof of Lemma 6.10.
-We argue by structured induction over the facial subspaces of CDC-D, as described in Proposition 6.1.
+Proof of Theorem 6.6.
+We argue by structured induction over the supporting subspaces of CDC-D, as described in Corollary 6.2.
 Fixing some $S\subseteq C\cap D$, the statement $A:\mathcal{F}(C-D;0)\to\{\text{true},\text{false}\}$ that we show is the following: $A(U)$ is true if, and only if, all of the following hold.
@@
     $C\cap(U+S)-D\cap(U+S)=(C-D)\cap U.$ The additional point 3) has a supportive role in this proof.
-The base case, where $U=\mathbb{R}^{n}$, holds because $\mathbb{R}^{n}$ is always a facial subspace. Point 2) is given by Lemma 6.11, and point 3) holds trivially.
+The base case, where $U=\mathbb{R}^{n}$, holds because $\mathbb{R}^{n}$ is always a supporting subspace. Point 2) is given by Lemma 6.6, and point 3) holds trivially.
 For the inductive step, let $V\in\mathcal{F}(C-D;0)$ be such that $A(V)$ is true.
@@
 In the rest of the proof, we show that $A(U)$ is true.
-Note that by point 2) of $A(V)$, $v\in N_{C}^{V}(S)\cap(-N_{D}^{V}(S))$, which by Lemma 6.3 implies that $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and this shows point 1) of $A(U)$.
+Note that by point 2) of $A(V)$, $v\in N_{C}^{V}(S)\cap(-N_{D}^{V}(S))$, which by Theorem 6.3 implies that $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and this shows point 1) of $A(U)$.
 We next show that for $C^{\prime}=C\cap(V+S)$, $D^{\prime}=D\cap(V+S)$,
@@
 Proof of Corollary 4.1.
-Considering Definition 4.1 in the light of Corollary 6.1, $T(C,D)=H_{C-D}(0)$.
-So $T\in\mathcal{F}_{C-D}(0)$, which by Lemma 6.10 implies that, taking $S=C\cap D$, $T\in\mathcal{F}_{C}(C\cap D)\cap\mathcal{F}_{D}(C\cap D)$.
-By the definition of facial subspaces (Definition 6.2), $C\cap D-C\cap D\subseteq T$.
+By Definition 6.6, $T(C,D)=H_{C-D}(0)$.
+So $T\in\mathcal{F}_{C-D}(0)$, which by Theorem 6.6 implies that, taking $S=C\cap D$, $T\in\mathcal{F}_{C}(C\cap D)\cap\mathcal{F}_{D}(C\cap D)$.
+By the definition of supporting subspaces (Definition 6.2), $C\cap D-C\cap D\subseteq T$.
 ∎
 7 Proofs
-We first show a smaller result, after which each subsection is dedicated to proving one result from Section 4.
+We begin with the proof of Proposition 5.1, after which each subsection is dedicated to one result from Section 4.
 Proof of Proposition 5.1.
@@
 where the first equality follows by Lemma 6.1 and the second by Lemma 6.2.
-Then by Lemma 6.9, $H_{C}(x)=\operatorname{\mathrm{span}}(C-C)$, and by Lemma 6.8, $F_{C}(x)=(\operatorname{\mathrm{span}}(C-C)+x)\cap C=C$.
-From Proposition 4.1, for $x\in\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))$ and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
+Then by Theorem 6.5, $H_{C}(x)=\operatorname{\mathrm{span}}(C-C)$, and by Lemma 6.5, $F_{C}(x)=(\operatorname{\mathrm{span}}(C-C)+x)\cap C=C$.
+From Lemma 4.1, for $x\in\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))$ and $T_{a}=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$,
 $T_{a}\cap\operatorname{dom}(f)=F_{\operatorname{dom}(f)}(\operatorname{dom}(f)\cap\operatorname{dom}(g))\supseteq F_{\operatorname{dom}(f)}(x)=\operatorname{dom}(f).$
 From this the results follows. ∎
-7.1 Regularity in the joint facial subspace
-We now show the main property of the joint facial subspace, Theorem 4.2.
-We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [10, Theorem 2.19].
-Lemma 7.1 (Equivalence of collisions and separation).
-For convex sets $C,D\subseteq\mathbb{R}^{n}$, and $S\subseteq C\cap D$ non-empty, the following statement holds.
+7.1 Regularity in the joint supporting subspace
+We now show the main property of the joint supporting subspace, Theorem 4.2.
+We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [mordukhovich_VariationalAnalysisApplications_2018, Theorem 2.19].
+Lemma 7.1 (Separation from local normal cones).
+For convex sets $C,D\subseteq\mathbb{R}^{n}$ and $S\subseteq C\cap D$ nonempty, the following equivalence holds:
 $N_{C}(S)\cap(-N_{D}(S))\neq\{0\}\iff C\text{ is separated from }D.$ (16)
@@
   1. 1.
-    $\forall s\in S,\langle v,s\rangle=\beta$.
+    $\langle v,s\rangle=\beta$ for all $s\in S$;
   2. 2.
-    $\forall c\in C$, $\langle y,v\rangle\geq\beta$.
+    $\langle v,c\rangle\geq\beta$ for all $c\in C$;
   3. 3.
-    $\forall d\in D$, $\langle d,v\rangle\leq\beta$.
-Then $\forall y\in C,s\in S$, $\langle v,y-s\rangle\geq 0,$ and so $v\in N_{C}(S)$.
-Similarly, $v\in N_{D}(S)$, so $v\in N_{C}(S)\cap(-N_{D}(S))$, which shows that $N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$.
+    $\langle v,d\rangle\leq\beta$ for all $d\in D$.
+Thus, for all $y\in C$ and $s\in S$, $\langle v,y-s\rangle\geq 0$, so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, whence $v\in N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$.\}\$. ∎
 Proof of Theorem 4.2.
-Recall that from Definition 6.6, $T=H_{C-D}(0)$. Then by Lemma 6.9, $N_{C-D}^{T}(C\cap D)=\{0\}$. Applying Lemma 6.10, we find that $N^{T}_{C}(C\cap D)\cap(-N^{T}_{D}(C\cap D))=\{0\}$.
+Recall that from Definition 6.6, $T=H_{C-D}(0)$. Then by Theorem 6.5, $N_{C-D}^{T}(C\cap D)=\{0\}$. Applying Theorem 6.6, we find that $N^{T}_{C}(C\cap D)\cap(-N^{T}_{D}(C\cap D))=\{0\}$.
 By applying Lemma 7.1 in the space $T_{a}$, the result follows.
-Note that Lemma 7.1 is stated in $\mathbb{R}^{n}$, yet we can apply it in $T_{a}$.
-Indeed, $T$ is a finite-dimensional subspace, and thereby isomorphic to $\mathbb{R}^{n}$, and Lemma 7.1 is invariant when translating $T$ to $T_{a}$.
+Note that Lemma 7.1 is stated in $\mathbb{R}^{n}$, yet we can apply it in $T_{a}$ because $T$ is a finite-dimensional subspace, thereby isomorphic to $\mathbb{R}^{n}$, and Lemma 7.1 is invariant when translating $T$ to $T_{a}$.
 ∎
-7.2 First characterization
-In this section we prove Proposition 4.1 and Theorem 4.1.
-Lemma 7.2 (nested constraint subspaces for nested normals).
-Let $C\subseteq\mathbb{R}^{n}$ be a convex set, $S\subseteq C$ be non-empty, and $U,V\subseteq\mathbb{R}^{n}$ be subspaces such that $S-S\subseteq U\subseteq V$ and $C\cap(V+S)\subseteq U+S$. Then
+7.2 Analytical characterizations
+In this section we prove Lemma 4.1 and Theorem 4.1.
+Lemma 7.2 (Nested normals of constrained set).
+Let $C\subseteq\mathbb{R}^{n}$ be a convex set and $S\subseteq C$ a nonempty set.
+Let $U,V\subseteq\mathbb{R}^{n}$ be subspaces such that $S-S\subseteq U\subseteq V$ and $C\cap(V+S)\subseteq U+S$.
+Then
 $N_{C}^{U}(S)+(U^{\perp}\cap V)=N_{C\cap(U+S)}^{V}(S).$
 Proof of Lemma 7.2.
-We start with the containment ($\supseteq$). Let $x\in N_{C}^{V}(S).$ Then $\forall s\in S,y\in C\cap(V+S)$, note that $y-s\in U$, and so
+We begin with the containment ($\supseteq$). Let $x\in N_{C}^{V}(S).$ Then $\forall s\in S,y\in C\cap(V+S)$, note that $y-s\in U$, and so
 $\langle\operatorname{\Pi}_{U}x,y-s\rangle=\langle x,y-s\rangle\leq 0.$
@@
 Therefore $q+z\in N_{C\cap(U+S)}^{V}(S)$, which shows the inclusion and completes the proof. ∎
-Lemma 7.3 (Conical projections yield facial subspaces).
+Lemma 7.3 (Conical projections yield supporting subspaces).
 Let $C\subseteq\mathbb{R}^{n}$ be convex, $S\subseteq C$ non-empty, and $U\in\mathcal{F}(C;S)$. Then for any $K\subseteq N^{U}_{C}(S)$,
@@
 Since $K$ is finite dimensional, there is a finite number of vectors $v_{1},\ldots,v_{m}\in K$ such that $\operatorname{\mathrm{span}}(K)=\operatorname{\mathrm{span}}(v_{1},\ldots,v_{m})$.
-Then since $v_{1},\ldots,v_{m}\in N_{C}^{U}(S)$, by repeated application of Lemma 6.3, $U\cap v_{1}^{\perp}\ldots\cap v_{m}^{\perp}$ is a facial subspace, which is the same set as $U\cap K^{\perp}$.
+Then since $v_{1},\ldots,v_{m}\in N_{C}^{U}(S)$, by repeated application of Theorem 6.3, $U\cap v_{1}^{\perp}\ldots\cap v_{m}^{\perp}$ is a supporting subspace, which is the same set as $U\cap K^{\perp}$.
 ∎
-Lemma 7.4 (Normal to the generated facial subspace of one convex set).
+Lemma 7.4 (Generated supporting subspace of one of the two convex sets).
 For convex sets $C,D\subseteq\mathbb{R}^{n}$, $T=T(C,D)$, $T_{a}=T+C\cap D$, we have
@@
 we first show that $U\in\mathcal{F}(C\cap D;C)$.
-Note that $T\in\mathcal{F}(C\cap D;C)$ by Lemma 6.10.
+Note that $T\in\mathcal{F}(C\cap D;C)$ by Theorem 6.6.
 Then we can use Lemma 7.3 because $N_{C}^{T}(C\cap T_{a})\subseteq N_{C}^{T}(C\cap D)$, which shows that $U\in\mathcal{F}(C\cap D;C)$ as well.
-In the remainder of the proof we show that the assumption of Lemma 6.9 holds for $U$, namely that $N_{C}^{U}(C\cap D)=\{0\}$.
+In the remainder of the proof we show that the assumption of Theorem 6.5 holds for $U$, namely that $N_{C}^{U}(C\cap D)=\{0\}$.
 We first show that $N_{C}^{U}(C\cap T_{a})=\{0\}$.
@@
 $N_{C}^{U}(C\cap T_{a})=N_{C}^{U}(x)\supseteq N_{C}^{U}(C\cap D).$
-Then $N_{C}^{U}(C\cap D)=\{0\}$, and so the result follows from Lemma 6.9. ∎
-Proof of Proposition 4.1.
+Then $N_{C}^{U}(C\cap D)=\{0\}$, and so the result follows from Theorem 6.5. ∎
+We are now in a position to prove the analytical characterizations in Section 4.
+Proof of Lemma 4.1.
 From Lemma 7.4, $H_{C}(C\cap D)=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}$.
@@
 $H_{C}(C\cap D)+C\cap D\supseteq C\cap T_{a}.$ (17)
-Additionally, since we know that $T\in\mathcal{F}(C\cap D;C)$ from Lemma 6.10, by definition of the generated facial subspace we know that $H_{C}(C\cap D)\subseteq T$. Therefore,
+Additionally, since we know that $T\in\mathcal{F}(C\cap D;C)$ from Theorem 6.6, by definition of the generated supporting subspace we know that $H_{C}(C\cap D)\subseteq T$. Therefore,
 $H_{C}(C\cap D)+C\cap D\subseteq T_{a}.$ (18)
@@
 Proof of Theorem 4.1.
-From Proposition 4.1, we have that $T\in\mathcal{F}(C;C\cap D)$ is a facial subspace of $C$ that contains $C\cap D$, and so by Definition 6.5, $H_{C}(C\cap D)\subseteq T$.
+From Lemma 4.1, we have that $T\in\mathcal{F}(C;C\cap D)$ is a supporting subspace of $C$ that contains $C\cap D$, and so by Definition 6.5, $H_{C}(C\cap D)\subseteq T$.
 Therefore, for $^{\prime}:=H_{C}(C\cap D)+H_{D}(C\cap D), TT$$^{\prime}\subseteq T$.
@@
 By a similar argument, $v,-v\in N_{D}^{T}(C\cap D)$.
-But then $v\in N_{C}^{T}(C\cap D)\cap(-N_{D}^{T}(C\cap D))$, which by Lemma 6.10 means that $v\in N_{C-D}^{T}(0)$.
-But by Lemma 6.9, $T$ cannot contain such a nested normal because by Definition 6.6, $T=H_{C-D}(0)$.
+But then $v\in N_{C}^{T}(C\cap D)\cap(-N_{D}^{T}(C\cap D))$, which by Theorem 6.6 means that $v\in N_{C-D}^{T}(0)$.
+But by Theorem 6.5, $T$ cannot contain such a nested normal because by Definition 6.6, $T=H_{C-D}(0)$.
 Therefore, by contradiction $T=T^{\prime}$.
 ∎
+Next, we present the proofs for the corollaries of Theorem 4.1, starting with a simple lemma.
+Lemma 7.5 (Formulation of the tangent space).
+For a convex set $C\subseteq\mathbb{R}^{n}$ and $S\subseteq C$ non-empty,
+$\operatorname{\mathrm{span}}(C-C)=\operatorname{\mathrm{span}}(C-S).$
+Proof of Lemma 7.5.
+That $\operatorname{\mathrm{span}}(C-C)\supseteq\operatorname{\mathrm{span}}(C-S)$ is trivial, so we only show $\operatorname{\mathrm{span}}(C-C)\subseteq\operatorname{\mathrm{span}}(C-S)$.
+It suffices to show that $C-C\subseteq\operatorname{\mathrm{span}}(C-S)$; let $u,v\in C$, then for any fixed $x\in S$,
+$u-v=(u-x)-(v-x)\in\operatorname{\mathrm{span}}(C-S).$
+Proof of Corollary 4.3.
+In this proof, we denote $F_{C}:=F_{C}(C\cap D)$ and $F_{D}:=F_{D}(C\cap D)$.
+$\operatorname{\mathrm{span}}(F_{C}-F_{D})$ $=\operatorname{\mathrm{span}}(F_{C}-F_{D})+(C\cap D-C\cap D)$ (19)
+$=\operatorname{\mathrm{span}}(F_{C}-F_{D}+(C\cap D-C\cap D))$ (20)
+$=\operatorname{\mathrm{span}}((F_{C}-C\cap D)-(F_{D}-C\cap D))$ (21)
+$=\operatorname{\mathrm{span}}(F_{C}-C\cap D)+\operatorname{\mathrm{span}}(F_{D}-C\cap D)$ (22)
+$=\operatorname{\mathrm{span}}(F_{C}-F_{C})+\operatorname{\mathrm{span}}(F_{D}-F_{D}),$ (23)
+$=H_{C}(C\cap D)+H_{D}(C\cap D).$ (24)
+where Equation 19 holds because
+$(C\cap D-C\cap D)\subseteq F_{C}-F_{D}\subseteq\operatorname{\mathrm{span}}(F_{C}-F_{D}),$
+and Equation 23 follows from Lemma 7.5. The result then follows from Theorem 4.1. ∎
+Proof of Corollary 4.2.
+Throughout this proof, we omit the argument $C\cap D$, i.e., $H_{C}:=H_{C}(C\cap D)$, and similarly for $H_{D},F_{C},F_{D}$.
+We also define $\bar{F}_{C}=F_{C}-C\cap D$ and $\bar{F}_{D}=F_{D}-C\cap D$.
+By Lemma 7.5,
+$H_{C}=\operatorname{\mathrm{span}}(F_{C}-F_{C})=\operatorname{\mathrm{span}}(\bar{F}_{C}).$
+With this, starting with $T_{a}$ as in Theorem 4.1,
+$T_{a}(C,D)$ $=H_{C}+H_{D}+C\cap D$
+$=\operatorname{\mathrm{span}}(\bar{F}_{C})+\operatorname{\mathrm{span}}(\bar{F}_{D})+C\cap D$
+$=\operatorname{\mathrm{span}}(\operatorname{\mathrm{span}}(\bar{F}_{C})\cup\operatorname{\mathrm{span}}(\bar{F}_{D}))+C\cap D$
+$=\operatorname{\mathrm{span}}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
+$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
+$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D}+C\cap D)$
+$=\operatorname{aff}(F_{C}\cup F_{D}).$
 7.3 Second characterization
@@
 Proof of Theorem 7.1.
-First, $(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])=N^{T_{i}}_{C-D}(S)$ by Lemma 6.10.
+First, $(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])=N^{T_{i}}_{C-D}(S)$ by Theorem 6.6.
 Applying Lemma 7.3 at each step, we find that $\forall i\in\{0,\ldots,\ell\},\,T_{i}\in\mathcal{F}(C-D;S)$.
@@
 On the other hand, the fact that $T_{\ell}=T_{\ell+1}$ implies that $N^{T_{\ell}}_{C-D}(S)=\{0\}$.
-Then from Lemma 6.9 it follows that $T_{\ell}=H_{C-D}(S)$, which is $T(C,D)$ from Definition 6.6.
+Then from Theorem 6.5 it follows that $T_{\ell}=H_{C-D}(S)$, which is $T(C,D)$ from Definition 6.6.
 ∎
-7.4 Aknowledgements
-The author would like to thank Sharvaj Kubal, prof.
+7.4 Acknowledgements
+The author would like to thank Sharvaj Kubal, Naomi Graham, prof.
 Yaniv Plan, and prof.
@@
 A special thanks to prof.
-Michael Friedlander for pointing out that a generalization of the Fenchel-Rockafellar dual would be a promising avenue of research.
+Michael Friedlander for pointing out that a generalization of the Fenchel-Rockafellar dual and optimality conditions would be a promising research direction.
 This paper was produced with the Latex Exporter plugin in Obsidian.
-range
-pages36 rangepages12 rangepages13 rangepages12 rangepages25 rangepages17 rangepages34

### v4

@@
 Our proofs are self-contained, and introduce a novel theoretical framework consisting of nested normals and supporting subspaces, which simultaneously describe both the boundary of convex sets and the lattice of faces.
+Keywords: facial reduction, subdifferential calculus, exact duality, qualification conditions, constraint qualifications, convex analysis
 1 Introduction
@@
 $\exists u\in\mathbb{R}^{n}\setminus\{0\}:F=\operatorname*{\arg\max}_{c\in C}\langle c,u\rangle.$
-For any set $S\subseteq C$, we denote by $F_{C}(S)$ the face of $C$ generated by $S$; it is the smallest face of $C$ that contains $S$ (it is also the intersection of all faces containing $S$, which, it can be shown, is itself a face).
+For any set $S\subseteq C$, we denote by $F_{C}(S)$ the face of $C$ generated by $S$; it is the smallest face of $C$ that contains $S$.
+(It is also the intersection of all faces containing $S$, which, it can be shown, is itself a face.)
 For $x\in\mathbb{R}^{n}$ and a convex set $C\subseteq\mathbb{R}^{n}$, the normal cone of $C$ at $x$ is
@@
 if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise.
-We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
+We also recall some objects from convex analysis.
+Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$.
+For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
 $\partial f(x):=\{v\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{n},f(x)+\langle v,y-x\rangle\leq f(y)\},$
@@
 Let $C,D\subseteq\mathbb{R}^{n}$ be convex sets and fix any $x\in C\cap D$. Beginning with $T_{0}=\mathbb{R}^{n}$, let
-$T_{i+1}=T_{i}\cap(N_{C}(x)\cap[-N_{D}(x)]\cap T_{i})^{\perp}.$
+$T_{i+1}=T_{i}\cap(N_{C\cap(T_{i}+x)}(x)\cap[-N_{D\cap(T_{i}+x)}(x)])^{\perp}.$
 Let $\ell$ be the smallest number such that $T_{\ell+1}=T_{\ell}$. Then:
@@
     $\ell\leq n$.
+We defer the proof to Section 7.3.
 Now that the joint supporting subspace has been thoroughly identified, we move on to describing its core property.
@@
 We thus generalize the following beyond qualification conditions: subdifferential additivity; normal cones of the intersection of convex sets; attained infimal convolution for the conjugate of a sum $(f+g)^{*}$; the subdifferential chain rule; Fenchel-Rockafellar duality; and KKT conditions.
-The next result implies that when standard qualification conditions hold, the classical formulation of our generalizations is trivially recovered.
+The next result implies that when standard qualification conditions hold, the classical formulations of our generalizations are trivially recovered.
 Proposition 5.1 (Recovery of the classical results).
 If the relative interiors of two convex sets $C,D\subseteq\mathbb{R}^{n}$ intersect, then
-$C,D\subseteq T_{a}(C,D).$
+$C\cup D\subseteq T_{a}(C,D).$
 5.1 Subdifferential calculus
@@
 Corollary 5.3 (Subdifferential chain rule).
-Let $g\in\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ a matrix. Then for $T:=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
+Let $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ a matrix. Then for $T:=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
 $\partial(g\circ A)(x)=A^{T}\partial(g+\iota_{T})(Ax).$
@@
 Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the infimal convolution is
-$(f\,\square g)(z):=\inf_{x+y=z}f(x)+g(y).$
+$(f\,\square g)(z):=\inf_{x,y:x+y=z}f(x)+g(y).$
 In the following results, we denote by $f^{*}$ the convex conjugate of $f$.
@@
 $A(\operatorname{relint}\operatorname{dom}(f))\cap\operatorname{relint}\operatorname{dom}(g)\neq\emptyset,$
-which is a weakening of Slater’s condition. A program with this property is called strongly consistent [rockafellarConvexAnalysis1970, p. 300].
+which is a weakening of Slater’s condition. A program with this property is called strongly consistent [rockafellarConvexAnalysis1970, Theorem 31.2].
 The next result shows the sense in which 5.1 can be localized to the joint facial subspace $T_{a}$, guaranteeing strong consistency whenever the program is feasible, while retaining all the crucial information of the original program.
@@
 Proof of Theorem 5.1.
-By Corollary 4.1, $(A\operatorname{dom}(f))\cap\operatorname{dom}(g)\subseteq T_{a}$, and therefore $\iota_{T_{a}}(Ax)=0$ for all $x\in\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)$.
+By Corollary 4.1, $A\operatorname{dom}(f)\cap\operatorname{dom}(g)\subseteq T_{a}$, and therefore $\iota_{T_{a}}(Ax)=0$ for all $x\in\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)$.
 Additionally,
@@
 $=$ $\operatorname{relint}(A(\operatorname{dom}(f)\cap A^{-1}(T_{a})))\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})$ (9)
-$=$ $\operatorname{relint}((A\operatorname{dom}(f))\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (10)
-where Equation 8 holds by [rockafellarConvexAnalysis1970, Theorem 6.6]; Equation 10 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and the last inequality is guaranteed by Corollary 4.5.
+$=$ $\operatorname{relint}(A\operatorname{dom}(f)\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (10)
+where Equation 8 holds by [rockafellarConvexAnalysis1970, Theorem 6.6]; Equation 10 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and that the set in Equation 10 is non-empty is guaranteed by Corollary 4.5.
 ∎
@@
 Proof of Corollary 6.1.
-1), $(\implies)$ we show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Theorem 6.1:
+1) ($\implies$) We show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Theorem 6.1:
   1. 1.
@@
 […]
-2), ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace.
+2) ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace.
 Indeed, by Theorem 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex.
@@
 Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
-$N_{C}^{U}(x):=N_{C\cap U}(x)\cap U.$
+$N_{C}^{U}(x):=N_{C\cap(x+U)}(x)\cap U.$
 We would also like for our normal cones to located at a set $S\subseteq C$, similarly to supporting subspaces.
@@
 The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$.
-Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [lu_NormalFansPolyhedral_2008], where normal cones are at faces of polytopes, instead at points.
-The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at any point in the relative interior of a convex set $S\subseteq C$ will match the normal cone of $S$.
+Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [lu_NormalFansPolyhedral_2008], where normal cones are defined for faces of polytopes, instead of at points.
 Combining the two generalizations, we get the following.
@@
 $N_{C}^{U}(S):=\{v\in U\mid\forall c\in C^{\prime},s\in S,\langle c-s,v\rangle\leq 0\}$
-is a nested normal cone of $C$ at $S$, and we label any $v\in N_{C}^{U}(S)$ to be a nested normal of $C$ at $S$ in $U$.
-The idea of normal cone defined at a set has an implicit relation with the notion of relative interior.
+is a nested normal cone of $C$ at $S$, and we label any $v\in N_{C}^{U}(S)$ a nested normal of $C$ at $S$ in $U$.
+The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at points in the relative interior match the normal cone of the set.
 Lemma 6.1 (Normal cones in the relative interior are representative).
-Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ a non-empty convex set. Let $U\subseteq\mathbb{R}^{n}$ be a subspace, and $x\in\operatorname{relint}(S)$. Then
+Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ a non-empty convex set. Let $U\subseteq\mathbb{R}^{n}$ be a subspace such that $S-S\subseteq U$, and $x\in\operatorname{relint}(S)$. Then
 $N_{C}^{U}(S)=N_{C}^{U}(x).$
@@
 Let $y\in N_{C}^{U}(S)$.
-By Definition 6.3, $\forall z,x\in S,$ since $x\in S$ and $z\in C$, $\langle x-z,y\rangle\leq 0,$ and since $x\in C$ and $z\in S$ $\langle x-z,y\rangle\geq 0$.
+By Definition 6.3, $\forall z,x\in S,$ since $x\in C$ and $z\in S$, $\langle x-z,y\rangle\leq 0,$ and since $x\in S$ and $z\in C$, $\langle x-z,y\rangle\geq 0$.
 Then $y\perp S-S$, which shows the inclusion.
@@
     $\langle v,d\rangle\leq\beta$ for all $d\in D$.
-Thus, for all $y\in C$ and $s\in S$, $\langle v,y-s\rangle\geq 0$, so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, whence $v\in N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$.\}\$. ∎
+Thus, for all $y\in C$ and $s\in S$, $\langle v,y-s\rangle\geq 0$, so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, whence $v\in N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$. ∎
 Proof of Theorem 4.2.
@@
 ∎
+Proof of Corollary 4.4.
+We consider the formula
+$T_{i+1}=T_{i}\cap(N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)])^{\perp}.$
+from Theorem 7.1. We show that the r.h.s. is the same set as
+$T_{i}\cap(N_{C\cap(T_{i}+x)}(x)\cap[-N_{D\cap(T_{i}+x)}(x)])^{\perp}.$
+We use a substitution as in Lemma 7.2 with $V=\mathbb{R}^{n}$.
+$T_{i}\cap(N_{C\cap(T_{i}+x)}(x)\cap[-N_{D\cap(T_{i}+x)}(x)])^{\perp}$
+$=$ $T_{i}\cap((N^{T_{i}}_{C}(x)+T_{i}^{\perp})\cap[-N^{T_{i}}_{D}(x)+T_{i}^{\perp}])^{\perp}$
+$=$ $T_{i}\cap(N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)]+T_{i}^{\perp})^{\perp}$
+$=$ $T_{i}\cap(N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)])^{\perp}.$
+The last equality holds because $N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)]\subseteq T_{i}$, so we may use the identity that $T_{i}\cap(S+T_{i}^{\perp})^{\perp}=T_{i}\cap S^{\perp}$ for any set $S\subseteq T_{i}$.
+∎
 7.4 Acknowledgements
@@
 This paper was produced with the Latex Exporter plugin in Obsidian.
+Declarations
+The author is supported by the University of British Columbia. The author has no relevant financial or non-financial interests to disclose.

### v5

@@
-Qualification-free convex analysis via the joint supporting subspace
+Bilateral facial reduction: qualification-free subdifferential calculus and exact duality
 Abstract
-In convex analysis, qualification conditions (also termed constraint qualifications) help avoid pathological behavior at domain boundaries.
-In this work we remove the need for such conditions by localizing to an affine subspace—the joint supporting subspace—that contains the feasible region and ensures qualification conditions hold after localization.
-Our theory generalizes Borwein and Wolkowicz’s facial reduction beyond conic programs to convex programs of the form $f(x)+g(Ax)$.
-Intuitively, the joint supporting subspace corresponds to a bilateral facial reduction between any two convex sets.
-It enables simple qualification-free generalizations for a host of central results of convex analysis.
-These include: an exact Fenchel-Rockafellar dual; Karush-Kuhn-Tucker (KKT) optimality conditions; attained infimal convolution for convex conjugates; subdifferential sum and chain rules; and a characterization of the normal cones of the intersection of two convex sets.
-All generalizations reduce seamlessly to their original formulations when qualification conditions hold.
-We offer a number of characterizations for the joint supporting subspace, one of which is constructive.
-Our proofs are self-contained, and introduce a novel theoretical framework consisting of nested normals and supporting subspaces, which simultaneously describe both the boundary of convex sets and the lattice of faces.
+Qualification conditions (also termed constraint qualifications) help avoid pathological behavior at domain boundaries in convex analysis.
+By generalizing facial reduction from conic programming to general convex programs of the form $f(x)+g(Ax)$, we provide qualification-free generalizations of several key results: an exact Fenchel-Rockafellar dual, KKT optimality conditions, an attained infimal convolution for the conjugate of a sum, subdifferential sum and chain rules, and normal cones of intersections.
+All our results reduce seamlessly to their original formulations when qualification conditions hold.
+The core insight is that for a sum of two convex functions, there is an affine subspace—the joint supporting subspace— that contains the feasible region, and such that qualification conditions hold when restricting the effective domain of each function to it.
+We offer a number of characterizations for the joint supporting subspace, including one that obtains the affine subspace via iterative, bilateral reduction between the two domains.
+In our proofs, which are self-contained, we develop a structured induction on faces where inductive steps are associated with normal vectors nested in supporting subspaces (a generalization of supporting hyperplanes).
+With this tool, we characterize the facial structure of the difference of two convex sets from the facial structures of the individual convex sets.
 Keywords: facial reduction, subdifferential calculus, exact duality, qualification conditions, constraint qualifications, convex analysis
 1 Introduction
-In convex analysis, many results depend on qualification conditions, of which one prominent example found in the seminal work [rockafellarConvexAnalysis1970] by Rockafellar, is that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$,
+In convex analysis, many results require so-called qualification conditions (sometimes called constraint qualifications), such as
 $\operatorname{relint}\operatorname{dom}(f)\cap\operatorname{relint}\operatorname{dom}(g)\neq\emptyset,$ (1)
-where $\operatorname{relint}$ denotes the relative interior.
-This condition ensures that the feasible region $\operatorname{dom}(f+g)$ is not contained in the relative boundary of either $\operatorname{dom}(f)$ or $\operatorname{dom}(g)$.
-When Equation 1 does not hold, the geometry of the boundary of the convex sets $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$ becomes important, and requires special treatment.
-Such cases are typically excluded from consideration with qualification conditions because of the added complexity, and also because such cases are often avoidable in practice.
-The sum rule of subdifferentials,
-$\partial(f+g)(x)=\partial f(x)+\partial g(x),$
-is guaranteed by [rockafellarConvexAnalysis1970, Theorem 23.8], subject to Equation 1 (amongst other possible sufficient conditions).
-Other results of convex analysis with similar conditions include, e.g., the subdifferential chain rule, attainment of the infimal convolution in convex conjugates, strong duality of the Fenchel-Rockafellar dual, and Karush-Kuhn-Tucker (KKT) optimality conditions.
-For results concerned with optimization problems, this type of condition is usually called constraint qualifications.
-Qualification conditions are not merely theoretical artifacts; they are justified by counterexamples, and therefore reflect more or less accurately the true scope of applicability of the associated rules.
-We introduce one such counterexample for the sum rule of subdifferentials, to later show how our theory resolves this case.
-Take the function $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=\infty$ otherwise.
-Consider the indicator function $g(x)=\iota_{\mathbb{R}_{-}}(x)$, which is 0 on $\mathbb{R}_{-}$ and $\infty$ elsewhere.
+This condition can be understood as “avoiding the boundary”; it ensures that the feasible region $\operatorname{dom}(f+g)$ is not contained in the relative boundary of either $\operatorname{dom}(f)$ or $\operatorname{dom}(g)$.
+When Equation 1 fails to hold, the geometry of the boundary of the convex sets $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$ starts to matter, and such cases may require special treatment.
+Qualification conditions are not merely theoretical artifacts; they are justified by counterexamples, and therefore reflect more or less accurately the true scope of applicability of the associated formulas.
+We introduce one such counterexample for the sum rule of subdifferentials.
+(We will later show how our theory resolves this case.) Take the function $f(x):=-\sqrt{x}$ when $x\geq 0$ and $f(x):=\infty$ otherwise.
+Consider the indicator function $g(x)=\iota_{(-\infty,0]}(x)$, which is 0 on $(-\infty,0]$ and $\infty$ elsewhere.
 Naive application of the sum rule yields
-$\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset.$
-In fact, $\partial(f+g)(0)=\partial\iota_{0}(0)=\mathbb{R}$ so the sum rule does not hold.
-There is no conflict with [rockafellarConvexAnalysis1970, Theorem 23.8] because qualification conditions are violated; the relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect.
-The standard sum rule fails in this case, though we will show that it is recovered after localization to the joint supporting subspace.
-We tackle the problem of qualification conditions directly by developing geometric results describing the boundary of convex sets, allowing us to generalize a collection of results from convex analysis in one fell swoop in Section 5, with the classical formulations trivially recovered when qualifications hold (see Proposition 5.1).
-The key geometric insight is as follows.
-For any two intersecting convex sets $C,D\subseteq\mathbb{R}^{n}$ in $\mathbb{R}^{n}$, there exists a special affine subspace $T_{a}$—the joint supporting subspace—that contains $C\cap D$ and that is uniquely determined by $C$ and $D$, i.e., $T_{a}:=T_{a}(C,D)$.
-We provide multiple characterizations of the joint supporting subspace in Section 4.
-Supporting subspaces (see Definition 6.1) are a generalization of supporting hyperplanes, first defined, to our knowledge, in [brondstedIntroductionConvexPolytopes1983, Exercise 5.4].
-As intersections with supporting hyperplanes reveal exposed faces of convex sets, so do intersections with supporting subspaces reveal faces—which need not be exposed.
-As the name suggests, $T_{a}$ is a supporting subspace of both $C$ and $D$, which means that two faces of $C$ and $D$ are simultaneously revealed by intersection with $T_{a}$.
-Each of the convex sets presents a face to the other, and they meet within $T_{a}$.
-The utility of $T_{a}$ in convex analysis stems from two simple facts:
-  1. 1.
-    $C\cap D\subseteq T_{a}.$
-  2. 2.
-    $\operatorname{relint}(C\cap T_{a})\cap\operatorname{relint}(D\cap T_{a})\neq\emptyset.$
-We prove these facts in Corollary 4.1 and Corollary 4.5 respectively.
-These two properties can be understood as follows: when reducing the ambient space to $T_{a}$, the the reduced ( localized, intersected) sets $C^{\prime}:=C\cap T_{a}$ and $D^{\prime}:=D\cap T_{a}$ preserve the original intersection ($C^{\prime}\cap D^{\prime}=C\cap D$) while satisfying the qualification condition $\operatorname{relint}(C^{\prime})\cap\operatorname{relint}(D^{\prime})\neq\emptyset$.
-Similarly, for convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, with $T_{a}:=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$, the localized functions $f^{\prime}:=f+\iota_{T_{a}},g^{\prime}:=g+\iota_{T_{a}}$ preserve the original sum ($f^{\prime}+g^{\prime}=f+g$) while satisfying Equation 1.
-Standard convex analysis results can then be applied to the localized functions $f^{\prime}$ and $g^{\prime}$, while still accurately describing the original sum $f+g$.
-For instance, in Corollary 5.1 we state that
-$\partial(f+g)(x)=\partial f^{\prime}(x)+\partial g^{\prime}(x),$ (2)
-without any condition on $f$ and $g$ other than they be proper convex functions. A host of other generalizations are obtained via minor variations on this procedure in Section 5.
-This modification to the sum rule resolves the previously encountered counterexample.
-In that problem, where $f(x)=-\sqrt{x}$ and $g(x)=\iota_{(-\infty,0]}$, the joint supporting subspace from Definition 4.1 is $T_{a}=\{0\}$, and so
-$\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x)=\partial\iota_{\{0\}}+\partial\iota_{\{0\}}=\mathbb{R},$
-which indeed matches $\partial(f+g)(x)$.
-1.1 Facial reduction and computational outlook
-As we show in Section 5.3.1, our theory is a generalization of facial reduction, a preprocessing technique with strong theoretical guarantees in conic programming.
-Facial reduction reduces the problem to the smallest face of the conic constraint that contains the feasible set [massamOptimalityConditionsConeConvex1979, borweinFacialReductionConeconvex1981, borweinRegularizingAbstractConvex1981].
-After reducing to this face, Slater’s condition is provably satisfied.
-As mentioned in [drusvyatskiyManyFacesDegeneracy2017], facial reduction leads to both a reduction in the size of the problem and improved numerical stability, and has consequently been implemented in solvers for semi-definite programs (e.g., in the MOSEK solver [permenterSolvingConicOptimization2017, fribergFacialReductionMOSEK]).
-The impact of facial reduction in computation inspires hope that our theory could lead to improvements in optimization algorithms in non-conic settings where constraint qualifications may fail.
-Where facial reduction reduces to the minimal face of a cone, we reduce jointly to the smallest faces of two convex sets, both selected by intersection with the single joint supporting subspace.
-For simplicity, we frame our results as a reduction of the relevant objects (convex sets or convex functions), while keeping the ambient space $\mathbb{R}^{n}$ fixed.
-Our results are however naturally suited to a reduction of the ambient space to the joint supporting subspace, i.e., the joint supporting subspace would then become the new ambient space (see, e.g., Theorem 5.1).
-This alternative mathematical formulation would better model a reduction of the problem in a preprocessing step.
-Reduction of the ambient space to the joint supporting subspace in a preprocessing step may seem counterintuitive, as results in convex analysis are commonly specified only for real vector spaces, not affine spaces.
-This discrepancy can be mitigated by translating the affine subspace jointly with all the objects it contains until it reaches origin.
-But even this minor adjustment seems redundant, because, we believe, affine spaces seem to be the natural home for many results in convex analysis.
-In full abstraction, an affine space is a pair $(L,U)$ consisting of a set $L$ and a real vector space $U$ (the tangent space), equipped with an addition operation $L\times U\to L$ with the property that $\forall l_{1},l_{2}\in L$ there exists a unique $u\in U$ such that $l_{1}+u=l_{2}$.
-For $L$ an affine subspace of $\mathbb{R}^{n}$, the tangent space is $U=L-L=\{l_{1}-l_{2}\mid l_{1},l_{2}\in L\}$.
-Results in convex analysis can then be generalized by distinguishing tangent vectors from points in the affine space.
-For example, that affine subspaces are closed under convex combinations follows from reformulating $\lambda x+(1-\lambda)y$ for $x,y\in L$ as $y+\lambda(x-y)$, by noticing that $y\in L$ and that $\lambda(x-y)$ is in the tangent space LLL-L.
-The tangent space appears in our definition for the separation of convex sets (Definition 4.3), and justifies our repeated use of self-differences of sets SSS-S (see, e.g., Definition 6.2, Corollary 4.1, or Definition 4.2).
-1.2 Theoretical contribution
-An important contribution of the present work is the theoretical framework that we build in Section 6.
-The central objects are nested normals (Definition 6.3) and supporting subspaces (Definition 6.2), which describe local subsets of the lattice of faces of a convex set.
-Their combination forms a mathematical object that is an exact dual for convex cones (see Theorem 6.2).
-As such, the nested normals can also be understood as a local description of the boundary of convex sets, an idea that we discuss in Section 6.
-With these objects, we derive results describing the intersection of convex sets, as well as the facial structure of differences of convex sets.
-Notable technical results include Theorem 6.3, Theorem 6.4, and Theorem 6.6.
-We therefore contribute a geometric framework describing the boundaries, intersections, and differences of convex sets.
+$\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset,$
+when in fact, $\partial(f+g)(0)=\partial\iota_{\{0\}}(0)=\mathbb{R}$, so the sum rule does not hold in this case.
+There is no contradiction in this failure because the qualification condition in Equation 1 is not satisfied in this case (otherwise, the sum rule would be guaranteed by [23, Theorem 23.8]).
+Indeed, the relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect.
+Our approach generalizes facial reduction from the conic setting to general convex programs of the form f(x)+g(Ax)f(x)+g(Ax)—let us here consider the special case $f(x)+g(x)$.
+We identify a single affine subspace $T_{a}\subseteq\mathbb{R}^{n}$ that simultaneously regularizes both functions and contains the feasible region.
+The joint supporting subspace $T_{a}$ is defined for any two convex sets (see Definition 4.1), which are usually the domains of two convex functions.
+We can then reduce the effective domains of each function to $T_{a}$: $f^{\prime}:=f+\iota_{T_{a}},g^{\prime}:=g+\iota_{T_{a}}$.
+For these reduced functions, qualification conditions are then guaranteed to hold, while preserving the original problem in the sense that $f+g\equiv f^{\prime}+g^{\prime}$.
+With these reduced functions we modify the sum rule to resolve our previously encountered counterexample, by using the reduced functions on the r.h.s: $\partial(f+g)(x)=\partial f^{\prime}(x)+\partial g^{\prime}(x)$.
+In our counter-example, the joint supporting subspace is $T_{a}=\{0\}$, and the modified sum rule yields
+$\partial f^{\prime}(0)+\partial g^{\prime}(0)=\partial\iota_{\{0\}}(0)+\partial\iota_{\{0\}}(0)=\mathbb{R},$
+which indeed matches $\partial(f+g)(0)$.
+We note that the reduced functions $f^{\prime}$, $g^{\prime}$ are trivial in this example only because the feasible region is itself trivial.
+In general, our reduction will preserve the functions’ behaviour inside—and to some extent outside—the feasible region.
+In this manner, in Section 5 we generalize a collection of results from convex analysis, with the classical formulations trivially recovered when qualifications hold (see Proposition 5.1).
+Specifically, we find a modified Fenchel-Rockafellar dual for which strong duality always holds, problem-independent Karush-Kuhn-Tucker (KKT) optimality conditions, always-attained infimal convolution for the convex conjugates of sums of pairs of functions, qualification-free subdifferential sum and chain rules, and a characterization of the normal cones of the intersection of two convex sets.
+We provide a number of characterizations for $T_{a}$ in Section 4.
+We specify the joint supporting subspace $T_{a}$ in terms of the difference of domains $\operatorname{dom}(f)-\operatorname{dom}(g)$, where the set difference is $C-D:=\{c-d\mid c\in C,d\in D\}$.
+Other characterizations relate $T_{a}$ to minimal faces of the domains containing the feasible region (Theorem 4.1, Corollary 4.2, and Corollary 4.3).
+One of our characterizations (Theorem 7.1) consists of an iterative reduction procedure which recovers $T_{a}$ and terminates in a number of steps no more than the ambient dimension.
+This reduction can be performed from any point in the feasible region, using so-called nested normal cones (see Definition 6.3).
+We show that the reduction at a point does not in fact depend on the choice of the point, which means that the reduction need only be performed once for a given optimization problem.
+It is known that facial reduction in the conic setting leads to a reduction of the size of the problem and improved numerical stability [7], which gives reason to believe that our generalization may have similar benefits.
+A contribution of the present work is in our proof technique, found in Section 6.
+The central objects are nested normals (Definition 6.3) and supporting subspaces (Definition 6.2), which together describe the local lattice of faces of a convex set.
+We establish a structured induction via nested normals, which traverses all supporting subspaces, with induction steps associated with nested normals (see Corollary 6.1).
+With nested normals and our structural induction, we characterize the minimal supporting subspace containing a set (see Theorem 6.3).
+We also characterize the facial structure of a difference of convex sets CDC-D at 0 exclusively from the local facial structure of the individual convex sets $C$ and $D$ at $C\cap D$ (see Corollary 6.2).
+These results enable us to prove our characterizations of the joint supporting subspace that are presented in Section 4.
 1.1 Paper outline
@@
 In Section 3 we introduce some notation.
-In Section 4 we state our results describing the joint supporting subspace, with proofs deferred to Section 7.
-There, we define the joint supporting subspace in Definition 4.1, characterize it with a distinct analytical formula in Theorem 4.1, and with a constructive iterative procedure in Corollary 5.7 which terminates in a number of steps no more than the ambient dimension.
+In Section 4 we state analytic and constructive characterizations of the joint supporting subspace, with proofs deferred to Section 7.
 In Theorem 4.2 we state the main property of the joint supporting subspace: that qualification conditions hold when localizing to it.
 In Section 5, we state and prove generalizations of central results in convex analysis.
-In Section 5.3.1, we show that results of facial reduction in conic programming can be recovered from our results.
-In Section 6, we provide a narrative, motivated introduction to our mathematical framework, introducing key concepts and proving associated results along the way.
-In Section 7, we prove the results of Section 4 with the tools from Section 6.
+In Section 6, we build the core theoretical framework.
+In Section 7, we prove the results of Section 4 with the tools of Section 6.
 2 Prior works
-Constraint qualifications (CQs) in finite dimensions evolved from Slater’s condition [slaterLagrangeMultipliersRevisited1959], later weakened to the algebraic Mangasarian-Fromovitz CQ [mangasarianFritzJohnNecessary1967] and to the geometric Abadie CQ [abadieNonlinearProgramming1967].
-The weak Slater condition, introduced by Rockafellar [rockafellarConvexAnalysis1970], uses the notion of relative interiors.
-Subsequent relaxed CQs include the constant rank constraint qualification condition [janinDirectionalDerivativeMarginal1984], further relaxed in [qiConstantPositiveLinear2000, andreaniRelaxedConstantPositive2012], and a CQ by [robinsonFirstOrderConditions1976], which takes the form of a stability condition now known as linear metric subregularity.
-We note that in his seminal work [rockafellarConvexAnalysis1970], Rockafellar used the same qualification conditions for optimization problems and subdifferential calculus, which enabled our unified treatment of optimization and subdifferential calculus.
-There has been two major research directions generalizing results beyond qualification conditions.
-The geometric stream, to which our work belongs, focuses on the geometry of constraints and feasible sets.
-The sequential stream considers limits of increasingly accurate approximations, with exact results in the limit, and is related to variational analysis [mordukhovich_VariationalAnalysisApplications_2018, rockafellarVariationalAnalysis1997].
-The geometric stream focuses on constraint qualifications for optimization problems, while the sequential stream emphasizes subdifferential calculus, with extensions to sequential optimality conditions.
-Our work extends the geometric stream to subdifferential calculus.
-We survey qualification-free results in finite dimensions that are comparable to our results in Section 5.
-2.1 Geometric stream
-In the setting of inequality-constrained convex programs, the Fritz-John condition [johnExtremumProblemsInequalities2014] is an optimality condition that holds without qualification conditions.
-Unlike the KKT condition, it is not a sufficient condition for optimality, and is satisfied either when KKT conditions hold, or when some degeneracy in the constraints is detected, akin to applying the KKT conditions with a suitable constraint qualification.
-A qualification-free necessary and sufficient condition was formulated by [ben-talCharacterizationOptimalityConvex1976], and extended to the nonsmooth case by [wolkowiczGeometryOptimalityConditions1980].
-These works add a term to the KKT condition to deal with pathological cases; see [dharaOptimalityConditionsConvex2011, Chapter 6] for a treatment.
-Our qualification-free KKT condition in Corollary 5.6 is formulated for programs of the form $f(x)+g(Ax)$, instead of inequality-constrained convex programs, and uses faces of convex sets instead of an additional term.
-Facial Reduction (FR) generalized preprocessing ideas from linear programming to the setting of conic programming [massamOptimalityConditionsConeConvex1979, borweinFacialReductionConeconvex1981, borweinRegularizingAbstractConvex1981], with subsequent developments providing efficient computational methods for the reduction.
-An early method iteratively reduces the dimension of the problem [borweinRegularizingAbstractConvex1981, luoDualityResultsConic1997, wakiFacialReductionAlgorithms2013] until the minimal face is reached.
-Ideas of facial reduction can alternatively be framed with exact dual problems [borweinCharacterizationOptimalityAbstract1981], with important applications to semidefinite programming [ramanaExactDualityTheory1997, patakiStrongDualityConic2013].
-For in-depth literature reviews of facial reduction and its numerous applications, we refer the reader to [tuncelStrongDualityMinimal2012, drusvyatskiyManyFacesDegeneracy2017, permenterReductionMethodsSemidefinite2017].
-Many results in our work have analogues in facial reduction.
-For instance, the notion of a minimal face appears in different forms in Definition 4.1 and Theorem 4.1.
-The iterative reduction procedure of [borweinRegularizingAbstractConvex1981] finds its analogue in Theorem 7.1.
-We derive an exact Fenchel-Rockafellar dual in Corollary 5.5, echoing exact duals in conic programming.
-Finally, our qualification-free optimality condition in Corollary 5.6 generalizes similar results in the conic setting.
-2.2 Sequential stream
-The sequential approach to relaxing qualification conditions began with subdifferential calculus.
-One approach is that of $\varepsilon$-subdifferentials, which defines subgradients satisfying the subgradient inequality up to an $\varepsilon$-error, recovering the exact subdifferential in the limit of vanishing $\varepsilon$ [hiriarturruty_SubdifferentialCalculusUsing_1993, penotSubdifferentialCalculusQualification1996, mordukhovich_VariationalAnalysisApplications_2018].
-A second, similar approach is that of fuzzy sum rules, which approximate the subdifferential by perturbing the point $x$ at which the subdifferential is taken, deriving exact additivity rules in the limit of vanishing perturbations [jules_FormulasSubdifferentialsSums_2002, ioffeFuzzyPrinciplesCharacterization1998].
-These ideas were then applied to optimality conditions, where the condition of optimality takes the form of the existence of some sequence converging to the optimal point.
-One approach is to consider sequences of points converging to the point of interest [thibaultSequentialConvexSubdifferential1997], similarly to fuzzy sum rules, while another is analogous to $\varepsilon$-subdifferentials [jeyakumarInequalitySystemsGlobal1996, jeyakumarNewSequentialLagrange2003] relating them to the epigraphs of conjugate functions.
-2.3 Theory of convex faces
-A geometric insight that was critical to the development of our theory was the lexicographical characterization of faces [martinez-legazLexicographicalSeparationRn1987, martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_].
-[martinez-legazLexicographicalSeparationRn1987] derive a separation theorem from which the characterization follows.
-It turned out that the lexicographical framework was not flexible enough for our needs, prompting us to develop an analogous yet independent and self-contained theory.
-In this paper, we recover the lexicographical characterization of convex faces in Theorem 6.4, though in different terms.
-Our proof is direct, and does not make use of a separation theorem, though some separation ideas are implicit in the definition of supporting subspaces (Definition 6.1).
-3 Preliminaries
+Constraint qualifications (CQs) in optimization started with Slater’s condition [24] and much work has since been aimed at finding relaxations of these conditions [15, 1, 22].
+Notably, a qualification condition using relative interiors (the weak Slater) was introduced by [23], who used the same qualification conditions for optimization problems and subdifferential calculus, which enabled our unified treatment.
+There are a number of qualification-free alternatives to classical results, which come at the cost of some added complexity.
+One approach is to consider approximations of the core objects, with qualification-free exact results obtained in the limit of increasingly accurate approximation.
+This type of result includes $\varepsilon$-subdifferentials [8], and fuzzy sum rules [9, 11].
+Similarly, sequential optimality conditions have been formulated [25, 10].
+Sequential approaches are related to the broader theory of variational analysis.
+Another kind of qualification-free result is obtained by considering the geometry of domain boundaries, with the notable example of facial reduction in conic optimization.
+Our work is the first, to our knowledge, to extend this geometric methodology to qualification-free subdifferential calculus.
+In this vein, a qualification-free necessary and sufficient condition was formulated by [2], and extended to the nonsmooth case by [27], which achieve qualification-free characterization of minimizers by adding a term to the KKT conditions.
+Our qualification-free KKT condition in Corollary 5.6 is formulated for programs of the form $f(x)+g(Ax)$, instead of inequality-constrained convex programs, and uses a reduction instead of an additional term.
+Our theory extends Facial Reduction (FR) from conic programs to programs of the form $f(x)+g(Ax)$.
+The field of facial reduction started by generalizing preprocessing ideas from linear programming to the setting of conic programming [18, 3, 5], with subsequent developments providing efficient computational methods.
+One computational method consists in iteratively reducing the dimension of the problem [5, 14, 26] until the minimal face is reached, an approach that we generalize in Theorem 7.1.
+Ideas of facial reduction can alternatively be framed as exact dual problems [4], with important applications to semidefinite programming [21, 20].
+From convex geometry, a line of work that enabled the development of our theory was the lexicographical characterization of faces [17, 16].
+It turned out that the lexicographical framework was not flexible enough for our needs, prompting us to develop an analogous self-contained theory, recovering the lexicographical characterization of convex faces in Theorem 6.2 in different terms.
+Specifically, our framework based on nested normals allows us to specify a special kind of structural induction on supporting subspaces (and thereby faces) in Corollary 6.1.
+This induction is powerful because each inductive step is associated with a single nested normal.
+Additionally, the use of nested normals allows us to specify a reduction procedure that uses nested normal cones at each step, instead of individual nested normals (see Corollary 4.4 and Theorem 7.1).
+This formulation provides structural insight and allows, in most cases, termination in fewer steps by dropping many dimensions at once in some steps.
+Shortly after the first version of our work was made public, an independent paper by [12] was made available which specifies a reduction for convex programs of the form $f(x)+g(x)$.
+Their work has an algorithmic focus: they show feasibility detection and fast reductions with extended duals, by defining a notion of niceness for the domains of the convex functions.
+In contrast, our work is more geometric, providing analytical characterizations of the joint supporting subspace in Section 4 and qualification-free results stated with this object in Section 5, including a reduction for programs in the slightly more general form $f(x)+g(Ax)$, and subdifferential calculus sum and chain rules.
+Additionally, the proof techniques of the two papers seem remarkably distinct: they lift to cones, whereas we introduce nested normals in the original space.
+3 Notation and Preliminaries
 For two sets $A,B\subseteq\mathbb{R}^{n}$, $A+B:=\{a+b\mid a\in A,b\in B\}$ is the Minkowski sum, with $A-B:=\{a-b\mid a\in A,b\in B\}$, and for $z\in\mathbb{R}^{n}$, let $A+z:=A+\{z\}$.
@@
 We also denote $\operatorname{cone}(S)$ for the conical hull of a set; the smallest cone containing $S$.
-Intuitively the notion of a face of a convex set $C$ is a subset of $C$ for which there is no line segment in $C$ that “goes through” $F$.
-The following definitions can be found inline in [schneider_ConvexBodiesBrunn_2013, p.
-16] and [schneider_ConvexBodiesBrunn_2013, p.75] respectively.
-Definition 3.1 (Face of a convex set).
-Let $C$ be convex. A subset $F\subseteq C$ is defined to be a face of $C$ when
-$\forall y,z\in C,(\exists x\in]y,z[\cap F\implies y,z\in F).$
-Definition 3.2 (Exposed face of a convex set).
-Let $C\subseteq\mathbb{R}^{n}$ be convex. A subset $F\subseteq C$ is defined to be an exposed face of $C$ when
-$\exists u\in\mathbb{R}^{n}\setminus\{0\}:F=\operatorname*{\arg\max}_{c\in C}\langle c,u\rangle.$
+Let $C\subseteq\mathbb{R}^{n}$ be convex.
+We denote by $\operatorname{relint}(C)$ the relative interior of $C$.
+A subset $F\subseteq C$ is defined to be a face of $C$ when $\forall y,z\in C$, $(\exists x\in]y,z[\cap F\implies y,z\in F)$.
+That is, a face $F\subseteq C$ is a subset of $C$ for which there is no line segment that “goes through” $F$ while remaining in $C$.
+A face $F\subseteq C$ is exposed when there is a $u\in\mathbb{R}^{n}\setminus\{0\}$ such that $F=\operatorname*{\arg\max}_{c\in C}\langle c,u\rangle$.
 For any set $S\subseteq C$, we denote by $F_{C}(S)$ the face of $C$ generated by $S$; it is the smallest face of $C$ that contains $S$.
 (It is also the intersection of all faces containing $S$, which, it can be shown, is itself a face.)
-For $x\in\mathbb{R}^{n}$ and a convex set $C\subseteq\mathbb{R}^{n}$, the normal cone of $C$ at $x$ is
-$N_{C}(x)=\{y\mid\forall c\in C,\langle y,c-x\rangle\leq 0\}$
-if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise.
-We also recall some objects from convex analysis.
-Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$.
-For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
+For $x\in\mathbb{R}^{n}$ and a convex set $C\subseteq\mathbb{R}^{n}$, the normal cone of $C$ at $x$ is $N_{C}(x)=\{y\mid\forall c\in C,\langle y,c-x\rangle\leq 0\}$ if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise.
+We denote $\mathbb{R}_{+}:=[0,\infty)$, $\mathbb{R}_{++}:=(0,\infty)$, and similarly for $\mathbb{R}_{-}$ and $\mathbb{R}_{--}$.
+The extended reals are $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$.
+Given a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ we denote by $\operatorname{dom}(f)$ the effective domain of $f$; $\operatorname{dom}(f):=\{x\in\mathbb{R}^{n}\mid f(x)\neq\infty\}$.
+The function $f$ is proper if it is nowhere -$\infty$ and has nonempty effective domain.
+The subdifferential is
 $\partial f(x):=\{v\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{n},f(x)+\langle v,y-x\rangle\leq f(y)\},$
-with $\partial f(x)=\emptyset$ when $x\notin\operatorname{dom}(f)$. The convex conjugate is
-$f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x).$
+with $\partial f(x)=\emptyset$ when $x\notin\operatorname{dom}(f)$. The convex conjugate is $f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x)$.
 4 The joint supporting subspace
@@
 We also define its affine counterpart: $T_{a}(C,D):=T(C,D)+C\cap D$.
-The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact an affine translation. This follows from the next corollary.
+We note that when $C\cap D=\emptyset$, we have $T_{a}=\emptyset$ also.
+This is an intentional convention which takes care of the infeasible case in Section 5.
+The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact an affine translation.
+This follows from the next corollary.
 Corollary 4.1 (Containment of the intersection).
@@
 $T(C,D)=H_{C}(C\cap D)+H_{D}(C\cap D).$
-We provide two other characterizations that follow from Theorem 4.1.
 Corollary 4.2 (Joint supporting subspace as hull of faces).
@@
 $\operatorname{relint}(C)\cap\operatorname{relint}(D)\neq\emptyset.$ (2)
-It was shown in [rockafellarConvexAnalysis1970, Theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains.
+It was shown in [23, Theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains.
 We define the notions of separation, and proper separation below.
@@
 Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in a single hyperplane.
-Since separation follows from proper separation, and lack of proper separation is equivalent to intersection of relative interiors (by [rockafellarConvexAnalysis1970, Theorem 11.3]), the contrapositive statement is that
+Since separation follows from proper separation, and lack of proper separation is equivalent to intersection of relative interiors (by [23, Theorem 11.3]), the contrapositive statement is that
 $\text{lack of separation}\implies\text{intersection of relative interiors}.$ (4)
@@
 Proof of Corollary 4.5.
-The result follows from Theorem 4.2 by Equation 5, with the caveat that we defined separation in affine subspaces, which are not necessarily vector spaces, while in [rockafellarConvexAnalysis1970, theorem 11.3] this notion is defined in $\mathbb{R}^{n}$.
+The result follows from Theorem 4.2 by Equation 4, with the caveat that we defined separation in affine subspaces, which are not necessarily vector spaces, while in [23, theorem 11.3] this notion is defined in $\mathbb{R}^{n}$.
 Yet the argument still holds because the notions of separation, proper separation, and relative interiors are invariant under translation of $T_{a}$ to $T$.
@@
 5 Main results
-With Corollary 4.5 we extend standard results of convex analysis by first localizing to $T_{a}$.
-We thus generalize the following beyond qualification conditions: subdifferential additivity; normal cones of the intersection of convex sets; attained infimal convolution for the conjugate of a sum $(f+g)^{*}$; the subdifferential chain rule; Fenchel-Rockafellar duality; and KKT conditions.
-The next result implies that when standard qualification conditions hold, the classical formulations of our generalizations are trivially recovered.
+With Corollary 4.5 we provide qualification-free generalizations by localizing to $T_{a}$.
+The next result shows that when standard qualification conditions hold, the classical formulations are trivially recovered.
 Proposition 5.1 (Recovery of the classical results).
@@
 $\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x).$
+We note that in the infeasible case, when $f+g\equiv\infty$, $T_{a}$ is $\emptyset$, and the result holds trivially. A similar pattern occurs for many results in this section.
 Proof of Corollary 5.1.
-By Corollary 4.5 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$, the result follows by application of [rockafellarConvexAnalysis1970, Theorem 23.8] to the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$.
-∎
+By Corollary 4.5 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$, the result follows by application of [23, Theorem 23.8] to the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$. ∎
 Corollary 5.2 (Normal cone of the intersection of convex sets).
@@
 Proof of Corollary 5.2.
-With Corollary 4.5 we can use [rockafellarConvexAnalysis1970, Corollary 23.8.1], from which we find that
+With Corollary 4.5 we can use [23, Corollary 23.8.1], from which we find that
 $N_{C\cap D}(x)=N_{(C\cap T_{a})\cap(D\cap T_{a})}(x)=N_{C\cap T_{a}}(x)+N_{D\cap T_{a}}(x).$
@@
 If $\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)=\emptyset,$ then the result holds trivially because $g(Ax)\equiv\infty$, and $(g+\iota_{T_{a}})(Ax)\equiv\infty$ as well.
-Otherwise, by Corollary 4.5, $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$, which lets us apply [rockafellarConvexAnalysis1970, Theorem 23.9] to the composition $g+\iota_{T_{a}}]\circ A$.
+Otherwise, by Corollary 4.5, $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$, which lets us apply [23, Theorem 23.9] to the composition $g+\iota_{T_{a}}]\circ A$.
 Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Lemma 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$.
@@
 Whenever $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, we have that $T_{a}=\emptyset$, in which case the equality holds trivially, with the infimum attained at any $y\in\mathbb{R}^{n}$.
-Otherwise, by Corollary 4.5 the result follows from [rockafellarConvexAnalysis1970, Theorem 20.1].
+Otherwise, by Corollary 4.5 the result follows from [23, Theorem 20.1].
 ∎
@@
 $A(\operatorname{relint}\operatorname{dom}(f))\cap\operatorname{relint}\operatorname{dom}(g)\neq\emptyset,$
-which is a weakening of Slater’s condition. A program with this property is called strongly consistent [rockafellarConvexAnalysis1970, Theorem 31.2].
+which is a weakening of Slater’s condition. A program with this property is called strongly consistent [23, Theorem 31.2].
 The next result shows the sense in which 5.1 can be localized to the joint facial subspace $T_{a}$, guaranteeing strong consistency whenever the program is feasible, while retaining all the crucial information of the original program.
 Theorem 5.1 (Strong consistency of the localized program).
-For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ lower semicontinuous proper convex functions, a matrix $A\in\mathbb{R}^{n\times m}$, $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$, and any $x\in\mathbb{R}^{n}$,
+For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ lower semicontinuous proper convex functions, a matrix $A\in\mathbb{R}^{m\times n}$, $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$, and any $x\in\mathbb{R}^{n}$,
 $f(x)+g(Ax)=(f(x)+\iota_{A^{-1}(T_{a})}(x))+(g(x)+\iota_{T_{a}}(x)).$ (5)
@@
 $=$ $\operatorname{relint}(A\operatorname{dom}(f)\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (9)
-where Equation 8 holds by [rockafellarConvexAnalysis1970, Theorem 6.6]; Equation 10 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and that the set in Equation 10 is non-empty is guaranteed by Corollary 4.5.
+where Equation 7 holds by [23, Theorem 6.6]; Equation 9 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and that the set in Equation 9 is non-empty is guaranteed by Corollary 4.5.
 ∎
@@
 For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ l.s.c.
-proper convex functions, and a matrix $A\in\mathbb{R}^{n\times m}$, let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$.
+proper convex functions, and a matrix $A\in\mathbb{R}^{m\times n}$, let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$.
 Then
@@
 Proof of Corollary 5.5.
-The result follows from Theorem 5.1 with the use of [rockafellarConvexAnalysis1970, Corollary 31.2.1]. ∎
+The result follows from Theorem 5.1 with the use of [23, Corollary 31.2.1]. ∎
 Definition 5.1 (Minimizer).
@@
 Corollary 5.6 (KKT conditions).
-Consider 5.1, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.1 if, and only if, the following conditions hold:
+Consider 5.1, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.1 if, and only if, there exists some $y\in\mathbb{R}^{m}$ such that
   1. 1.
@@
 Proof of Corollary 5.6.
-From Theorem 5.1, the result follows from application of [rockafellarConvexAnalysis1970, Corollary 31.3.1] to the functions $f+\iota_{A^{-1}(T_{a})}$ and $g+\iota_{T_{a}}$. ∎
-Infeasible case as convention
-In the above results, the domains of the function under consideration may not intersect, in which case $f+g\circ A\equiv\infty$.
-Intersection is required for the existence of a meaningful joint supporting subspace, so it might be reasonable to make this non-trivial intersection a requirement.
-This kind of requirement was dubbed a “universal qualification condition” in the context of conic programming [tuncelStrongDualityMinimal2012].
-We observe that such a condition is not a true limitation of the theory, because the behaviour outside this condition is known and simple.
-Therefore, we instead use a convention: whenever $C\cap D=\emptyset$, we let $T_{a}(C,D):=\emptyset$ as well.
-As a result, $f+\iota_{A^{-1}(T_{a})}\equiv\infty$, and the equality in Corollary 5.5 trivially holds because both sides evaluate to $\infty$, with the value of $\infty$ attained in the dual.
-5.3.1 Facial reduction
-We derive two known facial reduction results, thereby showing that the joint supporting subspace is indeed a generalization of facial reduction.
-Problem 5.2 (Conic program).
-For $K\subseteq\mathbb{R}^{n}$ a closed convex cone, $A\in\mathbb{R}^{m\times n}$ a matrix, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$,
-$\mathop{\mathrm{minimize}}_{x\in\mathbb{R}^{n}}\langle x,c\rangle\quad\mathrm{s.t.}\quad Ax=b,\quad x\in K.$
-We apply Corollary 5.5 and Corollary 5.6 to this problem, yielding analogues of [tuncelStrongDualityMinimal2012, Theorem 4.3] and [tuncelStrongDualityMinimal2012, Corollary 4.4].
-The derived results differ from those in [tuncelStrongDualityMinimal2012] only in that our primal problem is their dual.
-Corollary 5.7 (Exact conic dual).
-For $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, $c\in\mathbb{R}^{n}$, and a closed convex cone $K\subseteq\mathbb{R}^{n}$,
-$\inf\{\langle x,c\rangle\mid x\in\mathbb{R}^{n},Ax=b,\text{ and }x\in K\}$
-$\max\{\langle b,y\rangle\mid y\in\mathbb{R}^{m}\text{ and }A^{T}y-c\in F^{\circ}\},$
-where $F=K\cap A^{-1}(H_{AK}(\{b\}))$.
-Proof of Corollary 5.7.
-We use Corollary 5.5 with $f(x):=\langle c,x\rangle+\iota_{K}(x)$, and $g:=\iota_{\{b\}}$. We find
-$T:=T(A\operatorname{dom}(f),\operatorname{dom}(g))=T(AK,\{b\})=H_{AK}(\{b\}),$
-where the last equality follows from Theorem 4.1.
-Since $AK$ is a cone, any face includes the origin, and therefore $0\in H_{AK}(\{b\})=T_{a}$, which means that $T_{a}=T$.
-The functions in the dual program are
-$f(x)+\iota_{A^{-1}(T)}(x)$ $=\langle c,x\rangle+\iota_{K}(x)+\iota_{A^{-1}(H_{AK}(\{b\}))}(x)$
-$=\langle c,x\rangle+\iota_{K\cap A^{-1}(H_{AK}(\{b\}))}(x)$
-and
-$g(Ax)+\iota_{T}(Ax)=\iota_{\{b\}}(Ax)+\iota_{H_{AK}(\{b\})}(Ax)=\iota_{\{b\}}(Ax).$
-Let $F=K\cap A^{-1}(H_{AK}(\{b\}))$, the subset of $K$ selected by the joint supporting subspace.
-The set $F$ is a cone, because $H_{AK}(\{b\})\cap AK$ is a face of the cone $AK$ by Definition 6.5 and Corollary 6.1, and $F$ is therefore itself a cone.
-Then with [rockafellarConvexAnalysis1970, Theorem 12.3, Theorem 14.1] we compute that
-$(f+\iota_{A^{-1}(T)})^{*}(A^{T}y)+(g+\iota_{T})^{*}(-y)$ $=(\langle c,\cdot\rangle+\iota_{F})^{*}(A^{T}y)+(\iota_{\{b\}})^{*}(-y)$
-$=\iota_{F^{\circ}}(A^{T}y-c)-\langle b,y\rangle.$
-We thereby find the dual problem, for which, by Corollary 5.5, strong duality holds, and the dual is attained. ∎
-Corollary 5.8 (Conic KKT conditions).
-A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.2 if, and only if, there exists some $y\in\mathbb{R}^{m}$ such that all of the following hold.
-  1. 1.
-    $Ax=b$.
-  2. 2.
-    $A^{T}y-c\in F^{\circ}$.
-  3. 3.
-    $\langle x,A^{T}y-c\rangle=0$.
-Proof of Corollary 5.8.
-We apply Corollary 5.6 to the case where $f(x):=\langle c,x\rangle+\iota_{K}(x)$ and $g:=\iota_{\{b\}}$.
-As in the proof of Corollary 5.7, we find that $f(x)+\iota_{A^{-1}(T_{a})}=\langle c,x\rangle+\iota_{F}(x)$ and $g(\bar{y})+\iota_{T_{a}}(\bar{y})=\iota_{\{b\}}(\bar{y})$.
-Then $(g+\iota_{T_{a}})^{*}(y)=\langle b,y\rangle$, and the first KKT condition, that $\partial(g+\iota_{T_{a}})^{*}(y)=\{b\}$ reduces to $Ax=b$. The second KKT condition is
-$A^{T}y$ $\in\partial(f+\iota_{A^{-1}(T_{a})})(x)$
-$=\partial(\langle c,\cdot\rangle+\iota_{F})(x)=c+N_{F}(x).$
-But since $N_{F}(x)=F^{\circ}\cap x^{\perp}$, this second condition holds if, and only if, both $x\perp A^{T}y-c$ and $A^{T}y-c\in F^{\circ}$. ∎
+From Theorem 5.1, the result follows from application of [23, Corollary 31.3.1] to the functions $f+\iota_{A^{-1}(T_{a})}$ and $g+\iota_{T_{a}}$. ∎
 6 A framework for local face lattices
-Inspired by the lexicographic characterization of convex faces of [martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_], we develop a framework describing the lattice of convex faces at a point $x\in C$, which we will then use to show the results of Section 4.
+Inspired by the lexicographic characterization of convex faces of [16], we develop a framework describing the lattice of convex faces at a point $x\in C$, which we will then use to show the results of Section 4.
 We begin with a characterization of faces of convex sets.
 In the next few results, we denote $X$ for a real vector space.
-Theorem 6.1 (Characterization of convex faces by supporting subspaces).
+Proposition 6.1 (Characterization of convex faces by supporting subspaces).
 Let $C\subseteq X$ be a convex set. A subset $F\subseteq C$ is a face of $C$ if, and only if the two following conditions hold.
@@
     $C\setminus F$ is convex.
-Proof of Theorem 6.1.
+Proof of Proposition 6.1.
 Let $F$ be a face. We argue by contradiction that $\operatorname{aff}(F)\cap[C\setminus F]=\emptyset$; let us assume that there is some $z\in\operatorname{aff}(F)\cap[C\setminus F]$.
@@
 ∎
-Theorem 6.1 motivates the next mathematical object.
+Proposition 6.1 motivates the next mathematical object.
 Definition 6.1 (Supporting subspace).
 Let $C\subseteq X$ be convex. A supporting subspace $L\subseteq X$ of $C$ is an affine subspace such that $C\setminus L$ is convex.
-Though supporting subspaces are affine subspaces, we omit the adjective “affine” for brevity, mirroring the term “supporting hyperplane”.
-The notion of supporting subspaces seems to be surprisingly sparse in the literature; the only previous appearance of this notion the author could find is in [brondstedIntroductionConvexPolytopes1983, Exercise 5.4].
-This idea is of central importance to this paper; it allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set.
+Though supporting subspaces are affine subspaces, we omit the adjective “affine” for brevity.
+This notion appears in [6, Exercise 5.4], and is related to order ideals, a concept defined for convex cones.
+Supporting subspaces are of central importance to this paper; it allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set.
 The next result describes the relationship between supporting subspaces and faces.
-Corollary 6.1 (Relation of supporting subspaces to faces).
-Let $C\subseteq X$ be convex and $S\subseteq C$ non-empty.
+Proposition 6.2 (Relation of supporting subspaces to faces).
+Let $C\subseteq X$ be convex.
   1. 1.
@@
 […]
-Proof of Corollary 6.1.
-1) ($\implies$) We show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Theorem 6.1:
+Proof of Proposition 6.2.
+1) ($\implies$) We show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Proposition 6.1:
   1. 1.
@@
 2) ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace.
-Indeed, by Theorem 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex.
+Indeed, by Proposition 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex.
 Note that $C\setminus F=C\setminus(C\cap\operatorname{aff}(F))=C\setminus\operatorname{aff}(F)$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, while $F=C\setminus\operatorname{aff}(F)$, which shows that $\operatorname{aff}(F)$ is a supporting subspace that selects $F$ by intersection.
@@
 We find it useful to generalize this idea of “locality” to supporting subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all supporting subspaces of $C$ that also contain $S$.
-To emphasize this connection, we say that a supporting subspace is ”at SS” to describe its containment of $S$.
+To emphasize this connection, we say that a supporting subspace is “at SS” to describe its containment of $S$.
 The reader should keep in mind the important special case of $S$ being a singleton $S=\{x\}$, for which the results are localized at a point.
@@
 In other words, the supporting subspace $U+S$ contains $S$.
-Next, we introduce a generalization of the notion of a normal cone. A normal cone of a convex set $C\subseteq\mathbb{R}^{n}$ is defined as
-$N_{C}(x):=\{z\in\mathbb{R}^{n}\mid\forall y\in C\langle y-x,z\rangle\leq 0\}.$
-When $C$ is a cone and $x=0$, the normal cone $N_{C}(0)=:C^{\circ}$ is called the polar of $C$. The polar is a dual operation, in the sense that
-${C^{\circ}}^{\circ}=\operatorname{cl}(C).$
-We can see that a normal cone is a dual object which “describes” the original cone only up to its closure.
-We require a dual object which has sufficient descriptive power to go beyond the closure, by encoding information about the boundary.
-To this end, we introduce the notion of a nested normal cone.
+We introduce the notion of a nested normal cone.
 Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
@@
 The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$.
-Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [lu_NormalFansPolyhedral_2008], where normal cones are defined for faces of polytopes, instead of at points.
+Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [13], where normal cones are defined for faces of polytopes, instead of at points.
 Combining the two generalizations, we get the following.
@@
 ∎
-The nested normals at a set are orthogonal with that set.
+Next, we show that the nested normals at a set $S$ are orthogonal with SSS-S.
 Lemma 6.2 (Normal cone everywhere is the orthogonal subspace).
@@
 ∎
-As previously discussed, one motivating property of the nested normals is that they can encode information about the boundary of convex sets. We make this idea concrete.
-The tangent cone of a convex set $C$ at $x\in C$ is the smallest closed cone $T$ such that $C\subseteq x+T$.
-We can think of this cone as a local description of $C$ around $x$.
-The normal cone is the polar of the tangent cone, and therefore “describes” the convex set $C$ around $x$ to the same extent as the tangent cone.
-In contrast, nested normals in supporting subspaces of $C$ at $x$ describe the cone of feasible directions.
-Definition 6.4 (Cone of feasible directions).
-For a convex set $C\in\mathbb{R}^{n}$ with $x\in C$, we define the cone of feasible directions of $C$ at $x$ to be
-$D_{C}(x):=\operatorname{cone}(C-x).$
-The closure of the cone of feasible directions is the tangent cone.
-The difference is that the cone of feasible directions captures additional information about the boundary of $C$ around $x$.
-The following result shows that any convex cone (including the cone of feasible directions at any point) can be recovered from the nested normals in supporting subspaces.
-Theorem 6.2 (Local polar without closure).
-Given any convex cone $T$ with $0\in T$,
-$T=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\in N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}.$
-Proof of Theorem 6.2.
-Let $T^{\prime}:=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\in N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}$.
-By the definition of nested normals, $T\subseteq T^{\prime}$.
-For the containment ($\supseteq$), we argue by contradiction: we assume that there is some $x\in T^{\prime}\setminus T$.
-If $x\in\mathrm{int}(T^{c})$, then $x$ is strictly separated from $T$ by a supporting hyperplane of $T$ (e.g., see [rockafellarConvexAnalysis1970, Theorem 11.3]).
-Any such hyperplane induces a nested normal of the form $v\in N_{T}^{\mathbb{R}^{n}}(0)$, for which $v^{\perp}$ is the hyperplane, and $\langle v,x\rangle>0$.
-But then $x\notin T^{\prime}$, and so $x\notin\mathrm{int}(T^{c})$.
-Since we assumed that $x\notin T$, the only possibility is that $x\in\partial T$.
-Then $x$ is contained in a supporting hyperplane, so there is some nested normal $v\in N_{T}^{\mathbb{R}^{n}}(0)$ such that $\langle v,x\rangle=0$.
-Let $T_{1}=T\cap v^{\perp}$, and narrow the ambient space to the subspace $T_{1}$.
-Inside $T_{1}$, we repeat the same argument, and keep narrowing the ambient space in this manner.
-This process must eventually reduce the ambient space to $\operatorname{\mathrm{span}}(x)$, because $\mathbb{R}^{n}$ is finite-dimensional.
-In this one-dimensional space, with the narrowed cone $T_{\ell}=T\cap\operatorname{\mathrm{span}}(x)$, note that $T_{\ell}$ is a closed cone.
-Therefore, either $x\in T_{\ell}$ or $x\in\mathrm{int}(T_{\ell}^{c})$, both of which yields a contradiction, from which the result follows.
-∎
-Another important property of nested normals is that they describe the containment structure of supporting subspaces, which is closely related to the lattice structure of the faces.
-We make this precise in the next two results.
+The core property of nested normals is that they describe the lattice structure of supporting subspaces, which is closely related to the lattice structure of the faces.
+We make this idea precise in the next two results.
 Theorem 6.1 (Nested normals induce supporting subspaces).
@@
 $(V\cap v^{\perp}+S)\cap C=\operatorname*{\arg\max}_{y\in C\cap(V+S)}\langle y,v\rangle.$
-The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, which by Corollary 6.1 implies that $V\cap v^{\perp}\in\mathcal{F}(C;S)$. ∎
+The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, which by Proposition 6.2 implies that $V\cap v^{\perp}\in\mathcal{F}(C;S)$. ∎
 We next show that all supporting subspaces are “reached” by nested normals in supporting subspaces.
@@
 Then there is a sequence $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)$ with $k=\dim(V)-\dim(U)$ such that $V_{0}=V$, $V_{k}=U$, and for each $i\in\{1,\ldots,k\}$ there is a $v_{i}\in N_{C}^{V_{i-1}}(S)\setminus\{0\}$ such that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$.
-This result is similar to the lexicographic characterization of the faces of convex sets in [martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face.
+This result is similar to the lexicographic characterization of the faces of convex sets in [16, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face.
 Let us state a technical lemma that we will need for the proof.
@@
 Note that the line segment $[y,z]$ has endpoints $y,z\in C^{\prime}\setminus(U+S)$ with a point $w\in]y,z[$ that is also contained in $C^{\prime}\cap(U+S)$.
-This implies that $C^{\prime}\cap(U+S)$ is not a face of $C^{\prime}$, which, by Corollary 6.1, means that $U\notin\mathcal{F}(C^{\prime};S)$.
+This implies that $C^{\prime}\cap(U+S)$ is not a face of $C^{\prime}$, which, by Proposition 6.2, means that $U\notin\mathcal{F}(C^{\prime};S)$.
 This is a contradiction because in fact, $U\in\mathcal{F}(C^{\prime};S)$, as we now show.
@@
 is the intersection of two convex faces of $C$.
-Their intersection is then also a face (this is immediate from Definition 3.1).
-By Theorem 6.1, $C\setminus[(U\cap V+S)\cap C]$ is convex.
+Their intersection is then also a face (immediate from the definition of faces).
+By Proposition 6.1, $C\setminus[(U\cap V+S)\cap C]$ is convex.
 The convexity of this set, which we can write as $C\setminus((U\cap V)+S)$, implies that $U\cap V\in\mathcal{F}(C;S)$.
@@
 Proof of Lemma 6.5.
-By Theorem 6.1, $H_{C}(S)$ being the smallest supporting subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own supporting subspace (by Corollary 6.1), which would have to be properly contained in $H_{C}(S)$.
-For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a supporting subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is a supporting subspace, as implied by Theorem 6.1.
+By Proposition 6.1, $H_{C}(S)$ being the smallest supporting subspace implies that $C\cap(H_{C}(S)+S)$ is the smallest face containing $S$, because any properly contained, smaller face would have its own supporting subspace (by Proposition 6.2), which would have to be properly contained in $H_{C}(S)$.
+For the second statement, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is a supporting subspace because $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))+S=\operatorname{aff}(F_{C}(S))$, which is a supporting subspace, as implied by Proposition 6.1.
 Further, $\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$ is contained in any other supporting subspace in $\mathcal{F}(C;S)$, because given $U\in\mathcal{F}(C;S)$ such that $U\subsetneq\operatorname{\mathrm{span}}(F_{C}(S)-F_{C}(S))$, we would then have $C\cap(U+S)$ as a face of $C$ containing $S$ that is properly contained in $F_{C}(S)$, in contradiction with the definition of $F_{C}(S)$.
@@
 Theorem 6.3 (The generated supporting subspace is the only supporting subspace with no nested normals).
-Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated supporting subspace $H_{C}(S)$ is the only supporting subspace of $C$ at $S$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
+Given a convex set $C\subseteq\mathbb{R}^{n}$ with $S\subseteq C$ non-empty, the generated supporting subspace $H_{C}(S)$ is the only element in $\mathcal{F}(C;S)$ such that $N_{C}^{H_{C}(S)}(S)=\{0\}$.
 Proof of Theorem 6.3.
@@
 6.2 Nested normals of a difference of sets
-In this section, we show the following important result.
-Theorem 6.6 (Characterization of the nested normals of a difference of sets).
-Let $C,D\subseteq\mathbb{R}^{n}$ be convex with $S\subseteq C\cap D$ non-empty, and $U\in\mathcal{F}(C-D;0)$. Then $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$, and
+We characterize the supporting subspaces of a difference of sets CDC-D from the supporting subspaces and nested normals of the individual sets $C$ and $D$ at $C\cap D$.
+One might hope for the simple statement that $\mathcal{F}(C;S)\cap\mathcal{F}(D;S)=\mathcal{F}(C-D;0)$ for $S\subseteq C\cap D$.
+Unfortunately, though $(\supseteq)$ turns out to hold, $(\subseteq)$ does not in general.
+Example 6.1 (Mismatched nested normals).
+Let $h:\mathbb{R}\to\mathbb{R},h(x):=x^{2}$, $C:=\operatorname{epi}\min(\iota_{\mathbb{R}_{-}},h)$ and $D:=\mathrm{hypo}\max(-\iota_{\mathbb{R}_{-}},-h)$, where $\mathrm{hypo}$ denotes the hypograph of the concave function.
+Then $\{0\}$ is a face of both $C$ and $D$, yet $C-D=\mathbb{R}\times\mathbb{R}_{+}$, and therefore $\{0\}\notin\mathcal{F}(C-D;0)$.
+There is therefore a need for some additional condition on supporting subspaces shared by $C$ and $D$ for them to also be supporting subspaces of CDC-D.
+The next result allows us to derive such a condition by using nested normals.
+Theorem 6.4 (Decomposition of supporting subspaces of differences).
+Let $C,D\subseteq\mathbb{R}^{n}$ be convex with $S\subseteq C\cap D$ non-empty, and $U\in\mathcal{F}(C-D;0)$. Then
+  1. 1.
+    $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S).$
+  2. 2.
     $N^{U}_{C-D}(0)=N^{U}_{C}(S)\cap(-N^{U}_{D}(S)).$
+  3. 3.
+    $C\cap(U+S)-D\cap(U+S)=(C-D)\cap U.$
+The next corollary shows that the missing condition is the existence of a synchronized “path” of nested normals for $C$ and $D$.
+Corollary 6.2 (Characterization of the faces of differences of convex sets).
+Let $C,D\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C\cap D$ non-empty.
+The elements of $\mathcal{F}(C-D;0)$ are precisely the elements $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$ for which there exists a sequence $V_{0},\ldots,V_{k}\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S)$ such that $V_{0}=\mathbb{R}^{n}$, $V_{k}=U$, and such that for each $i\in\{1,\ldots,k\}$ there is a $v_{i}\in N_{C}^{V_{i-1}}(S)\cap-N_{D}^{V_{i-1}}(S)\setminus\{0\}$ such that $V_{i}=V_{i-1}\cap v_{i}^{\perp}$.
+Proof of Corollary 6.2.
+That any $U$ with such a sequence is in $\mathcal{F}(C-D;0)$ follows by first noticing that each $v_{i}$ in the sequence belongs to $N^{U}_{C-D}(0)$ by Theorem 6.4.
+By applying Theorem 6.1 sequentially, it follows that $U\in\mathcal{F}(C-D;0)$.
+That any element $U\in\mathcal{F}(C-D;0)$ has such a sequence follows from Theorem 6.2, which shows that there is a sequence of nested normals of CDC-D at 0 that, starting from $\mathbb{R}^{n}$, attains $U$.
+By Theorem 6.4, this sequence corresponds to a sequence $v_{i},\ldots,v_{k}$ as described in the statement.
+∎
+We can now better understand Example 6.1.
+The supporting subspace $\{0\}$ of $C$ and $D$ is obtained via the sequence of nested normals $\{(0,-1),(1,0)\}$ for $C$ and $\{(0,1),(1,0)\}$ for $D$.
+Only the first nested normals are opposite to each other, and therefore there is no synchronized path of nested normals to $\{0\}$.
+If we modify the example by taking $D:=\mathrm{hypo}\max(-\iota_{\mathbb{R}_{+}},-h)$, then $\{0\}$ is attained by $\{(0,1),(-1,0)\}$, and the sequence is synchronized with that of $C$.
+In this modified example, $\{0\}$ is indeed a supporting subspace of CDC-D because CDC-D is a convex set such that $C-D\subseteq\mathbb{R}\times\mathbb{R}^{+}$ and $C-D\cap(\mathbb{R}\times\{0\})=\mathbb{R}_{-}\times\{0\}$, which shows that $\{0\}\in\mathcal{F}(C-D;0)$.
 To show Theorem 6.4, we require the following lemma.
@@
 Proof of Theorem 6.4.
-We argue by structured induction over the supporting subspaces of CDC-D, as described in Corollary 6.2.
-Fixing some $S\subseteq C\cap D$, the statement $A:\mathcal{F}(C-D;0)\to\{\text{true},\text{false}\}$ that we show is the following: $A(U)$ is true if, and only if, all of the following hold.
-  1. 1.
-    $U\in\mathcal{F}(C;S)\cap\mathcal{F}(D;S).$
-  2. 2.
-    $N^{U}_{C-D}(0)=N^{U}_{C}(S)\cap(-N^{U}_{D}(S)).$
-  3. 3.
-    $C\cap(U+S)-D\cap(U+S)=(C-D)\cap U.$ The additional point 3) has a supportive role in this proof.
+We argue by nested normal induction (from Corollary 6.1) over the supporting subspaces of CDC-D.
+Fixing some $S\subseteq C\cap D$, let $A(U)$ be the statement that all three points in the statements hold for supporting subspace $U\in\mathcal{F}(C-D;0)$.
 The base case, where $U=\mathbb{R}^{n}$, holds because $\mathbb{R}^{n}$ is always a supporting subspace. Point 2) is given by Lemma 6.6, and point 3) holds trivially.
@@
 $=N_{C\cap(U+S)-D\cap(U+S)}(0)\cap U$ (12)
-$=N_{(C-D)\cap U}(0)\cap U$ (14)
-$=N_{C-D}^{U}(0),$ (15)
+$=N_{(C-D)\cap U}(0)\cap U=N_{C-D}^{U}(0),$ (13)
 where Equation 11 holds by definition of nested normals, Equation 12 follows from Lemma 6.6 applied to the sets $C\cap(U+S)$ and $D\cap(U+S)$, and Equation 13 follows from point 3) of $A(U)$.
@@
 We now show the main property of the joint supporting subspace, Theorem 4.2.
-We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [mordukhovich_VariationalAnalysisApplications_2018, Theorem 2.19].
+We first state and prove a lemma which connects the notion of separation of convex set to a statement reminiscent of the qualification condition by Mordukhovich in variational analysis: [19, Theorem 2.19].
 Lemma 7.1 (Separation from local normal cones).
@@
 $=\operatorname{\mathrm{span}}(\operatorname{\mathrm{span}}(\bar{F}_{C})\cup\operatorname{\mathrm{span}}(\bar{F}_{D}))+C\cap D$
-$=\operatorname{\mathrm{span}}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
-$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
-$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D}+C\cap D)$
-$=\operatorname{aff}(F_{C}\cup F_{D}).$
+$=\operatorname{\mathrm{span}}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
+$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D}+C\cap D)=\operatorname{aff}(F_{C}\cup F_{D}).$
 7.3 Second characterization
@@
 ∎
+7.4 Conclusion
+In this work we define and characterize the joint supporting subspace $T_{a}$, which exists between any two convex sets.
+By defining this affine subspace from the domains of two convex functions, and then reducing each function to it, qualification conditions are guaranteed to hold.
+We obtain a number of characterizations for $T_{a}$ by bridging between the facial structure of the difference of two convex sets CDC-D and the facial structure of the individual convex sets.
+By doing this, we show that “pathology on the boundary” of convex functions can be tamed by considering the facial structure of their domains.
+Future work avenues include specifying and implementing efficient reduction algorithms for convex functions, and pairing them with optimization solvers to obtain robust algorithms with strong theoretical guarantees.
 7.5 Acknowledgements
@@
 Declarations
-The author is supported by the University of British Columbia. The author has no relevant financial or non-financial interests to disclose.
+The author is supported by the University of British Columbia.